A Complete Bibliography of Publications in History and Technology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

24 May 2024 Version 1.22

Title word cross-reference

∘ [110]. ^{e} [100, 95].

11 [410]. 1650s [512]. 16c [725]. 1750s [841]. 1820s [28]. 1840s [282]. 1873 [732]. 1873-1947 [872]. 1880s [610]. 1881 [17]. 1890s [309, 311]. 18c [725]. 18th [360, 376, 173]. 18th-century [376]. 19 [836]. 1906 [480]. 1910 [567]. 1920s [309, 437]. 1930s [214, 601]. 1940s [311, 703]. 1940s-1960s [703]. 1950s [400, 641, 339, 854]. 1956 [707]. 1960s [400, 314, 763, 608, 703, 395, 841, 428]. 1970s [314, 214, 670, 521]. 1980s [670]. 1990s [704, 854]. 19th [497, 294, 202, 505, 141, 165, 169, 121, 509, 173].

2 [639, 559]. 20th [458, 202, 185, 165, 459, 509, 496, 457]. 2ème [12]. 2nd [153].

3 [833].

58 [838].

70* [415].

Aakash [721]. abatement [620, 734]. abbé [139]. Abbott [139]. absorption [196]. Abstract [148]. Abstracts [38, 47, 52, 58, 66, 72, 118, 129, 133, 137, 142, 155, 162, 174, 180, 192, 203, 230, 235, 240]. abuses [258]. academic [747, 746]. Academy [191, 153]. Accelerator [758, 395]. Accelerators [262, 214]. Accepting [508]. access [308]. according [7, 12]. account [773]. accumulation [358]. acid [704]. Acknowledgements [220]. activism [406]. activities [614]. activity [186, 749]. actor [700]. Ad [19]. Adelheid [684]. administration [519, 388, 481, 580]. adoption [345]. Advance [796]. Advancement [189]. Advertising [670, 734, 227]. advocacy [846]. AEC [331]. Aerial [665, 601]. aeroplane [263]. aerospace [621]. Aerospaces [669]. aesthetic [226]. aesthetics [722, 395, 97]. affairs [530]. Affiliates [481]. Afghan [534]. aflatoxin [804]. Africa [876, 575, 837, 796, 883, 770]. African [727, 576]. After [507, 422, 855, 575, 581, 873, 295]. aftermath [410]. against [840]. Age [467, 633, 705, 730, 372, 398, 830, 757, 781]. Agency [833]. agenda [838, 511]. agent [61, 796]. Ages [18]. agricoles [239]. Agricultural [820, 760, 883, 822, 847, 239, 681]. Agriculture [680, 795]. AI [886]. aids [809]. Air [228, 566, 736, 856]. air-space [856]. Aircraft [869, 217]. airline [371, 534]. Airlines [534]. airplanes [539]. alcohol [846]. Alexander [101]. Alexandre [101, 107, 84, 106, 78, 79, 81, 80]. Algeria [726]. alienated [271]. Aligning [826]. alizarin [264]. all-seeing [639]. alleged [700]. allemands [85]. alleys [812]. allies [371, 370]. Alois [748]. Along [23]. alternating [176]. alternative [581]. amateur [347]. ambassadors [834]. ambitions [826, 832]. America [482, 79, 49, 255, 80, 731, 847, 582]. américains [62]. American [371, 60, 438, 676, 281, 502, 370, 372, 63, 345, 754, 727, 481, 619, 278, 812, 412, 415, 795]. Americanise [481]. Ammonia [470]. among [104, 757]. anachronism [785]. analogue [244]. analysis [113, 813, 685, 729, 44, 693, 629, 688]. Analytical [21]. Ancient [161]. Andre [700, 197]. Androids [684]. anglaise [177]. Angleterre [140]. Anglo [371, 370, 372, 415]. Aniline [482]. animal [277, 773]. Annales [439]. années [198]. answer [365]. Antarctic [530]. Anthropocene [886]. Anthropophagy [775]. anti [580]. anti-ballistic [580]. antibiotic [546]. antilock [384]. antiquity [257]. any [785]. Apollo [483, 260]. apologetics [139]. apologetique [139]. apparatuses [583]. appliance [283]. applications [246]. Applied [34, 811, 274, 68, 673]. Approaches [329, 729, 528]. appropriation [516]. après [248]. après-guerre [248]. Aragon [540]. Araucanía [690]. Archeology [496]. architectes [125]. architects [125]. architecture [522, 669, 709]. Archival [332]. archives [147]. Argentina [869]. Ariana [534]. Army [839, 285]. Arnold [641]. Arpanet [245]. Arsenal [461]. art [395]. artefacts [221]. artesian [196]. artésiens [196]. artfulness [601]. Arthur [593, 727, 635]. articles [654]. artifacts [675]. artifactual [436]. artificial [868, 113, 236, 808, 806, 757, 264]. artificielle [236]. Artisans [684]. Arts [680, 122, 402, 121, 264]. Ashes [681]. Asia [512, 823, 822, 858]. Asian [716]. Asphalt [436]. assessing [846]. assessment [571]. Assistance [834, 826, 821, 825, 819, 824]. Association [132, 730]. assumptions [68]. astroculture [633]. Astronomical [206]. ateliers [7, 12]. Athénée [126]. Atlantic [628]. atom [259, 838]. Atomic [144, 833, 709, 834, 421, 464, 535, 420]. Atoms [418, 710, 591, 565]. attempts [211]. attendant [595]. attractions [630]. Attributing [471]. audio [737]. aujourd’hui [77]. Australia [259]. Australian [259, 272]. Austria [310]. Author [564, 626, 646, 600, 572, 586]. authority [783]. autographic [152]. autographique [152]. automata [686, 808]. automation [190, 641]. automobile [197]. Automobility [582, 583, 586]. ’automobilization [585]. Aviation [438, 485, 371]. awkward [708].

B [207, 258]. Back [505, 725, 216]. back-end [216]. backstage [831]. backwardness [397]. baffling [415]. Baker [291]. ballistic [580]. Baltic [781]. Bank [476]. barbers [552]. Bartolomé [772]. Based [473]. basic [741, 33]. basics [725]. Basin [842]. Bataan [798]. BBSes [793]. be [637, 535, 462]. bearers [663]. beautiful [462]. became [516]. Beck [505]. Becker [796]. Beckman [641]. Beckmann [233, 234]. becoming [343, 803]. Before [863, 507, 151, 396]. beginning [145, 590]. Beginnings [161]. behind [838, 448]. Belgian [268]. Belle [506, 615]. Belle-Epoque} [506]. Belleville [57]. belts [377]. Benedictus [93]. Benin [837]. Berlin [115]. Besson [131]. beton [170]. Between [84, 436, 788, 264, 485, 433, 337, 728, 294, 413, 518, 359, 61, 475, 243, 245, 341, 572, 532, 167, 364, 694, 521, 56, 258, 233, 234]. Beyond [643, 735, 15, 747, 591]. big [212, 210, 755, 376, 207, 457, 218]. biofuels [847]. Biogen [751]. biographies [191]. biological [517, 795, 258]. biology [751, 802]. biomedical [537]. biometric [718]. biopolitics [564]. Biosphere [559]. Birth [195, 29, 136, 396]. Bishop [101]. Black [870, 888, 602]. blame [624]. blanket [422]. blind [571, 809]. blots [324]. Board [587, 617, 695, 723, 744, 25, 32, 40, 48, 53, 59, 67, 74, 112, 119, 130, 157, 163, 175, 182, 193, 204, 287, 293, 298, 307, 317, 323, 328, 350, 355, 362, 368, 375, 381, 387, 393, 833]. boast [463]. Bob [532]. bodily [878]. bogus [743]. boiler [136, 57]. bollards [613]. bomb [423]. Bombay [881]. Bombay/Mumbai [881]. bombs [613]. Book [321, 335, 348, 373, 379, 391, 154, 416, 426, 431, 292, 297, 515, 117, 316, 305, 385, 344, 366, 531, 324]. books [44]. border [185]. borderline [1]. Borders [836]. Bosch [384]. Bossut [23]. boundaries [308, 428]. Boundary [433]. Bourbon [126]. Bourses [444]. box [523]. boyhood [539]. Brahe [783]. brains [168]. braking [384]. brand [659]. branding [653]. Bray [884]. Brazil [811]. bread [535]. breast [406]. breeding [868]. Bricks [436]. bridge [138, 135, 345]. bridges [125]. Britain [567, 818, 673, 81, 742, 320, 882, 27, 674, 360, 423, 636, 791, 734, 731, 738, 691, 222]. British [289, 132, 855, 730, 637, 749, 856, 253, 718, 733, 528]. broken [264]. brokers [444]. Buckets [613]. build [529, 733, 591]. Building [345, 706, 712, 436, 173, 740, 166, 731, 275, 604, 30, 752]. built [214]. Bulferretti [65]. Bumped [401]. bureau [122]. Business [177, 256, 250, 276, 124, 528, 290]. Butterfield [683]. buyer [372].

C [558, 635, 359, 805, 854]. calandre [177]. Calcutta [161]. California [669]. camera [722]. Cameralism [233]. camouflets [882]. Campaign [650]. can [463, 430]. Canada [809]. Canal [596, 42, 23, 42]. Canaries [882]. cancer [406]. Canton [752]. Canyon [438]. capitalism [517, 10]. car [620]. carbon [882]. card [246]. cardiac [655]. cards [527]. care [841]. career [818, 285]. carence [95]. Caribbean [767]. Carny [236]. cartel [697]. Carver [390]. carving [856]. Case [471, 705, 397, 676, 777, 271, 559, 213, 748, 382, 227, 207, 318, 595]. Catalonia [782]. Catastrophes [438]. categorisation [791]. category [195, 729]. cattle [868]. CB [792]. CCTV [666]. ce [109]. Celebration [228, 480]. cells [651]. cement [50]. censoring [739]. censures [101]. center [190, 409, 537]. centers [62]. Central [575, 770, 608, 650]. centres [62]. centrifuge [415]. Centuries [509, 496, 18, 202, 747, 188]. Century [497, 505, 539, 177, 458, 294, 673, 185, 273, 100, 837, 288, 516, 517, 269, 141, 165, 360, 424, 169, 274, 750, 166, 95, 524, 746, 459, 628, 11, 121, 753, 299, 376, 325, 614, 679, 540, 383, 167, 318, 605, 682, 457, 733, 795, 223, 173, 851]. CERN [609]. cervelle [168]. César [337]. chains [445]. challenge [248, 250, 635]. challenges [845]. Chandlerian [276]. Chandra [596]. change [356, 86, 742, 704, 231, 781, 151]. Changing [473, 99, 523, 34]. character [662]. Charles [338]. Chattopadhyaya [161]. chaudière [136, 57]. Chaussées [125]. cheaper [511]. Chemical [506, 479, 470, 265, 34, 365, 756, 261]. chemicals [278]. chemist [153]. chemistry [34, 56]. Chemists [261, 546, 847]. chemotherapy [56]. chemurgy [847]. Chernobyl [650]. Children [650]. Chile [844]. Chilean [690]. chimurenga [713]. China [864, 861, 806, 866, 752]. Chinese [707]. chocolate [592]. choice [208, 269, 215]. Choosing [212, 207, 218]. Choson [860, 862]. Chris [567]. chromolithography [615]. church [97]. Ciba [656]. Cie [135]. cinchona [697]. cinema [602]. circa [734, 731]. Circuits [790]. circulations [771]. citrus [775]. city [568, 224, 734, 752, 812]. Civil [440, 217, 138, 50]. Clarke [635]. classic [21]. classical [234]. classification [804]. classiques [103]. Clean [226, 543]. cleaning [326]. Clemens [153]. climate [704]. clinical [340]. clock [722]. cloning [775]. closed [643]. cloth [135, 773]. co [829]. co-production [829]. Co. [135, 290]. coal [268, 852, 365]. coal-mining [268]. code [453, 247]. code-switching [453]. Cognition [778]. Cold [464, 711, 378, 669, 714, 842, 700, 534, 707, 296, 591]. Cold-War [378]. Collective [598]. College [390]. Colliery [338]. Colonial [272, 630, 575, 711, 883, 717, 842, 838, 771, 774, 577, 798, 797, 697]. colonialism [458, 820]. colony [574]. combat [735]. comes [259]. comfort [788]. comme [64]. Comment [557]. commentary [583]. commerce [278]. commercial [749, 353, 253]. commercialization [751]. Commercializing [747, 746, 750]. Commission [608]. commodity [803]. commons [537]. communicating [732]. communication [737, 208, 792]. Communications [839, 209, 742, 114, 245, 739, 733]. communities [816, 698, 325]. Community [608]. companies [265]. company [413, 657]. comparative [308]. compared [208]. comparison [413, 320]. competence [608]. Competition [608, 436, 236, 236]. complex [711]. compliance [388, 691]. Component [353, 254, 545]. compromise [835]. Computation [886]. Computer [333, 473, 334, 529, 246, 245, 254, 429, 691]. Computer-Based [473]. computers [283, 330, 331, 244]. computing [329, 242]. conceit [832]. concept [100, 525, 28, 867]. Conceptions [498]. concepts [245]. concerned [372]. Concrete [436, 170, 171]. condenser [51]. Condorcet [23]. conference [248]. Confiscated [278]. confronts [423]. conjunction [208]. connecting [817, 768]. connection [304]. connects [773]. Connelly [560]. Conquest [705]. consequence [310]. Conservatoire [121]. Considerations [82, 82]. Consortium [609]. constructing [384]. Construction [501, 677, 138, 135, 653, 658, 165, 166, 523, 741, 167, 443, 173, 165]. constructions [172]. consultants [746]. consultation [122]. consumer [539, 419, 459]. consumers [855]. consumption [584]. container [823]. containers [841]. contamination [543]. contemporary [699, 721]. contentious [619]. contesté [237]. Context [477, 27, 357]. continents [773]. contingencies [877]. continued [857]. continuité [110]. Continuity [102, 110]. continuous [703]. continuum [406]. contraceptive [340]. Contractors [470]. contre [109]. contre-révolutions [109]. contribution [138]. Contributors [111, 411, 427, 432, 306, 322, 327, 336, 349, 354, 361, 367, 374, 380, 386, 392]. Control [560, 438, 562, 396]. controllability [832]. controlled [864, 697]. Controlling [508, 543]. Controversy [412, 23]. conventions [647]. conversion [738]. cooker [863]. cooking [703]. cooperation [403, 370, 260]. Copernic [97]. Copernicus [97]. core [218]. corollaries [598]. CORONALAG [879]. Corporate [227, 364, 786]. Corporation [481]. Corps [497]. corpse [774]. Correction [888, 887]. correlation [56, 140, 140]. Correspondence [425, 524]. cortisone [658]. COS [207]. COS-B [207]. cosmos [634, 139]. Cost [502]. Cotten [582, 583]. councilors [125]. counter [109]. counter-revolutions [109]. countries [825]. country [62, 755, 710]. County [338]. Cover [807, 853, 764, 784]. COVID [836]. COVID-19 [836]. cracked [666]. Crafting [636]. craftsman [238]. Craig [727]. Creating [825, 56]. creation [326, 673, 767, 619, 720]. creative [713]. credible [629]. credit [624, 360]. Criollos [501]. crisis [18, 876, 15, 666, 365, 581]. Critical [324, 479, 867, 257, 231]. critics [276]. critique [700]. crossing [308]. crossroads [534]. Crown [540]. Crunching [331]. crusade [629]. crystal [249]. Cuba [425, 405]. cult [272]. cultivated [804]. Cultivating [634]. Cultural [497, 582, 440, 62, 571, 685, 688]. culture [539, 378, 37, 420, 312, 376, 792]. culturelle [62]. Cultures [684]. Curie [300, 301]. Curies [304, 303]. current [231]. cybernetics [645]. cycle [216]. cycles [681]. cyphered [772].

d’absorption [196]. dairy [868]. D’Alembert [23]. danger [592]. dans [236, 11]. d’après [7, 12]. dark [737]. data [843, 880, 886]. daughters [311]. Daumas [19]. dawn [372]. day [461]. Days [662]. DDT [855]. Deal [847]. death [616]. d’eau [57]. débat [77]. debate [452, 77, 813, 429]. Debiprasad [161]. decade [422]. Decadence [480]. Decision [481, 213, 214]. decision-making [213, 214]. decisions [816]. Decline [480]. Decoding [312]. decolonization [711]. Deep [437, 383]. Defence [253, 736]. défi [248]. Defiance [691]. deficiency [95]. Defining [383, 708]. definitions [9]. Delaware [34]. demobilize [430]. Democracy [461]. demon [421]. d’encouragement [147, 124]. Denmark [246, 565]. dentistry [728]. departures [276]. Depression [18, 60]. Depth [399]. depuis [166]. dernières [198]. Desaguliers [238]. described [178]. desert [551]. deservingness [761]. design [864, 6, 21, 51, 55, 513, 271, 353, 869, 244]. Designed [686, 650]. designing [186]. desire [688]. destabilizing [61]. destruction [616]. Detachment [616]. determinism [840]. Deutsches [225]. deuxième [166]. develop [581]. developed} [825]. developing [535]. Development [490, 819, 489, 470, 452, 403, 87, 113, 6, 370, 513, 270, 89, 274, 799, 145, 121, 275, 811, 800, 364, 244, 35, 78, 10, 415, 820, 176, 313, 239]. developmental [707]. developments [371]. développement [239]. devices [44]. d’explication [95]. dialectic [86]. dialogue [96, 98]. did [209, 529]. Didimac [855]. didn’t [401]. diet [590]. difference [414]. Different [498, 654, 806]. diffusion [527]. Digital [334]. dikes [448]. dimensions [272]. Diogenes [538]. diplomacy [835, 832, 829, 833, 831]. diplomats [831]. directions [329]. disability [809]. Disaster [857, 409, 873]. discipline [572]. Disciplining [378]. discontinuity [102]. discourse [673, 408]. Discourses [551]. discovery [376, 377]. diseases [773]. disembodied [84]. disgust [774]. disinfestation [258]. dislocare} [768]. Disney [683]. Disneyland [700]. Displaying [419]. displays [674]. Dissertations [487, 465]. distinguished [740]. district [852]. divine [815]. division [511]. DNA [803]. Do [473]. doctrinal [80]. document [7, 12]. documentation [830]. Does [473, 479]. doing [810]. domain [267]. Domestic [124, 821, 788]. Domingo [92]. donating [675]. donations [675]. Dorado [800]. Doriden [656]. d’ouvrage [168]. Dove [593]. Dover [730]. down [843]. dramaturgy [536]. drap [135]. draw [619]. drawing [647]. dream [343]. dreams [699]. drive [585]. Drivers [582, 583]. Driving [584, 849]. drug [653, 658]. drugs [655, 516, 517]. Dual [477]. Dual-Use [477]. Duhem [77]. d’un [42]. d’une [62, 95, 57]. durant [11]. Durham [338]. during [737, 732, 370, 195, 249, 849, 877, 739, 614, 679, 540, 605, 153, 35, 707]. Dutch [512, 359, 284, 697]. Dutch-controlled [697]. Dye [479, 56]. dynamics [527, 605].

Early [211, 63, 539, 599, 785, 356, 655, 512, 864, 627, 320, 517, 767, 141, 736, 750, 257, 340, 747, 757, 540, 215, 733, 92, 795]. early-modern [512, 767]. Earth [843, 639, 168, 377]. East [714, 365, 858]. Eastern [650]. échec [134]. Echoes [438]. eclectic [351]. école [314, 13]. ecomuseums [229]. Economic [608, 719, 820]. economics [234, 788]. economy [251, 520, 525, 651, 790, 299]. ecosystem [551]. Edens [63]. édiles [125]. Editor [553, 765, 632, 660, 541, 548, 533, 555, 573, 578, 603, 612, 618, 640, 649, 667, 672, 689, 696, 701, 724]. Editorial [14, 25, 32, 40, 48, 53, 59, 67, 74, 112, 119, 130, 157, 163, 164, 175, 182, 193, 204, 287, 293, 298, 307, 317, 323, 328, 350, 355, 369, 362, 368, 375, 381, 387, 393, 587, 617, 695, 723, 744, 120, 745, 874, 184]. Editors [779, 801, 828]. Education [475, 309, 763, 270, 269, 267, 524, 310, 121, 721, 228, 35, 222]. educational [877]. Efficiency [491, 788]. Effort [484]. efforts [581]. egg [786]. église [97]. eighteenth [177, 166, 614, 679, 167, 682]. eighteenth-century [177, 682]. Eighty [662]. Einthoven [490]. either [263]. Eldo [211]. electric [863, 413, 704]. Electrical [294, 17, 873, 69, 727, 70]. electricity [630, 606, 873]. electrification [671]. électroaimant [29]. electrochemical [536]. electromagnet [29]. electromagnetic [28]. Electronic [370, 378, 248, 763, 244]. Electronics [255, 251, 256]. électronique [248]. elementi [90]. Eliminating [852]. Elitism [499]. embed [413]. embodied [626]. embracing [877]. embryos [805]. emergence [283, 87, 280, 620, 247, 781, 793]. Emerging [692, 698]. emotions [778]. empire [512, 711, 795, 856]. Empires [794, 796]. empiricism [89]. empirisme [105]. encounters [638, 756]. Encouragement [124]. Encyclopédie [647]. end [216]. Endeavour [290]. ends [586]. enemy [726]. Energizing [651]. Energy [887, 871, 833, 840, 403, 813, 850, 535, 702, 215, 262, 780, 629]. Engel [285]. Engine [372, 536, 11]. Engineer [501, 343, 727, 238]. Engineering [596, 186, 271, 475, 748, 549, 244, 457, 470, 818, 138, 309, 740, 510, 69, 308, 692, 453, 273, 749, 50, 195, 187, 860, 267, 272, 310, 34, 873, 275, 454, 70, 191]. Engineers [499, 497, 498, 495, 502, 273, 125, 454, 496, 650, 314, 339, 311, 270, 271, 272, 312, 295, 261, 325, 614, 346, 428]. engines [675]. England [537, 140, 28]. engravings [337]. enjeux [202]. enlightened [511]. Enlightenment [684, 647]. enough [26]. enquiry [338]. Enrichment [798]. ensemble [6]. Entangled [685]. entanglement [814]. Enterprise [401, 62]. enters [398]. entrepreneurial [614]. entrepreneurs [748, 746, 261]. entrepreneurship [750]. entreprise [62, 124]. entry [310]. Environing [776]. environment [759, 559, 290, 776]. Environmental [469]. environmentalism [786]. environments [757]. envisioning [636]. Enzyme [537]. Epistemic [878]. Epistemologic [159]. epistemological [810]. Epoque} [506, 615]. EPRI [704]. era [546, 883, 775]. Eric [867]. Eroticism [474]. Errata [39]. Erratum [73, 143, 183, 279, 648, 715]. ESA [345].

Escaped [490]. espingarda} [860]. esprit [171]. ESRF [762]. ESRO [345, 206, 207]. Essay [286]. Essence [29]. essor [78]. Est [109]. Est-ce [109]. Establishing [762]. estate [858]. Esthétique [97]. Ethernet [245]. ethnography [761]. Europe [458, 356, 209, 762, 208, 609, 527, 620, 607, 257, 455, 459, 511, 604, 605, 589, 10, 457, 780, 650, 591, 456]. European [676, 762, 211, 585, 27, 527, 608, 620, 606, 11, 376, 352, 260, 151]. Europeanism [604]. européenne [11]. évêque [101]. Every [468]. everyday [283]. Everyman} [792]. everyone [698]. evidence [813]. Evolution [199, 87, 185, 131, 62, 830, 262, 166, 62, 166]. evolutionary [318]. examining [606]. Example [500, 108]. examples [212]. exchange [273, 532]. exchanges [728, 444]. exclusion [309]. exemple [108]. Exhibiting [420]. exhibition [17]. exhibits [227, 226]. exodus [621]. expanding [796]. expectations [651]. expensive [710]. experience [520, 606, 415]. Experiment [28]. experimental [803, 627, 208, 628, 798]. experiments [872]. expert [657]. Expertise [506, 785, 820]. experts [800, 400]. Explaining [309, 318]. explanations [95, 158]. explanatory [110]. explicative [110]. exploration [796, 636]. explorers [377]. explosion [338]. Expo [838]. exposed [838]. Extraterrestrial [638]. Eye [567, 627, 639, 571]. eyes [370, 570].

F [796, 678]. fabric [446]. fabrication [236, 545]. faces [104]. facialization [718]. Facilitating [728]. Facility [762]. factor [61]. factories [808]. Factory [435, 782]. Fad [437]. failed [320, 703]. failure [134, 324]. Fairchild [353]. fall [206, 580]. familial [135]. family [135]. famines [708]. Faraday [338]. Farben [482, 265, 481]. farming [855]. Fashioning [631]. Fatal [560]. fate [7, 12]. fathers [311]. Fatigue [872]. FDA [406]. fear [732]. fears [832]. Federal [403, 382]. Female [315, 686, 717]. Fer [170]. Fermi [395]. Ferry [857]. Fez [630]. fibre [114]. fibre-optic [114]. fiction [513]. field [740, 831]. files [830]. Filière [116]. Film [474, 602, 482]. films [400]. findings [131]. fingers [865]. Finland [346]. Firma [161]. firms [653]. First [497, 334, 190, 403, 861, 675, 424, 538, 206, 727, 834, 299, 263, 728, 735, 737, 233, 790, 738, 258]. fit [528]. fix [754]. Fixing [761]. Flapping [277]. fleet [401]. flight [277, 263, 595]. Flows [836, 547]. fly [401]. Flying [539, 566, 639]. focus [693]. Following [314]. fonction [42]. fondations [168]. Fonds [43]. Food [491, 506, 588, 589]. foods [464, 591]. forces [140]. Ford [461, 226]. forecasts [295]. Foreign [270]. Foreword [194, 75, 266, 161]. forgotten [606, 846]. Form [513, 659, 333, 21, 752]. Forman [532]. Formation [70, 28, 125, 125]. forms [277]. Forum [410]. fossil [852]. Foucault [652]. foundation [233, 234]. Foundations [168, 810]. Four [10]. fourteenth [18]. Fowl [839]. Fragile [607]. framework [540]. francaise [100]. France [177, 196, 294, 655, 69, 202, 518, 140, 506, 398, 88, 269, 141, 848, 360, 507, 790, 615, 95, 282, 170, 439, 11, 295, 614, 738, 854, 78]. Franconia [364]. Frankenstein [408]. Franz [748]. free [62]. Freiberg [153]. French [83, 630, 599, 138, 196, 101, 96, 497, 201, 198, 248, 252, 64, 314, 202, 93, 82, 29, 339, 77, 36, 94, 140, 135, 160, 107, 127, 7, 122, 4, 134, 100, 575, 31, 110, 236, 62, 271, 50, 125, 168, 165, 195, 104, 358, 304, 280, 274, 166, 95, 108, 170, 37, 109, 254, 11, 136, 299, 496, 2, 124, 139, 325, 85, 682, 78, 239, 97, 42, 13, 57, 152, 12]. French-speaking [271]. friendship [123]. Front [839, 313]. Front-line [839]. frontier [690, 2, 678]. frontière [2]. frozen [607]. fuel [651, 833, 581, 846, 216]. fuel-cycle [216]. fuels [581, 852]. function [513, 659, 42]. fury [736]. fusion [835, 579]. future [677, 424, 712, 419, 581, 636, 462, 635]. Futures [706, 873]. futurity [661].

G [482, 481]. GAF [481]. Gaining [308]. Galérie [232]. Galilée [93]. Galileo [93, 93, 81, 98]. gallery [554]. Galvanometer [490]. gap [345]. garden [760]. gas [201, 849, 756, 415, 258]. Gasoline [29]. gauge [132]. gay [595]. GDR [313]. Gehry [522]. Gender [396, 552, 686, 389, 511]. Genealogy [505]. General [482, 413, 461]. generation [229]. genes [805]. Genesis [302, 69, 538]. genius [23]. Geographies [614]. geography [255, 720]. Geological [115]. geology [430]. geophysics [811]. geopolitics [714]. Geopostal [198]. géopostale [198]. George [434, 796, 390]. geostrategic [726]. geothermal [852]. Gerhard [616]. German [813, 265, 285, 787, 365, 85, 670, 278, 756, 733, 485, 697]. German- [697]. Germany [421, 655, 634, 403, 311, 517, 848, 169, 274, 295, 738, 153, 223, 792]. gets [619]. Getting [536]. ghost [808]. Gilbert [751]. give [624]. Glaxo [291]. Global [508, 467, 564, 759, 662, 408, 538, 621, 563, 720, 733, 534, 843]. Globalization [408, 394, 579]. GmbH [384]. goals [228]. Goddard [437]. Goddess [867]. good [238]. governance [607, 605, 752]. Governing [677]. governmentality [562]. Governors [833]. Gramophone [468]. Grand [438]. grandes [314]. Grave [468]. gravium [93]. great [781, 18, 60, 81, 550, 151]. Greece [449]. Greek [499]. Green [535, 821]. grounds [134]. group [96]. groupe [96]. Grumman [461]. Guangzhou [752]. Guayana [800]. Guerilla [435]. guerre [248]. Guerrilla [713]. guests [315]. guinea [340]. Gulf [551]. Guse [558].

H [423]. H-bomb [423]. H. [590]. Habitability [843]. Hahn [884]. hairdressers [552]. Half [497, 166, 11, 753, 299, 614, 679]. Han [842]. hands [865]. happen [209]. happened [450]. Harbingers [347]. Hardcastle [777]. hardware [579]. harmony [326]. harvesting [414]. Harvey [510]. Hassenfratz [236]. Haswell [338]. Hawai‘i [795, 820]. HBO [666]. Headlong [661]. Health [507, 518, 406, 407, 878, 824]. healthcare [713]. heap [694]. hearing [791]. heart [759]. heat [851]. heating [852, 851]. hegemony [676]. Heinz [590]. Henry [226]. here [390]. Herf [557]. Heritage [509, 92]. Hermétisme [105]. heroic [730]. hidden [455]. hier [77]. High [542, 55, 215, 262]. high-energy [215, 262]. higher [269]. highway [200, 197]. highways [125]. Hiroshima [422]. Histoire [202, 2, 152, 4, 31, 78, 57]. Historia [22]. historian [100]. Historians [487, 483, 465]. historic [630]. Historical [504, 150, 54, 509, 248, 332, 179, 64, 687, 269, 877, 221, 693, 159]. Historicizing [845, 810]. historien [100]. histories [613, 389, 324, 808, 773, 768]. Historiographical [341, 280, 802]. Historiography [475, 434, 351]. historique [248, 64]. History [329, 567, 582, 492, 553, 836, 1, 242, 333, 789, 885, 816, 867, 761, 202, 668, 200, 84, 3, 4, 398, 31, 525, 759, 725, 115, 187, 702, 276, 602, 857, 860, 409, 703, 20, 382, 727, 102, 272, 846, 340, 247, 729, 80, 621, 41, 454, 635, 68, 563, 645, 858, 694, 191, 78, 840, 184, 688, 683, 223, 880, 886, 663, 514, 161, 2, 152, 8]. hole [759]. Holl [522]. Holocaust [258]. Holography [434, 471, 189]. home [326, 674, 788, 338]. honest [538]. horse [773]. horse-cloth [773]. hospital [712]. Hosting [609]. House [678]. household [400, 36]. housewife [400]. Housewive [400]. Hughesian [276]. Human [340, 510, 865, 277, 597]. humanism [395]. Humanist [97]. humaniste [97]. humanities [840]. humanoid [778]. humble [841, 881]. Hungary [824]. hydraulic [510, 42]. hydraulique [42]. hydroelectric [861]. hydrogen [651]. hygiene [326]. Hypothesis [440, 233, 234]. hyuji [862].

I. [482, 187, 481]. IAEA [826, 830, 825, 834]. Iberian [753]. IBM [545]. ice [530]. Iceland [852]. Icon [479]. iconic [754]. Iconography [91]. ICT [444]. ideal [233, 234]. Ideals [496]. ideas [84]. Identities [495, 566, 607]. Identity [499, 501, 453, 594, 620, 390, 595]. Ideology [492, 20, 820]. if [281]. IG [265]. II [251, 132, 520, 370, 249, 481, 741, 485]. Illuminating [568, 569, 812]. illustrations [754, 783]. Image [473, 553]. Images [553, 433, 734, 601]. imaginaries [714, 721]. imaginary [716]. Imagined [706, 729]. imagining [731]. Imitation [277]. Impact [437, 243, 261]. Imperial [634, 864]. implantation [544]. Importance [96, 191, 654, 727, 96]. Imported [760, 859]. importers [278]. Impossible [596]. improvement [68, 681]. in-the-making [518]. inch [538]. Inciteful [770]. Increase [491]. index [161]. India [161, 709, 717, 718, 721, 563, 719, 827]. Indian [856]. Indians [772]. indigenous [577]. individualization [524]. Indonesia [817]. Industrial [439, 485, 818, 888, 229, 252, 870, 339, 446, 6, 16, 134, 27, 517, 208, 50, 711, 551, 538, 402, 734, 872, 590, 511, 866, 35, 10, 681, 145]. industrialising [346]. industrialism [688]. industrialization [414, 153, 537]. industrializing [866]. industrie [147, 36, 11, 124]. industriel [134]. industries [60, 365]. Industry [485, 433, 470, 289, 36, 288, 844, 284, 406, 703, 592, 268, 353, 253, 254, 11, 811, 261, 301, 670, 547, 521, 881, 697, 124]. Influence [273, 270, 274]. influencing [253]. informatic [329]. informatics [31]. information [242, 398, 525, 527, 629]. informatique [31]. Infrastructural [604]. infrastructure [887, 871, 823, 330, 662, 611, 804, 842, 841]. infrastructures [604, 445]. Inga [772]. Ingénieur [237]. Ingénieurs [125]. ingenios [628]. ingredients [526]. initiative [843, 17]. Innovation [282, 489, 291, 151, 319, 816, 17, 351, 27, 146, 288, 358, 859, 703, 255, 363, 713, 318, 547, 290]. innovations [320, 511]. INRIA [609]. inscriptions [566]. insecticide [855]. insects [795]. insemination [868]. inseparability [149]. instigator [197]. Institut [302]. Institute [704, 261]. institution [62]. institutional [208, 360, 540]. Institutions [356, 440, 740, 69]. Instrument [490, 283, 682]. Instruments [783, 294, 116, 132, 566]. insulin [519, 806]. insurers [831]. integrated [30]. integrating [620, 690]. integration [455, 443]. Intellectual [777, 623, 382]. intelligence [726, 113, 598]. Inter [499, 436, 289, 791, 718, 797]. Inter-artifactual [436]. inter-colonial [797]. Inter-war [499, 289, 791, 718]. Interactions [364]. intercontinentales [199]. intercultural [273]. Interdisciplinarity [469, 574]. Interest [487, 465, 519]. interests [702]. International [8, 371, 243, 824, 833, 620]. internationalism [530]. InterNyet [529]. Interpreting [483, 601]. intersected [752]. interspecies [878]. interventions [819]. Introducing [553]. Introduction [329, 814, 449, 356, 248, 504, 495, 567, 582, 560, 596, 622, 642, 241, 417, 455, 44, 802, 875, 766, 404, 442, 289, 512, 165, 794, 679, 604, 588, 248]. introuvable [107]. inventeurs [239]. Inventing [702, 455, 628, 500]. Invention [357, 114, 360, 249, 363, 743, 805, 540, 738, 526, 158, 57]. inventors [239]. inventory [147]. investigations [189, 857, 409]. investment [365]. ion [544]. ionogram [378]. ionosphere [295, 296]. Iowa [390]. Iridium [467]. Iron [141, 170]. irradiated [464, 591]. Irreplicability [877]. Islam [817]. issue [802, 9]. issues [202, 208, 845, 145, 218]. Italie [82]. Italy [498, 82, 167]. ITER [835, 579]. Ivory [749].

J [590]. J. [238]. Jack [818]. Japan [865, 270, 208, 832, 833, 859, 419]. Japanese [778, 213]. Jeffrey [557]. Jésuites [93]. Jesuits [93, 92]. Jet [489, 372, 675]. Jewish [816]. John [678, 558, 532]. Joint [842]. Joliot [304, 303]. Joliot-Curies [304, 303]. Joseph [161]. jour [171]. Journal [8]. journey [759]. Juanelo [44]. Jubilee [480]. Jules [662]. Julien [57]. junction [459]. junctions [444]. jusqu’en [42].

Keiser [818]. Kennedy [678]. Kepler [102]. Khankhoje [821]. kind [806]. KLM [161]. Know [473, 268]. know-how [268]. knowing [576, 770]. Knowledge [453, 706, 836, 786, 644, 84, 626, 186, 794, 647, 799, 804, 747, 34, 231, 451, 787, 774, 809, 625, 858, 798, 688]. koji [804]. Konfederatov [187]. Korea [760, 863, 857, 860, 862, 842]. Korean [435, 692, 823, 824]. Koyré [101, 86, 82, 77, 107, 84, 106, 89, 104, 85, 78, 79, 81, 80]. Krige [532]. kyongun’gi} [760].

L. [65]. labor [872, 511]. laboratories [295, 331]. Laboratory [490, 439, 433, 413, 787, 395, 390, 300, 758, 395]. labors [887, 871]. labour [256, 865]. Lachine [42]. Lag [440]. Lake [797]. lamps [630]. Landings [483]. landscape [693]. language [770]. lantern [663]. lantern-bearers [663]. Lanz [128]. Large [206]. Largo [768]. last [198]. late [750, 357, 747, 395]. later [89]. lathe [113]. Latin [49, 815]. launcher [211]. Lavoisier [236]. Law [626, 518, 625]. layers [341]. ldots [202]. lead [284]. Leblanc [337]. lecturers [750]. Led [489]. Legal [440]. legibility [671]. LEGO [523]. length [538]. Lenormand [402]. lens [666, 577]. Leo [676]. Leo-penicillin [676]. Leonardo [90]. less [825]. Lessons [409, 846]. Let [535]. letters [732]. liability [831]. liberalization [606]. liberated [405]. liberty [572]. libre [62]. life [559, 808, 95, 565, 590]. Light [567, 535]. Lightening [762]. ligne [196]. Like [311, 782]. lime [50]. l’Industrie} [439]. line [337, 196, 839, 185]. Linearity [489]. link [728]. links [607]. Linnaeus [233, 234]. Listening [735, 737, 739]. live [757]. lives [876]. lobbyists [831]. Local [498, 869]. localization [863]. Locating [577]. location [256]. locomotive [6, 21, 51, 55, 136]. logistic [176]. logistics [197]. London [631, 479, 526]. Long [49, 509, 677, 846, 679, 605, 428]. Look [467]. loss [791]. lost [107]. Louis [99, 402]. Louis-Sébastien [402]. Louvre [13]. loves [887, 871]. Low [502]. Lyell [338].

Macadam [436]. macchinali [90]. machine [6, 90, 547, 528]. machine-ensemble [6]. machinery [463, 239]. Machines [466, 232, 337, 760, 44, 11, 239, 33]. Madagascar [841]. madder [264]. made [615]. Magazine [680]. Mahal [771]. main [214]. mainframes [545]. major [206]. make [209]. maker [810]. makers [294, 628]. makes [599, 414]. Making [803, 855, 467, 674, 859, 862, 210, 363, 858, 680, 396, 539, 888, 870, 518, 644, 882, 213, 810, 712, 544, 214, 787, 774, 786, 776, 457, 319, 707, 880, 462]. makings [861]. male [312, 595]. man [425, 90, 538, 405, 882]. man-machine [90]. Management [435, 345, 363, 788, 682, 795]. managing [708]. manly [315]. manufacture [864, 236, 694]. manufacturers [254]. manufactures [264]. Manufacturing [865, 545, 60, 546, 542, 544, 511]. map [619]. mapping [823, 614]. Marc [140, 139]. Marconi [730]. marginal [881]. Maria [602]. Marie [301]. maritime [682]. mark [884]. Market [506, 34]. marketing [655, 653, 658, 654, 656, 657]. marketplace [537]. Markets [509, 655, 256, 653, 658, 353, 547]. Marly [463]. Marshallese [711]. martin} [682]. Marx [146, 233, 234, 15]. masculinity [730]. mask [756]. masks [104]. masques [104]. Mass [502, 421, 281, 353, 524, 721, 485]. mass-individualization [524]. mastery [781]. Masud [714]. material [761, 694, 443]. materialities [769]. Materiality [486, 659, 851]. Materializing [752]. Materials [169, 141, 165, 611, 172, 859, 166, 167, 547, 537]. matériaux [165, 166]. mathématiques [103]. Matthew [560]. Maurice [19]. Mauve [479, 480]. May [291]. McOndo [645]. me [171]. meal [464]. Meaning [357, 326, 673, 687]. means [835]. measurement [525, 108]. Mécanique [168]. mechanical [128]. Mechanics [684, 6, 21, 51, 55, 168, 274, 238]. mechanization [760]. Mechanizing [389]. Media [433, 705, 615]. Mediation [267, 459]. medicine [797]. medieval [102, 257, 357]. meeting [8, 449, 730]. meetings [833]. meets [885, 408]. Meiji [859]. Mekong [799]. Meldola [480]. memoriam [19]. memories [333]. men [566]. Ménage [36]. mentality [177]. Mersenne [96, 81]. Messenger [839]. messy [226]. mesure [108]. metallurgists [888, 870]. Metallurgy [22]. Methanol [581]. method [64, 860, 528]. méthode [64]. Methodological [280, 158, 159]. methodology [877, 41]. methods [21]. métiers [121]. Metrics [502, 453]. metrology [300]. Mexicanos} [501]. Mexico [551, 712, 501, 710, 825, 239]. Mexique [239]. Meyerson [86]. mice [805]. Michael [338]. Michelin [197]. microbes [546]. Microelectronics [486]. microhistories [768]. mid [524, 407, 383, 167]. mid-1960 [407]. mid-eighteenth [167]. mid-nineteenth [383, 167]. mid-twentieth [524]. middle [121, 18]. Midi [596]. migratory [549]. Military [466, 257, 217, 711, 378, 295, 151, 457, 296]. milk [445]. Milky [445]. Millennial [330]. millet [887, 871]. mineral [796]. Mines [491]. Mining [153, 772, 268, 881]. Ministère [239]. Ministry [253, 239]. Minitel [398]. Misconception [560]. missile [580]. Missing [884, 611]. mission [818, 834]. missions [826]. mobilization [738]. mobilized [792]. Mobilizing [836]. MOD [253, 631]. mode [654]. Model [267, 539]. Models [231, 864, 883, 277, 318, 33]. Modern [630, 539, 599, 785, 400, 512, 864, 87, 320, 767, 712, 275, 540, 56, 356]. modernisation [284]. Modernism [499]. Modernity [498, 492, 539, 772]. Modernization [439, 400, 848, 822, 534]. Modernizing [844]. Moigno [139]. moitié [166, 11]. molecular [751]. monopoly [540]. monoxide [882]. monuments [630]. Moon [483]. Moorkultur} [781]. moped [631]. moral [520, 299]. Morocco [630]. mortality [880]. Moscow [188]. mosque [630]. Mosquito [425, 405]. most [463]. Mother [473]. motion [339]. Motors [461]. motu [93]. mouvance [94]. move [353]. movement [641, 94, 810, 872, 847]. Mozhaiskii [263]. Mrs [678]. Mukerji [596]. multimedia [763]. Multiple [861, 805]. Mumbai [881]. murals [712]. Museet [225]. Museum [479, 224, 674, 225, 228]. Museums [687, 221, 229, 223, 222]. Mutant [822]. Muybridge [722]. mystique [105]. myth [727].

n [110]. Nîmes [177]. naissance [29, 136]. name [785]. naming [770]. nanoelectronics [692]. nanotech [462]. Napoleon [99]. Napoleonic [88]. narratives [719]. NASA [843, 382]. Nation [498, 706, 124, 500, 844, 712, 755, 720, 124]. Nation-Building [706]. National [499, 495, 758, 395, 228, 484, 655, 777, 826, 121, 873, 295, 811, 262]. nationale [147, 124]. nationalism [675, 709, 419, 755, 718, 782]. nationalist [576]. Nations [535]. nationwide [529]. natural [93, 233, 234, 191, 881]. naturalis [22]. Nature [473, 452, 277, 549, 682, 448, 442, 447]. naturelle [93]. Navigating [711]. Navigation [608, 755]. Navy [363, 430]. Nazi [556, 285, 621, 570]. NCI [406]. need [825]. Needham [161]. négoce [135]. negotiating [592]. negotiation [267]. negotiations [260]. neoliberalism [719]. Netherlands [677, 452, 320, 172, 436, 318, 319, 443]. Network [195, 196, 201, 446, 529, 733, 152]. networked [812]. Networking [448, 456, 442]. Networks [575, 289, 202, 607, 303, 793]. never [252, 430]. Nicolas [337]. nightmare [683]. Nightmares [700]. nineteenth [673, 273, 100, 269, 274, 750, 747, 746, 11, 753, 325, 383, 167, 318, 605, 223]. nineteenth- [746]. nineteenth-century [673, 325]. Nixon [580]. nodes [200]. noise [620]. Non [477, 489, 700, 812]. non-actor [700]. Non-Linearity [489]. non-networked [812]. Nordic [484]. North [435, 281]. northern [848]. northwest [841]. Norway [251, 296]. Norwegian [289, 592]. nostalgia [702]. Note [632, 618, 640, 649, 667, 672, 689, 701, 660, 541, 548, 533, 555, 573, 578, 603, 612, 696, 724, 553, 700, 765, 779, 801, 828]. notebook [787]. Notes [306, 322, 327, 336, 349, 354, 361, 367, 374, 380, 386, 392, 411, 427, 432, 158]. nouveau [7, 12]. nouveaux [125]. Nuclear [403, 813, 699, 674, 826, 702, 832, 304, 829, 833, 830, 594, 424, 831, 544, 709, 259, 670, 854, 780, 415, 216]. nuclearity [408]. number [110]. numbers [813, 331]. numerous [315]. Nyegaard [290].

Oak [420]. oasis [551]. objections [110]. objects [687]. observation [566, 872]. Observations [159, 131, 184, 17]. observers [566]. observing [843]. Occulted [769]. Occupation [313]. ocean [399, 383]. ocean-scape [383]. oceans [850]. Oeconomia} [233]. oeconomic [682, 681, 680]. oeconomy [679, 682]. Office [122, 389, 518, 742, 338]. officers [313]. offices [527]. oil [852, 811, 365, 581]. oiseau [682]. Olympic [225]. One [467, 461, 712, 862, 879]. ontologies [761]. open [698]. operations [339]. opportunities [692, 802]. Oppositions [110]. optic [114]. optics [102]. oral [340]. Oratoire [94]. Oratory [94]. order [644, 517, 817]. ordinary [876]. organic [278]. Organisation [528, 256]. organization [265, 788, 149]. organizing [883]. Origin [42]. Original [211, 479]. Origine [42]. Origins [179, 190, 115, 543, 821, 414, 9, 827]. Orthodox [816]. other [785]. otherwise [880]. Otter [567]. our [710, 462]. overproduction [847]. overview [407]. overviews [514]. own [862]. ozone [759].

Pacific [376]. Packaging [247]. pages [161]. painting [593]. paintings [554]. Pakistan [826, 709, 714]. Pandemic [879]. Pandurang [821]. Panic [732]. pape [97]. paper [830, 862, 319]. Papert [810]. parable [211]. parables [702, 754]. Paracelse [104]. Paracelsus [104]. Parallels [276]. Paris [17, 122, 126, 304, 123, 610, 232, 302]. Parish [680]. parks [812]. parmi [104]. part [233, 12, 116, 132]. participation [677]. particle [214]. partie [12]. parts [234]. party [831]. past [161]. Pasteur [302]. patent [337, 518, 518]. patentable [516]. patentees [359, 746]. patenting [750, 520, 519]. Patents [518, 359, 521, 519, 382, 540, 517]. patina [334]. Patronage [98]. Patrons [221]. patterns [99]. Paulo [775]. pavements [188]. Pavillon [302]. pays [62]. peace [591, 710]. peaceful [565, 838]. peatlands [781]. pendulum [131]. Penicillin [676, 520]. peninsula [798]. people [536, 680]. Pererius [93]. performance [253]. period [320]. périodisation [103]. peripheral [755]. Perkin [479]. Perrache [134]. personal [283, 79, 80]. personnel [854]. Perspective [504, 509, 248, 179, 308, 585, 27, 54, 786]. perspectives [454]. Pesticide [507]. petroleum [430]. Pharmaceutical [653, 289, 288, 517, 659, 656, 406, 521, 290, 657, 518, 546, 657]. pharmaceuticals [516]. pharmacy [56]. Phase [176, 153]. Phase-alternating [176]. phases [10]. phenomenologists [85]. phénoménologues [85]. phenomenology [756]. Philadelphia [761]. Philippines [796]. Philips [413, 446, 250]. Philomathique [123]. philosophie [83, 93]. philosophy [93, 628, 83]. phonography [593]. photograph [601]. photographic [627]. photographing [389]. photography [728, 639, 747, 722]. physical [140, 331]. Physicists [295]. physics [91, 421, 674, 749, 750, 544, 215, 262, 139]. physiological [872]. physiology [510]. Physique [139]. physiques [140]. piano [787]. piano-making [787]. Picardy [23]. picture [611, 315]. piece [463]. Pigeons [839]. pigs [340]. pill [396]. pills [656]. pioneers [314]. pipeline [743]. Pistes [95]. pits [338]. Place [574, 879]. Plains [550]. Plan [707]. planes [461]. plantation [798]. Plantationocene [775]. planting [550]. Plants [795]. plate [722]. plates [647]. play [523]. Playing [523]. Plinius [22]. plutonium [832]. pneumatics [197]. Point [196, 196]. Poland [166]. policies [462]. Policing [682]. Policy [433, 677, 452, 252, 280, 213, 382, 579, 540, 262]. Political [567, 440, 251, 644, 398, 208, 857, 790, 299, 451, 223]. Politics [836, 198, 127, 675, 88, 576, 579, 303, 260, 296]. politique [198, 127]. Pologne [166]. polyglot’ [867]. pont [135, 227]. Ponts [125]. Pope [97]. popular [714, 227]. popularization [402]. population [562, 560]. porcelain [864]. Pornography [473]. porter [526]. Portraying [554]. ports [823]. Portugal [500]. Portuguese [860]. position [11]. possibility [329]. possible [95]. Post [706, 818, 251, 248, 676, 823, 865, 255, 883, 636, 771, 260, 244, 577, 798, 742, 532]. Post-1945 [706]. post-Apollo [260]. post-colonial [883, 577]. post-Second [244]. post-war [818, 248, 676, 823, 865, 255, 636, 798]. post-World [251]. postcolony [574]. Poster [671]. postérité [110]. posterity [110]. Postmodernity [492, 469]. postwar [403, 674, 691, 419]. potential [337]. Pour [31, 147, 122, 100, 236, 124]. Power [144, 497, 704, 624, 760, 536, 561, 575, 861, 671, 848, 879]. powered [848]. practical [647]. practicality [272]. Practice [475, 433, 813, 342, 284, 20, 238, 390, 412]. practices [797]. Practicing [679]. pre [511, 595]. pre-industrial [511]. pre-Second [595]. Preface [76, 205, 219, 478, 97, 97]. preliminary [301, 147, 834]. preparations [347]. Preparing [704]. presence [325]. Presencing [815]. present [8, 771, 161]. pressure [55]. prestressed [171]. Pretzfeld [364]. prevention [406]. Primacy [492]. Primitive [358]. princely [856]. principles [159]. printing [158]. Prison [798]. Prisoners [491]. Privas [43]. private [17]. Probabilistic [813]. problem [785, 885, 686, 16, 626, 128, 872]. Problems [144, 187, 280, 454]. process [654, 268, 30]. Processes [470, 214]. Processing [880]. procurement [545]. producer [820]. product [513, 247]. Production [116, 485, 179, 446, 637, 516, 829, 353, 862, 869, 549, 858, 589]. Productivity [491, 68]. products [855, 384, 589]. Professional [289, 497, 517, 453, 527]. Professionalisation [100]. professionalization [311, 268]. professionals [623, 429]. professionnalisation [100]. professors [749, 748]. Profiling [718, 399]. profiloscope [718]. program [551, 212]. Programmatic [110]. programmatiques [110]. programme [259, 260]. programs [210, 825]. Progress [502, 471, 453, 61, 54, 717]. Project [484, 134, 206, 604, 168, 483, 842, 798]. Projects [693, 207]. projet [134]. Proliferation [477, 832]. promise [800]. promises [639]. promoting [763]. propaganda [377, 658]. property [777, 623, 382, 625]. propos [64]. Propulsion [489]. protest [428]. Proto [882]. provincial [325]. provocation [450]. Prussian [115]. Pseudo [44]. psychic [731]. psychologist [743]. Psychology [866, 627]. Public [438, 507, 780, 677, 452, 673, 518, 519, 227, 325, 670]. publication [98]. Publications [24, 46, 71, 156, 181]. publics [221]. published [191]. puits [196]. pulp [703]. pumps [851]. Punched [246, 527]. Punched-card [246]. Purity [590]. Pursuing [740]. Putting [850]. Pyramid [30].

quantification [872]. Québec [271]. Query [110]. quest [536, 638]. question [655]. Questioning [68]. questions [685]. qu’il [109]. quinine [697].

R&D [256, 365]. race [718, 820]. Racial [390]. radar [370]. Radiation [762, 591, 377]. radio [347, 844, 243, 792]. radio-telephone [792]. Radioactivity [300, 508]. radiocommunications [295]. Radium [302, 301]. railway [605]. railways [753]. rain [704]. Rana [714]. Raphael [480]. Rationalization [499]. Rationalizing [435]. raw [859]. rays [698]. RCA [844]. Re [324]. Re-writing [324]. Reactionary [499]. reactor [403, 833]. Reactor-3 [833]. Reactors [650]. Read [585]. realities [699]. Really} [473]. realm [876]. reappraisal [288]. Reassembling [842]. Reassessment [479, 257]. received [24, 46, 71, 156, 181]. recenter [768]. reception [267]. recherche [2]. reciprocity [835]. reclamation [140, 140]. Reconfiguring [823]. reconstruction [873]. records [399]. recovery [61]. rectifier [364]. rectifiers [249]. Rectors [680]. Redefining [589]. Reefs [551]. reference [88]. refinement [365]. refining [725]. Reflections [3, 259, 330, 324, 80, 802, 4, 408]. Réflexions [4]. Reform [435, 97]. réforme [97]. réformés [95]. Reforms [95]. Reframed [802]. regenerated [873]. Regime [477]. region [800]. regional [883]. Regression [474]. Regulation [508, 506, 507]. Regulations [509]. Rehabilitation [809]. Reich [491]. reinforced [171]. reinvention [758]. relation [670]. relations [227, 415]. relationship [341]. reliability [384, 378]. reliable [384]. Religion [282, 814, 885, 358, 815]. religious [590]. remap [768]. Remember [816]. reminiscences [79]. remote [210]. Renaissance [18, 89]. renewables [845]. repair [761]. repercussions [738]. replace [313]. replies [558]. Reply [65]. Report [423]. repression [285]. reprocessing [832]. Reproducing [561]. Republic [403, 359, 582, 583]. Republican [752]. reputation [747]. Research [249, 704, 833, 836, 265, 1, 289, 740, 413, 692, 546, 250, 16, 304, 280, 877, 883, 741, 2, 207, 331, 364, 797, 251]. réseau [196, 152]. réseaux [202]. reservoirs [878]. Resetting [722]. reshaping [447]. resilience [713]. resource [881]. resources [740, 710]. response [60, 564, 676, 626, 646, 600, 572, 884, 586]. Responses [493]. restauration [127]. resting [862]. restoration [127, 195, 126]. Rethinking [633, 475, 410]. retrospective [526]. Reuleaux [748]. Reverse [440, 860]. Review [515, 161, 154, 867, 286, 231, 531]. Reviews [373, 426, 431, 441, 460, 472, 488, 503, 45, 494, 321, 335, 348, 379, 391, 416, 292, 297, 117, 316, 305, 385, 344, 366]. revisited [556]. revisiting [527]. revolt [847]. Revolution [87, 485, 888, 870, 107, 49, 50, 106, 197, 35, 107, 535, 821, 145]. revolutionary [88]. revolutions [108, 109, 108, 109]. rhetoric [754, 647, 395, 700]. rhetorical [670]. Rhine [608]. Rice [883, 798, 863, 822]. rich [710]. Richter [616]. Ridge [420]. Riedler [748]. Rigs [551]. Rigs-to-Reefs [551]. rise [206, 419, 718, 719, 5, 580]. Risk [504, 505, 406, 422, 813]. Risks [508]. rivalry [371, 733]. River [842, 799, 549]. Road [436, 167, 619, 719]. road/map [619]. roads [575]. Robert [384, 437]. robot [865, 566]. robotics [778]. rocketry [285]. rockets [639]. role [414, 253, 121, 68]. Rolf [285]. Roman [37]. Romania [876]. room [543, 712, 396]. Roosevelt [481]. root [264]. roumaines [37]. rounds [606]. Royal [363]. rubbers [396]. rubble [409]. rural [671, 848]. Russia [269, 424, 699].

sa [57]. Sabbath [816]. sacred [885]. sad [211]. sadness [775]. Safeguard [580]. Safety [650, 813, 804, 591]. Sage [245]. sales [372]. Salvarsan [654]. sand [881]. São [775]. Sarony [722]. Sarrià [261]. Satellite [206, 484, 207, 827]. satellites [843, 637, 208]. Saxony [145, 35]. scale [526]. Scandinavia [739]. scape [383]. Scarcity [491]. scenarios [49]. scene [793]. Schatzberg [867]. Schering [658]. school [763, 877, 13]. schools [273, 275]. Schumpeter [15]. Science [94, 127, 492, 88, 475, 338, 706, 123, 303, 161, 352, 720, 707, 476, 433, 222, 697, 835, 83, 677, 520, 673, 772, 87, 730, 637, 100, 638, 28, 674, 759, 62, 187, 284, 341, 829, 819, 746, 402, 717, 80, 451, 509, 395, 771, 376, 738, 5, 262, 364, 70, 768, 78, 683, 296, 450, 94, 127, 674, 479, 430]. Sciences [469, 332, 233, 234, 191, 78, 191]. Scientific [644, 471, 275, 190, 86, 740, 655, 116, 132, 283, 107, 653, 658, 106, 654, 206, 95, 108, 109, 788, 376, 207, 781, 290, 191, 530, 537]. scientific-technological [781]. scientifique [107, 95]. scientifiques [108, 109]. scientists [422, 189, 99, 746, 299, 429, 428]. screens [665]. screw [116, 132]. screws [179]. sea [383, 880]. seams [412]. search [538, 883, 800]. Searching [339]. Sébastien [402]. second [166, 11, 753, 614, 679, 35, 234, 244, 595, 12]. seconde [11]. secrecy [772, 257, 357]. secret [628, 526]. sector [397]. Secularism [885]. Security [594, 422, 372, 777, 262]. Seeing [782, 639]. Seguin [140, 135, 160, 43, 136, 139]. Seguins [160]. Seiler [582, 583]. Self [688, 388, 684]. self-administration [388]. Self-knowledge [688]. Selling [734, 657]. semantics [774]. semi [151]. semi-technical [151]. semiconductor [252, 544, 253, 364, 353]. Senegal [887, 871, 576]. sensing [210]. sentiment [590]. September [410]. Serengeti [448]. serfdom [719]. serial [869]. Serve [210]. service [511]. settler [820]. seventeenth [628, 540]. seventeenth-century [628]. Sewol [857]. sexuality [717]. Seymour [810]. shades [315]. shaped [790, 831]. shaping [594, 376, 445]. shelterbelts [550]. Sherry [810]. shift [697]. Shilongba [861]. shipbuilding [270]. shipping [363]. Shippingport [144]. ships [755]. shipyards [412]. Shockley [641]. Shoddy [694]. shop [761]. Shortwave [243]. shot [843]. should [528]. shows [547]. shuttle [401]. siècle [100, 199, 165, 166, 95, 11]. siècles [202]. Siemens [250, 364]. silence [312]. Silent [507]. Silicon [353, 641]. silk [511]. silver [725]. Simmons [522]. simulacrum [743]. since [201, 166, 407, 223]. sins [211]. SITE [827]. sixteenth [18]. sketch [301]. skidder [414]. skill [268]. SLAC [758]. Slavery [485]. sleeping [656]. slept [390]. slowly [346]. sluice [510]. Small [755, 462, 820]. smoke [734]. smokeless [734]. Social [668, 496, 233, 234, 717, 325, 781, 428]. socialist [821, 806, 824]. Société [124, 147, 123]. Society [504, 440, 505, 124, 421, 54, 459, 451, 10, 442, 447]. sociology [84]. sociotechnical [817]. soda [236]. Soil [168]. Sols [168]. Some [131, 269, 80, 158, 184, 79, 145]. Something [698]. son [96]. sort [7, 12]. Soto [92]. soude [236]. soul [743]. Sound [736, 620]. sounding [399]. Sounds [738]. sources [288]. Souris [549]. sous [127]. south [770, 692, 760, 863, 876, 823, 796, 857, 716]. southern [848, 669]. sovereignty [711, 856]. Soviet [491, 347, 189, 529, 703, 437, 650, 313]. Soviet-designed [650]. Space [209, 633, 228, 437, 296, 705, 637, 674, 213, 212, 636, 856, 757, 401, 851, 788, 207, 352, 260, 705]. space-age [705]. spaceflight [634]. Spacelab [352]. spaceship [635]. Spain [725, 873, 540]. Spanish [397, 261, 680, 854]. sparrow [887, 871]. Speaking [429, 271]. special [88, 802]. specialists [594]. Sperry [545]. spirit [334]. spirits [731]. spiritualité [94]. spirituality [94]. spot [571]. Spreading [732]. Spring} [507]. Sputnik [347]. Sputniks [377]. stadium [225]. Stag [474]. stage [35]. Stages [245]. Standard [116, 132, 538]. standardization [342, 589]. Standardizing [620, 690, 657]. standards [339]. star [429]. stars [757]. startup [751]. Stata [522]. State [251, 497, 17, 435, 540, 390, 864, 857, 739, 690, 275, 707, 697]. state-controlled [864]. States [430, 856, 16, 61, 671, 527, 790, 409, 877, 846, 247, 589, 629]. station [403, 861, 144]. Status [526]. steam [55, 363, 11, 614]. steam-engine [11]. steamboat [282]. Stephenson [136]. stewardship [681]. stock [444]. stones [767]. story [863, 57]. storytelling [719]. Strategic [365]. Strategies [265, 256, 311, 670, 364]. strategy [251, 250]. Strath [423]. Streams [799]. street [752]. streets [812]. strength [141]. String [490]. Stroke [434]. Strontium [616]. Strophanthin [655]. Structural [413]. Structure [232]. structures [709]. Struggle [560]. Struggles [403]. STS [449]. students [311, 774]. Studies [295, 339, 611, 451, 450, 313]. studio [602]. Study [440, 705, 357]. studying [191]. subject [613, 756]. submarine [737]. suburbs [812]. success [384, 265]. Successful [320]. sugar [628]. sugar-makers [628]. Suharto [817]. Suisse [116]. superheater [777]. supply [848]. support [99, 559]. surgery [728]. Surveillance [388, 843, 735, 737]. Survey [842, 750, 151, 601, 115]. surveying [693]. suspendu [135]. suspension [138, 135]. Suspicious [371]. Sustainable [452]. Svedberg [433]. Sweden [677, 400, 740, 225, 581]. Swedish [309, 868, 849, 741]. Sweet [592]. swinging [631]. switching [453]. Switzerland [851]. Synchrotron [762]. Synthesis [546, 806]. synthetic [278]. system [160, 559, 690, 619, 395, 428]. système [160]. systems [848, 345, 384, 414, 457].

T [238]. tablet [721]. tachistoscope [627]. tailoring [837]. Tain [135]. Taj [771]. tale [597]. Tales [434]. Tancarville [138]. taste [678]. teaching [727]. teamwork [520]. tech [542]. techné} [717]. Technical [497, 229, 202, 61, 146, 54, 269, 360, 238, 511, 824, 356, 160, 644, 813, 342, 186, 556, 826, 358, 821, 268, 357, 825, 819, 121, 540, 70, 35, 151, 313, 108]. technique [202, 160, 798]. Techniques [165, 108, 116, 113, 4]. Techno [509, 675, 576, 419]. techno-nationalism [675, 419]. techno-politics [576]. Technocracy [499, 754]. Technocratic [497, 496, 782]. technodiplomacy [753]. Technological [742, 471, 754, 424, 750, 282, 439, 743, 500, 474, 816, 397, 536, 16, 63, 146, 49, 675, 324, 213, 727, 755, 231, 783, 721, 376, 809, 781, 260, 5, 226, 601]. Technologie [37, 2, 13, 233]. Technologies [482, 512, 438, 388, 881, 201, 735, 737, 772, 837, 613, 576, 698, 212, 718, 841, 858, 776, 812, 457, 258, 218]. technologies-in-use [837]. Technology [8, 487, 652, 18, 867, 552, 839, 596, 668, 440, 477, 492, 466, 553, 706, 394, 15, 402, 455, 436, 562, 802, 161, 1, 383, 489, 215, 716, 595, 10, 415, 707, 476, 814, 177, 465, 599, 458, 242, 789, 144, 885, 421, 332, 864, 320, 281, 211, 185, 351, 250, 513, 3, 4, 233, 234, 270, 389, 550, 28, 759, 571, 527, 399, 543, 341, 276, 651, 602, 378, 859, 20, 647, 382, 877, 224, 410, 225, 414, 210, 554, 819, 579, 37, 574, 817, 729, 459, 149, 41, 2, 9, 866, 364, 26, 244, 580]. technology [815, 396, 780, 184, 688, 223, 296, 663, 528, 176, 450, 313, 588, 770, 514, 442, 867, 13]. Technology-Led [489]. Technometrics [159, 64]. technométrie [64]. Technometry [41, 150]. Technopedagogies [524]. technopolitical [408]. technoscience [633, 577]. technoscientists [741]. Technospatial [714]. Tekniska [225]. telecommunication [607]. telecommunications [31, 199, 243, 31]. telegrams [732]. telegraph [202, 690]. télégraphe [202]. telegraphs [731]. Telegraphy [69, 730]. Telephone [467, 664, 793, 397, 790, 792]. telephonic [789]. telephony [791]. televised [637]. Television [678, 637, 844, 674, 827]. telex [202, 202]. Telling [434]. Tempier [101]. Temple [263]. temps [108]. Ten [351]. terabytes [830]. term [677, 49, 509]. terms [700]. terrains [134]. Terre [168]. Territoriality [596]. Territory [820]. terrorism [408]. terrors [613]. test [263]. Textile [264]. textiles [394]. Texts [334]. Thalidomide [656]. theft [357]. their [497]. them [263]. Themes [272]. theoretical [238]. theory [23, 776, 412]. there [535, 109]. thermodynamics [6, 21, 51, 55]. theses [351]. thick [530]. things [829]. Thinking [625]. Third [234, 831, 491]. Thirty [451]. though [195]. Thought [496, 634, 281, 556]. Thoughts [731, 269]. thousand [461]. threads [116, 132]. threat [737]. Three [833, 514, 178]. throughout [214]. Thumba [720]. tidal [850]. tiller [760]. time [339, 108, 712, 879, 151]. Tinkering [805]. Tocqueville [585]. today [77, 606]. tomorrow [763]. tons [833]. tool [547]. tools [837]. topside [378]. Toronto [519]. Tour [678]. tourist [630]. Tournon [138, 135]. Tournon-Tain [135]. towers [749]. toxicity [804]. toys [523, 710]. Tracing [878]. trade [409]. trademark [659]. trading [512, 135]. trading-cloth [135]. tradition [102]. traditional [268]. Traditions [498, 702]. traffic [620]. trail [314]. training [27, 268, 854]. trajectories [878]. Trans [484]. Trans-National [484]. transcendence [638]. Transfer [436, 177, 144, 415]. transform [511]. Transformability [33]. transformations [876, 862]. transgenic [805]. transhumant [773]. Transistor [250, 249]. transit [281]. transition [849, 852]. transitions [320, 851]. Translation [717]. translations [513]. transmission [357]. Transnational [836, 829, 621, 753, 605, 793]. transport [823, 662]. tray [786]. treatment [388]. tree [414]. Trees [550]. trials [340]. tropical [795, 797]. troublesome [565]. trust [732, 592, 591, 588]. Trypanosomiasis [797]. Tubab [576]. tube [136, 57]. tuberculosis [388]. tubes [57]. tubulaire [136]. tunnelling [882]. turbine [51]. turbine-condenser [51]. turbulent [880]. Turkle [810]. Turning [855]. Turriano [44]. TVA [799]. Twelve [707]. Twelve-Year [707]. twentieth [539, 837, 288, 516, 517, 750, 524, 747, 746, 733, 795, 851]. twentieth-century [746, 795, 851]. twenty [424]. twenty-first [424]. Twiggy [631]. Two [774, 346, 655, 518, 61, 606, 849]. Tycho [783]. type [281, 233, 234]. types [787, 346]. Typing [342].

U [727]. U.K. [736]. U.N. [8]. U.S. [839, 212, 210, 312, 228, 407, 800]. U.S.A. [150]. UFOs [638]. Uganda [773]. UK [256, 217, 521]. Ukraine [699]. umma} [817]. underground [726]. Understanding [213, 829, 574]. Understood [484]. unemployment [16]. uniform [247]. unintended [310]. Union [347, 529]. United [16, 61, 671, 535, 527, 790, 409, 877, 846, 247, 430, 589, 629]. Unity [498]. Univac [545]. Universe [497, 637]. university [692, 519]. Unpacking [384]. unsuccessful [284]. until [265]. untrained [627]. UPC [247]. Upheaval [520]. Upper [364, 566]. upper-air [566]. Uranium [838, 833]. urban [281, 734, 666, 752]. urban-industrial [734]. USA [850, 208]. usage [179]. Use [868, 502, 477, 887, 871, 351, 837, 846, 729]. usefulness [318]. user [210]. users [400, 294, 868, 792]. uses [258]. Using [491, 113, 611]. USSR [850, 191, 184]. utilisation [352]. utilitarian [869]. utopianism [424]. Utopias [389].

V [639]. V-2 [639]. Valley [887, 641, 871, 34]. value [761, 881]. values [522, 647]. vapeur [11]. Vaucanson [7, 12]. vehicles [869]. Venezuela [800]. venture [670]. vernacularization [714]. Verne [662]. version [710]. versus [141]. very [597]. Vesting [481]. veterinary [773]. via [245]. viaducts [610]. Victor [844]. Victoria [797]. Victorian [567, 571, 568]. Victorians [569]. Victory [790, 284]. video [577]. vie [95]. view [420]. Vinci [131, 90]. vingt [198]. vint [171]. violence [857]. viral [878]. Virtual [705]. visages [104]. Vision [567, 627, 736]. Visions [691, 781, 733]. visual [735, 674, 759, 611, 670]. visualization [671]. Visualizing [647, 610]. Vitruvius [178]. voice [574]. volumes [399]. Voskuhl [684]. Voting [440]. voyages [376]. vs [285, 352].

waiting [712]. Wakefield [700]. Walter [751]. War [491, 251, 520, 370, 575, 466, 249, 849, 739, 481, 842, 741, 756, 485, 818, 289, 499, 248, 726, 676, 823, 882, 865, 255, 636, 791, 718, 738, 798, 728, 735, 737, 464, 711, 790, 378, 669, 714, 700, 738, 244, 151, 595, 534, 824, 707, 258, 296, 591]. warez [793]. wars [518, 61, 243, 358, 429]. Wartime [371]. Was [263, 252]. Washington [390]. waste [862, 710, 694, 681]. Water [447, 512, 463, 848, 23, 841, 57]. water- [512]. water-tube [57]. water-works [463]. waterclocks [178]. waterfowl [549]. watery [551]. waves [363]. Way [445]. ways [835, 413, 576]. Weak [412]. weather [566]. Web [609]. Weimar [570]. welding [412]. wells [196]. were [843, 615]. West [837, 421, 403, 883, 365, 815, 792]. Western [466, 10]. wet [722, 447]. wet-plate [722]. while [690]. White [678, 284]. Who [619, 372, 824]. Whose [827]. Wide [609]. Wild [804]. Will [462, 637]. Willem [490]. William [641]. Wind [848]. wind-powered [848]. windmill [320]. windmill-technology [320]. windmills [512]. wings [869]. Winkler [153]. winter [755]. Wire [666]. wireless [730, 731, 733]. Wires [664]. Wirtschaftswunder} [364]. within [242]. without [869]. witnessing [705]. wizards [447]. Women [309, 314, 308, 310, 312, 313, 686, 311, 407]. wonderful [463]. Wonderland [418]. Wood [141, 849]. Wooden [188]. word [389]. Work [325, 876, 850, 389, 187, 788, 149, 625, 528]. worker [343]. working [337]. works [463, 273]. workshops [7, 12]. World [728, 735, 737, 467, 560, 662, 790, 738, 244, 595, 258, 463, 772, 518, 61, 675, 535, 643, 409, 523, 669, 214, 315, 682, 733, 534, 251, 520, 370, 575, 609, 249, 849, 739, 481, 741, 756, 485, 476]. worlds [767]. writing [732, 389, 324].

X [698]. X-rays [698]. XI [188]. XIV [99]. XIX [100, 188]. XIXe [202, 165, 11]. XVII [95]. XVIIIème [166]. XVIIth [95]. XX [165]. XXe [202]. XXeme [199]. XXth [299].

Y2K [330]. Ya [187]. Year [879, 707, 877]. years [289, 198, 451, 708]. yesterday [77].

Zealand [552]. Zimbabwe [713]. Zone [313]. Zyklon [258].

References

Redondi:1983:HTRa

[1] Pietro Redondi. History and technology: Research on the borderline. History and Technology, 1(1):1–6, 1983. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Redondi:1983:HTRb

[2] Pietro Redondi. Histoire et technologie: Une recherche de frontière. (French) [History and technology: a frontier research]. History and Technology, 1(1):7–12, 1983. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Febvre:1983:RHTa

[3] Lucien Febvre. Reflections on the history of technology. History and Technology, 1(1):13–18, 1983. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Febvre:1983:RHTb

[4] Lucien Febvre. Réflexions sur l’histoire des techniques. (French) [Reflections on the history of technology]. History and Technology, 1(1):19–24, 1983. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sebestik:1983:RTS

[5] Jan Sebestik. The rise of the technological science. History and Technology, 1(1):25–43, 1983. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Duffy:1983:MTL

[6] M. C. Duffy. Mechanics, thermodynamics and locomotive design: The machine-ensemble and the development of industrial thermodynamics. History and Technology, 1(1):45–78, 1983. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

DePlace:1983:SAV

[7] Dominique De Place. Le sort des ateliers de Vaucanson, 1783–1791, d’après un document nouveau. (French) [The fate of workshops in Vaucanson, 1783–1791, according to a new document]. History and Technology, 1(1):79–100, 1983. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1983:HTI

[8] Anonymous. History and Technology, an International Journal in the present: A U.N. meeting. History and Technology, 1(1):101–111, 1983. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Salomon:1984:WTI

[9] Jean-Jacques Salomon. What is technology? The issue of its origins and definitions. History and Technology, 1(2):113–156, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Torstendahl:1984:TDS

[10] Rolf Torstendahl. Technology in the development of society 1850–1980. Four phases of industrial capitalism in Western Europe. History and Technology, 1(2):157–174, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Payen:1984:PFI

[11] Jacques Payen. La position de la France dans l’industrie européenne des machines à vapeur durant la seconde moitié du XIXe siècle. (French) [The position of France in European steam-engine industry in the second half of the nineteenth century]. History and Technology, 1(2):175–211, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

dePlace:1984:SAV

[12] Dominique de Place. Le sort des ateliers de Vaucanson, 1783–1791, d’après un document nouveau. 2ème partie. (French) [The fate of workshops in Vaucanson, 1783–1791, according to a new document. Second part]. History and Technology, 1(2):213–237, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Woronoff:1984:TED

[13] Denis Woronoff. La technologie à l’École du Louvre. (French) [Technology at the Louvre School]. History and Technology, 1(2):239–241, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1984:E

[14] Anonymous. Editorial. History and Technology, 1(3–4):243–247, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mackenzie:1984:TCM

[15] Michael Mackenzie. Technology and crisis; Marx, Schumpeter and beyond. History and Technology, 1(3–4):249–275, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fano:1984:PTU

[16] Ester Fano. The problem of “technological unemployment” in the industrial research of the 1930’s in the United States. History and Technology, 1(3–4):277–306, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Caron:1984:EIS

[17] François Caron and Christine Berthet. Electrical innovation: State initiative or private initiative? Observations on the 1881 Paris exhibition. History and Technology, 1(3–4):307–318, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Benoit:1984:TCG

[18] Paul Benoit. Technology and crisis: The Great Depression of the Middle Ages and the technology of the Renaissance (fourteenth to sixteenth centuries). History and Technology, 1(3–4):319–334, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Payen:1984:AMM

[19] Jacques Payen. Ad memoriam: Maurice Daumas, 1910–1984. History and Technology, 1(3–4):335–341, 1984. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Konig:1985:IPT

[20] Wolfgang König. Ideology and practice of technology in history. History and Technology, 2(1):1–15, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Duffy:1985:MTL

[21] Michael C. Duffy. Mechanics, thermodynamics and locomotive design: Analytical methods and the classic form. History and Technology, 2(1): 17–61, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Piaskowski:1985:MPH

[22] Jerzy Piaskowski. Metallurgy in Plinius’ Historia naturalis. History and Technology, 2(1):63–75, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Redondi:1985:AWG

[23] Pietro Redondi. Along the water: The genius and the theory D’Alembert, Condorcet and Bossut and the Picardy Canal controversy. History and Technology, 2(1):77–110, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1985:PR

[24] Anonymous. Publications received 1983–1984. History and Technology, 2 (1):113–114, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1985:EBa

[25] Anonymous. Editorial board. History and Technology, 2(1):ebi, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sigaut:1985:MET

[26] François Sigaut. More (and enough) on technology! History and Technology, 2(2):115–132, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fox:1985:BPE

[27] Robert Fox and Anna Guagnini. Britain in perspective: The European context of industrial training and innovation, 1880–1914. History and Technology, 2(2):133–150, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gooding:1985:ECF

[28] David Gooding. Experiment and concept formation in electromagnetic science and technology in England in the 1820s. History and Technology, 2(2):151–176, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cazenobe:1985:ENE

[29] Jean Cazenobe. Essence et naissance de l’électroaimant. (French) [Gasoline and birth of the electromagnet]. History and Technology, 2(2):177– 201, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sleeswyk:1985:PBI

[30] André Wegener Sleeswyk. Pyramid building as an integrated process. History and Technology, 2(2):203–230, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Garcon:1985:PHT

[31] François Garçon. Pour une histoire des télécommunications et de l’informatique. (French) [For a history of telecommunications and informatics]. History and Technology, 2(2):231–234, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1985:EBb

[32] Anonymous. Editorial board. History and Technology, 2(2):ebi, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wislicki:1985:TBM

[33] Alfred Wiślicki. Transformability of basic models of machines. History and Technology, 2(3):235–244, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Misa:1985:CMC

[34] Thomas J. Misa. The changing market for chemical knowledge: Applied chemistry and chemical engineering in the Delaware Valley, 1851–1929. History and Technology, 2(3):245–268, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Stutzner:1985:DTE

[35] Heinz Stützner and Dagmar Szöllösi. The development of technical education during the second stage of the industrial revolution in Saxony. History and Technology, 2(3):269–282, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cloarec:1985:MIF

[36] Gisèle Cloarec. Ménage et industrie. (French) [Household and industry]. History and Technology, 2(3):283–284, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mihail:1985:TCR

[37] Dr. Zamfira Mihail. Technologie et culture roumaines. (French) [Roman technology and culture]. History and Technology, 2(3):285–288, 1985. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1985:A

[38] Anonymous. Abstracts. History and Technology, 2(3):289–291, 1985. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1985:E

[39] Anonymous. Errata. History and Technology, 2(3):293, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1985:EBc

[40] Anonymous. Editorial board. History and Technology, 2(3):ebi, 1985. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Piaskowski:1986:TMH

[41] Jerzy Piaskowski and Helena Piaskowska. Technometry as a methodology of the history of technology. History and Technology, 2(4):295–308, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Willis:1986:CLJ

[42] John Willis. Le canal de Lachine jusqu’en 1870: Origine et fonction d’un canal hydraulique. (French) [The Lachine Canal up to 1870: Origin and function of a hydraulic canal]. History and Technology, 2(4):309–329, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gillispie:1986:FSP

[43] Charles C. Gillispie. The Fonds Seguin at Privas. History and Technology, 2(4):331–334, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Nordon:1986:IAB

[44] Marcel Nordon. Introduction and analysis of the 21 books of devices and of machines by Pseudo Juanelo Turriano. History and Technology, 2(4): 335–365, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cardwell:1986:R

[45] Donald S. L. Cardwell, Charles C. Gillispie, and André Guillerme. Reviews. History and Technology, 2(4):367–372, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1986:PR

[46] Anonymous. Publications received 1984–1985. History and Technology, 2 (4):373–374, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1986:Aa

[47] Anonymous. Abstracts. History and Technology, 2(4):375–377, 1986. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1986:EBa

[48] Anonymous. Editorial board. History and Technology, 2(4):ebi, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Furtado:1986:LTS

[49] Andre Furtado. Long term scenarios for Latin America and the new technological revolution. History and Technology, 3(1):1–23, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guillerme:1986:LCI

[50] André Guillerme. From lime to cement: The industrial revolution in French civil engineering (1770–1850). History and Technology, 3(1):25–85, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Duffy:1986:MTL

[51] M. C. Duffy. Mechanics, thermodynamics & locomotive design: The turbine-condenser locomotive. History and Technology, 3(1):87–122, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1986:Ab

[52] Anonymous. Abstracts. History and Technology, 3(1):123–126, 1986. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1986:EBb

[53] Anonymous. Editorial board. History and Technology, 3(1):ebi, 1986. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gille:1987:TPS

[54] Bertrand Gille. Technical progress and society: Historical perspective. History and Technology, 3(2):127–153, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Duffy:1987:MTL

[55] M. C. Duffy. Mechanics, thermodynamics & locomotive design: The high pressure steam locomotive. History and Technology, 3(2):155–192, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Stolz:1987:CBD

[56] Rüdiger Stolz and Rüdiger Schwaiberger. The correlation between dye chemistry and pharmacy in creating the modern chemotherapy. History and Technology, 3(2):193–203, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zelbstein:1987:LIJ

[57] Uri Zelbstein. L’histoire d’une invention: Julien Belleville et sa chaudière à tubes d’eau. (French) [The story of an invention: Julien Belleville and his water-tube boiler]. History and Technology, 3(2):205–218, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:Aa

[58] Anonymous. Abstracts. History and Technology, 3(2):221–223, 1987. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:EBa

[59] Anonymous. Editorial board. History and Technology, 3(2):ebi, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bernstein:1987:RAM

[60] Michael A. Bernstein. The response of American manufacturing industries to the Great Depression. History and Technology, 3(3):225–248, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fano:1987:TPD

[61] Ester Fano. Technical progress as a destabilizing factor and as an agent of recovery in the United States between the two world wars. History and Technology, 3(3):249–274, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guay:1987:CSA

[62] Yves Guay. Les centres de science américains: Évolution d’une institution culturelle au pays de la libre entreprise. (French) [US science centers: evolution of a cultural institution in the country of free enterprise]. History and Technology, 3(3):275–328, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fano:1987:EAT

[63] Ester Fano. Early American technological Edens. History and Technology, 3(3):329–338, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bulferetti:1987:PTC

[64] Luigi Bulferetti. À propos de technométrie comme méthode historique. (French) [On technometrics as a historical method]. History and Technology, 3(3):339–341, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See reply [65].

Piaskowski:1987:RB

[65] Jerzy Piaskowski and Helena Piaskowska. Reply to L. Bulferretti. History and Technology, 3(3):343, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See [64].

Anonymous:1987:Ab

[66] Anonymous. Abstracts. History and Technology, 3(3):345–348, 1987. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:EBb

[67] Anonymous. Editorial board. History and Technology, 3(3):ebi, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Presutti:1987:PQA

[68] William D. Presutti, Jr. Productivity: Questioning the assumptions for improvement — a role for applied history. History and Technology, 3(4): 349–364, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Butrica:1987:TGE

[69] Andrew J. Butrica. Telegraphy and the genesis of electrical engineering institutions in France, 1845–1895. History and Technology, 3(4):365–380, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Simonenko:1987:FTS

[70] O. D. Simonenko. Formation of the technical science in electrical engineering. History and Technology, 3(4):381–396, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:PR

[71] Anonymous. Publications received 1986–1987. History and Technology, 3 (4):397–398, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:Ac

[72] Anonymous. Abstracts. History and Technology, 3(4):399–401, 1987. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:E

[73] Anonymous. Erratum. History and Technology, 3(4):403, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:EBc

[74] Anonymous. Editorial board. History and Technology, 3(4):ebi, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Redondi:1987:F

[75] Pietro Redondi. Foreword. History and Technology, 4(1–4):1–6, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Canguilhem:1987:P

[76] Georges Canguilhem. Preface. History and Technology, 4(1–4):7–10, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Clavelin:1987:DKD

[77] Maurice Clavelin. Le débat Koyré–Duhem, hier et aujourd’hui. (French) [The Koyré–Duhem debate, yesterday and today]. History and Technology, 4(1–4):13–35, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Taton:1987:AKE

[78] René Taton. Alexandre Koyré et l’essor de l’histoire des sciences en France (1933 à 1964). (French) [Alexandre Koyré and the development of the history of science in France (1933–1964)]. History and Technology, 4(1– 4):37–53, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cohen:1987:AKA

[79] I. Bernard Cohen. Alexandre Koyré in America — some personal reminiscences. History and Technology, 4(1–4):55–70, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Murdoch:1987:AKH

[80] John. E. Murdoch. Alexandre Koyré and the history of science in America: Some doctrinal and personal reflections. History and Technology, 4(1– 4):71–79, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Crombie:1987:AKG

[81] Alistair C. Crombie. Alexandre Koyré and Great Britain: Galileo and Mersenne. History and Technology, 4(1–4):81–92, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Casini:1987:CKI

[82] Paolo Casini. Considérations sur Koyré et l’Italie. (French) [considerations on Koyré and Italy]. History and Technology, 4(1–4):93–99, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Amsterdamski:1987:PSF

[83] Stefan Amsterdamski. La philosophie en science. (French) [Philosophy in science]. History and Technology, 4(1–4):103–113, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Elkana:1987:AKB

[84] Yehuda Elkana. Alexandre Koyré: Between the history of ideas and sociology of disembodied knowledge. History and Technology, 4(1–4):115–148, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schuhmann:1987:KPA

[85] Karl Schuhmann. Koyré et les phénoménologues allemands. (French) [Koyré and German phenomenologists]. History and Technology, 4(1– 4):149–167, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Biagioli:1987:MKT

[86] Mario Biagioli. Meyerson and Koyré: Toward a dialectic of scientific change. History and Technology, 4(1–4):169–182, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cohen:1987:REE

[87] I. Bernard Cohen. Revolution, evolution and emergence in the development of modern science. History and Technology, 4(1–4):183–211, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gillispie:1987:SPS

[88] Charles C. Gillispie. Science and politics, with special reference to revolutionary and Napoleonic France. History and Technology, 4(1–4):213–223, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hall:1987:KDE

[89] Marie Boas Hall. Koyré and the development of empiricism in the later Renaissance. History and Technology, 4(1–4):225–233, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Galluzzi:1987:LVE

[90] Paolo Galluzzi. Leonardo da Vinci: From the “elementi macchinali” to the man-machine. History and Technology, 4(1–4):235–265, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ashworth:1987:INP

[91] William B. Ashworth, Jr. Iconography of a new physics. History and Technology, 4(1–4):267–297, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wallace:1987:EJH

[92] William A. Wallace. The early Jesuits and the heritage of Domingo de Soto. History and Technology, 4(1–4):301–320, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Carugo:1987:JPN

[93] Adriano Carugo. Les Jésuites et la philosophie naturelle de Galilée: Benedictus Pererius et le De motu gravium de Galilée. (French) [Jesuits and the natural philosophy of Galileo: Benedictus Pererius and galileo’s De motu gravium]. History and Technology, 4(1–4):321–333, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Costabel:1987:SSM

[94] Pierre Costabel. Science et spiritualité: La mouvance de l’Oratoire. (French) [Science and spirituality: The movement of the Oratory]. History and Technology, 4(1–4):335–340, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Labrousse:1987:RVS

[95] Elisabeth Labrousse. Les réformés et la vie scientifique en France au XVII^{e} siècle: Pistes d’explication d’une carence. (French) [Reforms and scientific life in France in the XVIIth century: possible explanations of deficiency]. History and Technology, 4(1–4):341–349, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Beaulieu:1987:IDD

[96] Armand Beaulieu. Importance du dialogue: Mersenne et son groupe. (French) [Importance of dialogue: Mersenne and his group]. History and Technology, 4(1–4):351–364, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Westman:1987:PCA

[97] Robert S. Westman. La préface de Copernic au pape: Esthétique humaniste et réforme de l’église. (French) [The preface of Copernicus to the Pope: Humanist aesthetics and church reform]. History and Technology, 4(1–4): 365–384, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Westfall:1987:PPG

[98] Richard S. Westfall. Patronage and the publication of Galileo’s Dialogue. History and Technology, 4(1–4):385–399, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hahn:1987:CPS

[99] Roger Hahn. Changing patterns for the support of scientists from Louis XIV to Napoleon. History and Technology, 4(1–4):401–411, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fox:1987:PCP

[100] Robert Fox. La professionnalisation: Un concept pour l’historien de la science française au XIX^{e} siècle. (French) [Professionalisation: a concept for the historian of French science in the nineteenth century]. History and Technology, 4(1–4):413–422, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Beaujouan:1987:AKE

[101] Guy Beaujouan. Alexandre Koyré, l’évêque Tempier et les censures de 1277. (French) [Alexander Koyré, Bishop Tempier and the censures of 1277]. History and Technology, 4(1–4):425–429, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lindberg:1987:CDH

[102] David C. Lindberg. Continuity and discontinuity in the history of optics: Kepler and the medieval tradition. History and Technology, 4(1–4):

431–448, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Rashed:1987:PMC

[103] Roshdi Rashed. La périodisation des mathématiques classiques. History and Technology, 4(1–4):449–454, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Halleux:1987:KPM

[104] Robert Halleux. Koyré parmi les masques et les visages de Paracelse. (French) [Koyré among masks and the faces of Paracelsus]. History and Technology, 4(1–4):455–464, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zambelli:1987:HME

[105] Paola Zambelli. Hermétisme, mystique, empirisme. History and Technology, 4(1–4):465–483, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hall:1987:AKS

[106] A. Rupert Hall. Alexandre Koyré and the scientific revolution. History and Technology, 4(1–4):485–496, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Coumet:1987:AKR

[107] Ernest Coumet. Alexandre Koyré: La révolution scientifique introuvable?. (French) [Alexandre Koyré: the lost scientific revolution?]. History and Technology, 4(1–4):497–529, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Landes:1987:TRS

[108] David S. Landes. Techniques et révolutions scientifiques: L’exemple de la mesure du temps. (French) [Technical and scientific revolutions: The example of the measurement of time]. History and Technology, 4(1–4): 531–541, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Moscovici:1987:ECQ

[109] Serge Moscovici. Est-ce qu’il y a des contre-révolutions scientifiques?. (French) [Are there scientific counter-revolutions?]. History and Technology, 4(1–4):543–559, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gil:1987:PQO

[110] Fernando Gil. La postérité de la “query n^{∘} 31”: Oppositions programmatiques et continuité explicative. (French) [The posterity of “query number 31”: Programmatic objections and explanatory continuity]. History and Technology, 4(1–4):561–573, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:C

[111] Anonymous. Contributors. History and Technology, 4(1–4):575–581, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1987:EBd

[112] Anonymous. Editorial board. History and Technology, 4(1–4):ebi, 1987. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Danko:1988:ALD

[113] George M. Danko and Friedrich B. Prinz. An analysis of lathe development using artificial intelligence techniques. History and Technology, 5(1):1–29, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Faltas:1988:IFO

[114] Sami Faltas. The invention of fibre-optic communications. History and Technology, 5(1):31–49, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guntau:1988:HOP

[115] Martin Guntau. The history of the origins of the Prussian Geological Survey in Berlin (1873). History and Technology, 5(1):51–58, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Brooks:1988:SSTa

[116] Randall C. Brooks. Standard screw threads for scientific instruments. Part I: Production techniques and the Filière Suisse. History and Technology, 5(1):59–76, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guillerme:1988:BR

[117] André Guillerme, Catherine Cardinal, and Antonio E. Ten. Book reviews. History and Technology, 5(1):77–85, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:Aa

[118] Anonymous. Abstracts. History and Technology, 5(1):87–90, 1988. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:EBa

[119] Anonymous. Editorial board. History and Technology, 5(1):ebi, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fox:1988:E

[120] Robert Fox and Pietro Redondi. Editorial. History and Technology, 5 (2–4):91–94, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Payen:1988:RCN

[121] Jacques Payen. The role of the Conservatoire national des arts et métiers in the development of technical education up to the middle of the 19th century. History and Technology, 5(2–4):95–138, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

DePlace:1988:BCP

[122] Dominique De Place. Le bureau de consultation pour les arts, Paris, 1791– 1796. (French) [The Consultation Office for the Arts, Paris, 1791–1796]. History and Technology, 5(2–4):139–178, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mandelbaum:1988:SFS

[123] Jonathan Mandelbaum. Science and friendship: The Société Philomathique de Paris, 1788–1835. History and Technology, 5(2–4):179–192, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Redondi:1988:NES

[124] Pietro Redondi. Nation et entreprise. La Société d’encouragement pour l’industrie nationale, 1801–1815. (French) [Nation and business. The Society for Encouragement of Domestic Industry, 1801–1815]. History and Technology, 5(2–4):193–222, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guillerme:1988:FNE

[125] André Guillerme. La formation des nouveaux édiles: Ingénieurs des Ponts et Chaussées et architectes, 1804–1815. (French) [Formation of new councilors: Engineers of bridges and highways and architects, 1804–1815]. History and Technology, 5(2–4):223–247, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Haines:1988:APB

[126] Barbara Haines. The Athénée de Paris and the Bourbon Restoration. History and Technology, 5(2–4):249–271, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

DeBertierDeSauvigny:1988:SPS

[127] Guillaume De Bertier De Sauvigny. Science et politique sous la restauration. (French) [Science and politics under the restoration]. History and Technology, 5(2–4):273–299, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Garcia-Diego:1988:MPL

[128] Jose A. Garcia-Diego and Eduardo L. Ortiz. On a mechanical problem of Lanz. History and Technology, 5(2–4):301–313, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:Ab

[129] Anonymous. Abstracts. History and Technology, 5(2–4):315–321, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:EBb

[130] Anonymous. Editorial board. History and Technology, 5(2–4):ebi, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Foley:1988:BVE

[131] Vernard Foley, Darlene Sedlock, Carole Widule, and David Ellis. Besson, da Vinci, and the evolution of the pendulum: Some findings and observations. History and Technology, 6(1):1–43, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Brooks:1988:SSTb

[132] Randall C. Brooks. Standard screw threads for scientific instruments. Part II: The British Association screw gauge. History and Technology, 6 (1):45–59, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:Ac

[133] Anonymous. Abstracts. History and Technology, 6(1):61–62, 1988. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fontanon:1988:LDP

[134] Claudine Fontanon. L’échec du projet industriel sur les terrains de Perrache, 1826–1832. (French) [The failure of the industrial project on the grounds of Perrache, 1826–1832]. History and Technology, 6(2):63–94, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cotte:1988:SCD

[135] Michel Cotte. Seguin et Cie (1806–1824): Du négoce familial de drap à la construction du pont suspendu de Tournon-Tain. (French) [Seguin and Co. (1806–1824): From trading-cloth family to the construction of the Tournon-Tain suspension bridge]. History and Technology, 6(2):95–144, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Payen:1988:SSN

[136] Jacques Payen. Seguin, Stephenson et la naissance de la locomotive à chaudière tubulaire (1828–1829). (French) [Seguin, Stephenson, and the birth of the locomotive boiler tube (1828–1829)]. History and Technology, 6(2):145–171, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:Ad

[137] Anonymous. Abstracts. History and Technology, 6(2):173–176, 1988. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Barjot:1988:TTC

[138] Dominique Barjot. From Tournon to Tancarville: The contribution of French civil engineering to suspension bridge construction, 1824–1959. History and Technology, 6(3):177–201, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Redondi:1988:PAC

[139] Pietro Redondi. Physique et apologetique. Le Cosmos de l’abbé Moigno et de Marc Seguin. (French) [Physics and apologetics. the Cosmos of Abbott Moigno and Marc Seguin]. History and Technology, 6(3):203–225, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Costabel:1988:CFP

[140] Pierre Costabel. La corrélation des forces physiques en France et an Angleterre (1843) et la réclamation de Marc Seguin en 1847. (French) [Correlation of physical forces in France and England (1843) and the reclamation of Marc Seguin in 1847]. History and Technology, 6(3):227–238, 1988. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guillerme:1988:WVI

[141] André Guillerme. Wood versus iron: The strength of materials in early 19th century France. History and Technology, 6(3):239–252, 1988. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:Ae

[142] Anonymous. Abstracts. History and Technology, 6(3):253–256, 1988. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:E

[143] Anonymous. Erratum. History and Technology, 6(3):257, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Beaver:1988:SAP

[144] William Beaver. The Shippingport Atomic Power Station: Problems in technology transfer. History and Technology, 6(4):257–273, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Muller:1988:SID

[145] Klaus Müller and Eberhard Wächtler. On some issues of the development of Saxony at the beginning of the Industrial Revolution. History and Technology, 6(4):275–298, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Frison:1988:TTI

[146] Guido Frison. Technical and technological innovation in Marx. History and Technology, 6(4):299–324, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Butrica:1988:SEP

[147] Andrew J. Butrica. Société d’encouragement pour l’industrie nationale. Preliminary inventory of the archives. History and Technology, 6(4):325– 334, 1988. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1988:Af

[148] Anonymous. Abstract. History and Technology, 6(4):335–337, 1988. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Perrin:1990:ITW

[149] Jacques Perrin. The inseparability of technology and work organization. History and Technology, 7(1):1–13, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bulferetti:1990:HTU

[150] Luigi Bulferetti. Historical technometry in the U.S.A. History and Technology, 7(1):15–29, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Teagarden:1990:ITS

[151] Ernest M. Teagarden. Innovation over time: A semi-technical survey of European military change before the Great War. History and Technology, 7(1):31–40, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zelbstein:1990:HDR

[152] Uri Zelbstein. Histoire du réseau autographique. (French) [History of the autographic network]. History and Technology, 7(1):41–49, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sorms:1990:CWC

[153] Bernhard Sorms. Clemens Winkler: A chemist at the Freiberg Mining Academy during the 2nd phase of industrialization in Germany. History and Technology, 7(1):51–61, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1990:BR

[154] Anonymous. Book review. History and Technology, 7(1):63–65, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1990:Aa

[155] Anonymous. Abstracts. History and Technology, 7(1):67–69, 1990. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1990:PR

[156] Anonymous. Publications received. History and Technology, 7(1):71–72, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1990:EBa

Tucker:1990:SMN

[158] Aviezer Tucker. Some methodological notes on the explanations to the invention of printing. History and Technology, 7(2):73–89, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zanzi:1990:OEM

[159] Luigi Zanzi. Observations on the epistemologic and methodological principles of historical technometrics. History and Technology, 7(2):91–117, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cotte:1990:STS

[160] Michel Cotte. Le système technique des Seguin en 1824–25. (French) [The technical system of the Seguins in 1824–25]. History and Technology, 7 (2):119–147, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Raj:1990:PPR

[161] Kapil Raj. The present in the past: A review of Debiprasad Chattopadhyaya, History of Science and Technology in Ancient India. The Beginnings (Calcutta: Firma KLM, 1986). Foreword by Joseph Needham; 532 pages and index. History and Technology, 7(2):149–156, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1990:Ab

[162] Anonymous. Abstracts. History and Technology, 7(2):157–159, 1990. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1990:EBb

[163] Anonymous. Editorial board. History and Technology, 7(2):ebi, 1990. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1991:E

[164] Anonymous. Editorial. History and Technology, 7(3–4):163–164, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guillerme:1991:TMC

[165] André Guillerme. Techniques et matériaux de construction XIXe–XX siècle, une introduction. (French) [Construction techniques and materials 19th–20th century, an introduction]. History and Technology, 7(3–4): 165–178, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Krolikowski:1991:LMC

[166] Lech Królikowski. L’Évolution des matériaux de construction en Pologne depuis la deuxième moitié du XVIIIème siècle. (French) [Evolution of building materials in Poland since the second half of the eighteenth century]. History and Technology, 7(3–4):179–196, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Scavizzi:1991:RMR

[167] C. Paola Scavizzi. Road materials and road construction in Italy between the mid-eighteenth and the mid-nineteenth century. History and Technology, 7(3–4):197–210, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guillerme:1991:CTM

[168] André Guillerme. La cervelle de la terre: La mécanique des sols et les fondations d’ouvrage de 1750 à 1830. (French) [The brains of the Earth: the Soil Mechanics and Foundations Project 1750–1830]. History and Technology, 7(3–4):211–254, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kahlow:1991:MCG

[169] Andreas Kahlow. Materials in 19th century Germany. History and Technology, 7(3–4):255–266, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lemoine:1991:FBF

[170] Bertrand Lemoine. Fer et beton en France (1850–1920). (French) [Iron and concrete in France (1850–1920)]. History and Technology, 7(3–4):267–278, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ordonez:1991:RCP

[171] Jose A. Fernandez Ordoñez. From reinforced concrete to prestressed concrete: “Un jour il me vint à l’esprit”. History and Technology, 7(3–4): 279–294, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Janse:1991:NMC

[172] H. Janse, G. G. Nieuwmeijer, and C. H. Van Eldik. New materials and constructions in The Netherlands. History and Technology, 7(3–4):295–320, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wislicki:1991:BCC

[173] Alfred Wiślicki. Building and construction of the 18th and 19th century. History and Technology, 7(3–4):321–341, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1991:Aa

[174] Anonymous. Abstracts. History and Technology, 7(3–4):343–347, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1991:EBa

[175] Anonymous. Editorial board. History and Technology, 7(3–4):ebi, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wislicki:1991:PAL

[176] Alfred Wiślicki. Phase-alternating logistic development of technology. History and Technology, 8(1):1–7, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Allen:1991:BMT

[177] Edward A. Allen. Business mentality and technology transfer in eighteenth-century France: The calandre anglaise at Nîmes, 1752–1792. History and Technology, 8(1):9–23, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sleeswyk:1991:TWD

[178] André W. Sleeswyk and Bjarue Huldén. The three waterclocks described by Vitruvius. History and Technology, 8(1):25–50, 1991. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Brooks:1991:OUP

[179] Randall C. Brooks. Origins, usage and production of screws: An historical perspective. History and Technology, 8(1):51–76, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1991:Ab

[180] Anonymous. Abstracts. History and Technology, 8(1):77–78, 1991. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1991:PR

[181] Anonymous. Publications received. History and Technology, 8(1):79, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1991:EBb

[182] Anonymous. Editorial board. History and Technology, 8(1):ebi, 1991. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1992:E

[183] Anonymous. Erratum. History and Technology, 8(2):80, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Voronkov:1992:ESO

[184] Yuri S. Voronkov. Editorial: Some observations on the history of technology in the USSR. History and Technology, 8(2):81–85, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Dorfman:1992:TCB

[185] V. F. Dorfman. The technology of the 20th century as a border line of evolution. History and Technology, 8(2):87–105, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gorokhov:1992:EAD

[186] V. G. Gorokhov. Engineering activity, designing and technical knowledge. History and Technology, 8(2):107–112, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gvozdetskh:1992:PHS

[187] V. L. Gvozdetskh. Problems of history of science and engineering in the work of I. Ya. Konfederatov (1902–1975). History and Technology, 8(2): 113–117, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Scherbo:1992:WPM

[188] G. M. Scherbo. Wooden pavements of Moscow (XI–XIX centuries). History and Technology, 8(2):119–125, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Denisyuk:1992:AHI

[189] Yu. N. Denisyuk and V. A. Gurikov. Advancement of holography, investigations by Soviet scientists. History and Technology, 8(2):127–131, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Apokin:1992:OFS

[190] I. A. Apokin and A. Z. Chapovski. The origins of the first scientific center for automation. History and Technology, 8(2):133–138, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sokolovskaya:1992:ISB

[191] Z. K. Sokolovskaya. Importance of scientific biographies published by USSR Academy of Sciences for studying the history of engineering and natural sciences. History and Technology, 8(2):139–143, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1992:Aa

[192] Anonymous. Abstracts. History and Technology, 8(2):147–150, 1992. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1992:EBa

[193] Anonymous. Editorial board. History and Technology, 8(2):ebi, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Dupuy:1992:F

[194] Gabriel Dupuy. Foreword. History and Technology, 8(3–4):iii–v, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guillerme:1992:NBC

[195] André Guillerme. Network: Birth of a category in engineering though during the French restoration. History and Technology, 8(3–4):151–166, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Barles:1992:PLO

[196] Sabine Barles. Point, ligne, ou réseau: Les puits artésiens d’absorption en France — 1820–1840. (French) [Point, line, or network: The artesian wells of absorption in France — 1820–1840]. History and Technology, 8(3– 4):167–191, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ribeill:1992:PHL

[197] Georges Ribeill. From pneumatics to highway logistics: André Michelin, instigator of the “automobile revolution”. History and Technology, 8(3– 4):193–216, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Blin:1992:PGV

[198] Eric Blin. La politique géopostale des vingt dernières années. (French) [Geopostal politics of the last 20 years]. History and Technology, 8(3– 4):217–229, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Griset:1992:ETI

[199] Pascal Griset. L’ evolution des telecommunications intercontinentales au XXeme siècle. History and Technology, 8(3–4):231–245, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Desportes:1992:HHN

[200] Marc Desportes. The history of highway nodes. History and Technology, 8(3–4):247–261, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Beltran:1992:FGN

[201] Alain Beltran. The French gas network and new technologies since 1946. History and Technology, 8(3–4):263–273, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Carre:1992:DTA

[202] Patrice A. Carré. Du télégraphe au télex: Histoire technique, histoire des réseaux, enjeux … (France XIXe–XXe siècles). (French) [From the telegraph to the telex: Technical history, history of networks, issues ldots (France 19th–20th centuries)]. History and Technology, 8(3–4):275–294, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1992:Ab

[203] Anonymous. Abstracts. History and Technology, 8(3–4):295–299, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1992:EBb

[204] Anonymous. Editorial board. History and Technology, 8(3–4):ebi, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Krige:1992:P

[205] John Krige. Preface. History and Technology, 9(1–4):iii–viii, 1992. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Krige:1992:RFE

[206] John Krige. The rise and fall of ESRO’s first major scientific project, the Large Astronomical Satellite (LAS). History and Technology, 9(1–4):1–26, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Russo:1992:CBP

[207] Arturo Russo. Choosing big projects in space research: The case of ESRO’s scientific satellite COS-B. History and Technology, 9(1–4):27–61, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Giget:1992:CPI

[208] Marc Giget. The conjunction of political, institutional and industrial issues in the choice of experimental communication satellites in Europe, compared to the USA and Japan. History and Technology, 9(1–4):63–82, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collette:1992:SCE

[209] René Collette. Space communications in Europe. How did we make it happen? History and Technology, 9(1–4):83–93, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mack:1992:MBT

[210] Pamela E. Mack. Making big technology serve the user: U.S. remote sensing programs. History and Technology, 9(1–4):95–107, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

DeMaria:1992:EEA

[211] Michelangelo De Maria and John Krige. Early European attempts in launcher technology: Original sins in Eldo’s sad parable. History and Technology, 9(1–4):109–137, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Logsdon:1992:CBT

[212] John M. Logsdon. Choosing big technologies examples from the U.S. space program. History and Technology, 9(1–4):139–150, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Johnson-Freese:1992:UJT

[213] Joan Johnson-Freese. Understanding Japanese technological decisionmaking: The case of space policy. History and Technology, 9(1–4):151–162, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pestre:1992:DMP

[214] Dominique Pestre. The decision-making processes for the main particle accelerators built throughout the world from the 1930s to the 1970s. History and Technology, 9(1–4):163–174, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Seidel:1992:TCE

[215] Robert W. Seidel. Technology choice in early high-energy physics. History and Technology, 9(1–4):175–187, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Walker:1992:BEN

[216] William Walker. The back-end of the nuclear fuel-cycle. History and Technology, 9(1–4):189–201, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gummett:1992:CMA

[217] Philip Gummett. Civil and military aircraft in the UK. History and Technology, 9(1–4):203–222, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Williams:1992:CBT

[218] Roger Williams. Choosing big technologies: The core issues. History and Technology, 9(1–4):223–234, 1992. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lesgards:1993:P

[219] Roger Lesgards. Preface. History and Technology, 10(1):v, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1993:Aa

[220] Anonymous. Acknowledgements. History and Technology, 10(1):vii, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schroeder-Gudehus:1993:PPM

[221] Brigitte Schroeder-Gudehus. Patrons and publics: Museums as historical artefacts. History and Technology, 10(1):1–3, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Williams:1993:SEM

[222] Mari Williams. Science, education and museums in Britain, 1870–1914. History and Technology, 10(1):5–11, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Weber:1993:PHM

[223] Wolfhard Weber. The political history of museums of technology in Germany since the nineteenth century. History and Technology, 10(1):13–25, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Levin:1993:CMT

[224] Miriam R. Levin. The city as a museum of technology. History and Technology, 10(1):27–36, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lindqvist:1993:OST

[225] Svante Lindqvist. An Olympic stadium of technology: Deutsches Museum and Sweden’s Tekniska Museet. History and Technology, 10(1):37– 54, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Staudenmaier:1993:CEM

[226] John M. Staudenmaier, S.J. Clean exhibits, messy exhibits: Henry Ford’s technological aesthetic. History and Technology, 10(1):55–65, 1993. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Rhees:1993:CAP

[227] David J. Rhees. Corporate advertising, public relations and popular exhibits: The case of Du Pont. History and Technology, 10(1):67–75, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Roland:1993:CEG

[228] Alex Roland. Celebration or education? The goals of the U.S. National Air and Space Museum. History and Technology, 10(1):77–89, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bergeron:1993:NGM

[229] Louis Bergeron. The new generation of museums: Technical, industrial and “ecomuseums”. History and Technology, 10(1):91–95, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1993:Ab

[230] Anonymous. Abstracts. History and Technology, 10(1):97–101, 1993. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Parayil:1993:MTC

[231] Govindan Parayil. Models of technological change: A critical review of current knowledge. History and Technology, 10(2–3):105–126, 1993. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Stamper:1993:SGM

[232] John W. Stamper and Robert Mark. Structure of the Galérie des Machines, Paris, 1889. History and Technology, 10(2–3):127–138, 1993. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Frison:1993:LBMa

[233] Guido Frison. Linnaeus, Beckmann, Marx and the foundation of technology. Between natural and social sciences: A hypothesis of an ideal type. First part: Linnaeus and Beckmann, cameralism, Oeconomia and technologie. History and Technology, 10(2–3):139–160, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Frison:1993:LBMb

[234] Guido Frison. Linnaeus, Beckmann, Marx and the foundation of technology. Between natural and social sciences: A hypothesis of an ideal type. Second and third parts: Beckmann, Marx, technology and classical economics. History and Technology, 10(2–3):161–173, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1993:Ac

[235] Anonymous. Abstracts. History and Technology, 10(2–3):175–177, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Grison:1993:HCL

[236] E. Grison. Hassenfratz, Carny et Lavoisier dans la compétition pour la fabrication de la soude artificielle en 1790. (French) [Hassenfratz, Carny and Lavoisier in the competition for the manufacture of artificial soda in 1790]. History and Technology, 10(4):179–197, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Olmo:1993:LC

[237] Carlo Olmo. L’ingénieur contesté. History and Technology, 10(4):199–214, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Poni:1993:CGE

[238] Carlo Poni. The craftsman and the good engineer: Technical practice and theoretical mechanics in J. T. Desaguliers. History and Technology, 10 (4):215–232, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Villasenor:1993:MDD

[239] Alejandro Tortolero Villaseñor. Le Ministère du développement au Mexique et les inventeurs des machines agricoles (1852–1903). (French) [The Ministry of Development of Mexico and inventors of agricultural machinery (1852–1903)]. History and Technology, 10(4):233–247, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1993:Ad

[240] Anonymous. Abstracts. History and Technology, 10(4):249–251, 1993. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kranakis:1994:I

[241] Eda Kranakis. Introduction. History and Technology, 11(1):1–5, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Aspray:1994:HCW

[242] William Aspray. The history of computing within the history of information technology. History and Technology, 11(1):7–19, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Headrick:1994:SRI

[243] Daniel R. Headrick. Shortwave radio and its impact on international telecommunications between the wars. History and Technology, 11(1): 21–32, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Small:1994:ETD

[244] James S. Small. Engineering, technology and design: The post-Second World War development of electronic analogue computers. History and Technology, 11(1):33–48, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hellige:1994:SAE

[245] Hans Dieter Hellige. From Sage via Arpanet to Ethernet: Stages in computer communications concepts between 1950 and 1980. History and Technology, 11(1):49–75, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Heide:1994:PCC

[246] Lars Heide. Punched-card and computer applications in Denmark 1911– 1970. History and Technology, 11(1):77–99, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Morton:1994:PHE

[247] Alan Q. Morton. Packaging history: The emergence of the uniform product code (UPC) in the United States, 1970–75. History and Technology, 11 (1):101–111, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Botelho:1994:ICD

[248] Antonio Botelho, François Jacq, and Dominique Pestre. Introduction to the conference “Le défi électronique de l’après-guerre, une perspective historique”. (French) [Introduction to the conference “The post-war electronic challenge, a historical perspective”]. History and Technology, 11 (2):113–119, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hoddeson:1994:RCR

[249] Lillian Hoddeson. Research on crystal rectifiers during World War II and the invention of the transistor. History and Technology, 11(2):121–130, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Erker:1994:CNT

[250] Paul Erker. The challenge of a new technology: Transistor research and business strategy at Siemens and Philips. History and Technology, 11(2): 131–143, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Andersen:1994:SSE

[251] Håkon With Andersen. State, strategy and electronics. Research in the political economy of post-World War II Norway. History and Technology, 11(2):145–164, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Botelho:1994:IPN

[252] Antonio José J. Botelho. The industrial policy that never was: French semiconductor policy, 1945–1966. History and Technology, 11(2):165–180, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Morris:1994:RMD

[253] Robin Morris. The role of the Ministry of Defence (MOD) in influencing the commercial performance of the British semiconductor industry. History and Technology, 11(2):181–193, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mounier-Kuhn:1994:FCM

[254] Pierre-E. Mounier-Kuhn. French computer manufacturers and the component industry, 1952–1972. History and Technology, 11(2):195–216, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Leslie:1994:EGI

[255] Stuart W. Leslie and Robert H. Kargon. Electronics and the geography of innovation in post-war America. History and Technology, 11(2):217–231, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Charles:1994:LER

[256] David R. Charles. The location of electronics R&D in the UK: Business strategies, organisation and labour markets. History and Technology, 11 (2):233–257, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Long:1994:MSA

[257] Pamela O. Long and Alex Roland. Military secrecy in antiquity and early medieval Europe: A critical reassessment. History and Technology, 11 (2):259–290, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Weindling:1994:UAB

[258] Paul Weindling. The uses and abuses of biological technologies: Zyklon B and gas disinfestation between the First World War and the Holocaust. History and Technology, 11(2):291–298, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

MacLeod:1994:ACA

[259] Roy MacLeod. The atom comes to Australia: Reflections on the Australian nuclear programme, 1953 and 1993. History and Technology, 11 (2):299–315, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sebesta:1994:PTC

[260] Lorenza Sebesta. The politics of technological cooperation in space: US– European negotiations on the post-Apollo programme. History and Technology, 11(2):317–341, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Raposo:1994:CEE

[261] Núria Puig Raposo and Santiago M. López García. Chemists, engineers and entrepreneurs: the Chemical Institute of Sarrià’s impact on Spanish industry (1916–1992). History and Technology, 11(3–4):345–359, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Seidel:1994:ANS

[262] Robert W. Seidel. Accelerators and national security the evolution of science policy for high-energy physics, 1947–1967. History and Technology, 11(3–4):361–391, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sobolev:1994:DTM

[263] D. A. Sobolev. Du Temple and Mozhaiskii: Was either of them the first to flight test an aeroplane? History and Technology, 11(3–4):393–401, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Travis:1994:BBR

[264] Anthony S. Travis. Between broken root and artificial alizarin: Textile arts and manufactures of madder. History and Technology, 12(1):1–22, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Marsch:1994:SSR

[265] Ulrich Marsch. Strategies for success: Research organization in German chemical companies and IG Farben until 1936. History and Technology, 12(1):23–77, 1994. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schroeder-Gudehus:1995:F

[266] Brigitte Schroeder-Gudehus. Foreword. History and Technology, 12(2):79, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Karvar:1995:MRD

[267] Anousheh Karvar. Model reception in the domain of engineering education: Mediation and negotiation. History and Technology, 12(2):81–93, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Leboutte:1995:TKH

[268] René Leboutte. From traditional know-how to technical skill. The process of training and of professionalization in the Belgian coal-mining industry, 1700–1850. History and Technology, 12(2):95–108, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gouzevitch:1995:THE

[269] I. Gouzévitch. Technical higher education in nineteenth century Russia and France: Some thoughts on a historical choice. History and Technology, 12(2):109–117, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fukasaku:1995:FID

[270] Yukiko Fukasaku. Foreign influence in the development of shipbuilding technology and the education of engineers in Japan, 1855–1940. History and Technology, 12(2):119–127, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guedon:1995:EAD

[271] Jean-Claude Guédon. Engineering by alienated design: The case of French-speaking engineers in Québec. History and Technology, 12(2): 129–146, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Macleod:1995:CEC

[272] Roy Macleod. Colonial engineers and the “cult of practicality”: Themes and dimensions in the history of Australian engineering. History and Technology, 12(2):147–162, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Dreicer:1995:IIE

[273] Gregory K. Dreicer. Influence and intercultural exchange: Engineers, engineering schools and engineering works in the nineteenth century. History and Technology, 12(2):163–177, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kahlow:1995:FID

[274] Andreas Kahlow. French influence on the development of applied mechanics in Germany in the nineteenth century. History and Technology, 12(2): 179–189, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Petitjean:1995:SDE

[275] Patrick Petitjean. Scientific development, engineering schools and the building of a modern state. History and Technology, 12(2):191–204, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hounshell:1995:HHT

[276] David A. Hounshell. Hughesian history of technology and Chandlerian business history: Parallels, departures, and critics. History and Technology, 12(3):205–224, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Nahum:1995:INF

[277] Andrew Nahum. The imitation of nature? Flapping flight and animal forms as models for human flight. History and Technology, 12(3):225–260, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Steen:1995:CCA

[278] Kathryn Steen. Confiscated commerce: American importers of German synthetic organic chemicals, 1914–1929. History and Technology, 12(3): 261–284, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1995:E

[279] Anonymous. Erratum. History and Technology, 12(3):285, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jacq:1995:EFR

[280] François Jacq. The emergence of French research policy: Methodological and historiographical problems (1945–1970). History and Technology, 12 (4):285–308, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Davis:1995:NAU

[281] Donald F. Davis. North American urban mass transit, 1890–1950: What if we thought about it as a type of technology? History and Technology, 12 (4):309–326, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lagree:1995:RTI

[282] Michel Lagrée. Religion and technological innovation: The steamboat in 1840s France. History and Technology, 12(4):327–359, 1995. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ceruzzi:1996:SIE

[283] Paul Ceruzzi. From scientific instrument to everyday appliance: The emergence of personal computers, 1970–77. History and Technology, 13(1): 1–31, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Homburg:1996:VPS

[284] Ernst Homburg and Johan H. De Vlieger. A victory of practice over science: The unsuccessful modernisation of the Dutch white lead industry (1780–1865). History and Technology, 13(1):33–52, 1996. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Neufeld:1996:REV

[285] Michael J. Neufeld. Rolf Engel vs. The German Army: a Nazi career in rocketry and repression. History and Technology, 13(1):53–72, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mandelbrote:1996:ER

[286] Scott Mandelbrote. Essay review. History and Technology, 13(1):73–81, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1996:EBa

[287] Anonymous. Editorial board. History and Technology, 13(1):ebi, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Galambos:1996:PIT

[288] Louis Galambos and Jeffrey L. Sturchio. The pharmaceutical industry in the twentieth century: A reappraisal of the sources of innovation. History and Technology, 13(2):83–100, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Amdam:1996:PNI

[289] Rolv Petter Amdam. Professional networks and the introduction of research in the British and Norwegian pharmaceutical industry in the interwar years. History and Technology, 13(2):101–114, 1996. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sogner:1996:PIB

[290] Knut Sogner. A pharmaceutical innovation. Business environment and scientific endeavour in Nyegaard & Co. History and Technology, 13(2): 115–131, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slinn:1996:IGM

[291] Judy Slinn. Innovation at Glaxo and May & Baker, 1945–1965. History and Technology, 13(2):133–147, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ceruzzi:1996:BR

[292] Paul Ceruzzi, Nathalie Jas, Rainer Karlsch, and Anthony S. Travis. Book reviews. History and Technology, 13(2):149–155, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1996:EBb

[293] Anonymous. Editorial board. History and Technology, 13(2):ebi, 1996. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Blondel:1997:EIC

[294] Christine Blondel. Electrical instruments in 19th century France, between makers and users. History and Technology, 13(3):157–182, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pestre:1997:SIF

[295] Dominique Pestre. Studies of the ionosphere and forecasts for radiocommunications. Physicists and engineers, the military and national laboratories in France (and Germany) after 1945. History and Technology, 13 (3):183–205, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wicken:1997:SST

[296] Olav Wicken. Space science and technology in the Cold War: The ionosphere, the military and politics in Norway. History and Technology, 13 (3):207–229, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fox:1997:BR

[297] Robert Fox, Ilana Löwy, Michael Mende, and Lorenta Sebesta. Book reviews. History and Technology, 13(3):231–240, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1997:EBa

Pestre:1997:MPE

[299] Dominique Pestre. The moral and political economy of French scientists in the first half of the XXth century. History and Technology, 13(4):241–248, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Boudia:1997:CLR

[300] Soraya Boudia. The Curie Laboratory: Radioactivity and metrology. History and Technology, 13(4):249–265, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Roque:1997:MCR

[301] Xavier Roqué. Marie Curie and the radium industry: A preliminary sketch. History and Technology, 13(4):267–291, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Vincent:1997:GPP

[302] Bénédicte Vincent. Genesis of the Pavillon Pasteur of the Institut du Radium of Paris. History and Technology, 13(4):293–305, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pinault:1997:JCS

[303] Michel Pinault. The Joliot-Curies: Science, politics, networks. History and Technology, 13(4):307–324, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hughes:1997:FCJ

[304] Jeff Hughes. The French connection: The Joliot-Curies and nuclear research in Paris, 1925–1933. History and Technology, 13(4):325–343, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Neufeld:1997:BR

[305] Michael J. Neufeld, Keith Pavitt, Antoine Picon, and Ulrich Wengenroth. Book reviews. History and Technology, 13(4):345–353, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1997:NC

[306] Anonymous. Notes on contributors. History and Technology, 13(4):355, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1997:EBb

Canel:1997:GAC

[308] Annie Canel and Karin Zachmann. Gaining access, crossing boundaries: Women in engineering in a comparative perspective. History and Technology, 14(1–2):1–5, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Berner:1997:EEW

[309] Boel Berner. Explaining exclusion: Women and Swedish engineering education from the 1890s to the 1920s. History and Technology, 14(1–2):7–29, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mikoletzky:1997:UCW

[310] Juliane Mikoletzky. An unintended consequence: Women’s entry into engineering education in Austria. History and Technology, 14(1–2):31–48, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fuchs:1997:LFL

[311] Margot Fuchs. Like fathers–like daughters professionalization strategies of women students and engineers in Germany 1890s to 1940s. History and Technology, 14(1–2):49–64, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Oldenziel:1997:DSW

[312] Ruth Oldenziel. Decoding the silence: Women engineers and male culture in the U.S., 1878–1951. History and Technology, 14(1–2):65–95, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zachmann:1997:WRF

[313] Karin Zachmann. Women to replace the ‘front officers of technology’? On the development of technical studies for women in the Soviet Occupation Zone and the GDR (1946–1971). History and Technology, 14(1–2):97–122, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Canel:1997:FTP

[314] Annie Canel. Following the trail of the pioneers: Women engineers in the French ‘grandes écoles’ in the 1960s and 1970s. History and Technology, 14(1–2):123–145, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schmidt:1997:FGM

[315] Dorothea Schmidt. Female guests in a manly world: A picture with numerous shades. History and Technology, 14(1–2):147–157, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Konig:1997:BR

[316] Wolfgang König, William Walker, Robert W. Seidel, and Gerhard Mener. Book reviews. History and Technology, 14(1–2):159–171, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1997:EBc

[317] Anonymous. Editorial board. History and Technology, 14(1–2):ebi, 1997. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schot:1998:UEM

[318] Johan Schot. The usefulness of evolutionary models for explaining innovation. The case of The Netherlands in the nineteenth century. History and Technology, 14(3):173–200, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

VanLente:1998:IPM

[319] Dick Van Lente. Innovation in paper making: The Netherlands 1750– 1850. History and Technology, 14(3):201–224, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Davids:1998:SFT

[320] Karel Davids. Successful and failed transitions. A comparison of innovations in windmill-technology in Britain and The Netherlands in the early modern period. History and Technology, 14(3):225–247, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:1998:BRa

[321] Jennifer Alexander, Mikael Hård, Robert Fox, John Krige, Xavier Roqué, and A. J. Turner. Book reviews. History and Technology, 14(3):249–264, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1998:NCa

[322] Anonymous. Notes on contributors. History and Technology, 14(3):265, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1998:EBa

[323] Anonymous. Editorial board. History and Technology, 14(3):ebi, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gooday:1998:RWB

[324] Graeme Gooday. Re-writing the ‘book of blots’: Critical reflections on histories of technological ‘failure’. History and Technology, 14(4):265–291, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ringrose:1998:WSP

[325] Daniel Ringrose. Work and social presence: French public engineers in nineteenth-century provincial communities. History and Technology, 14 (4):293–312, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Berner:1998:MCC

[326] Boel Berner. The meaning of cleaning: The creation of harmony and hygiene in the home. History and Technology, 14(4):313–352, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1998:NCb

[327] Anonymous. Notes on contributors. History and Technology, 14(4):353, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1998:EBb

[328] Anonymous. Editorial board. History and Technology, 14(4):ebi, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Agar:1998:IHC

[329] Jon Agar. Introduction: History of computing: Approaches, new directions and the possibility of informatic history. History and Technology, 15(1–2):1–5, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Edwards:1998:YMR

[330] Paul N. Edwards. Y2K: Millennial reflections on computers as infrastructure. History and Technology, 15(1–2):7–29, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Seidel:1998:CNC

[331] Robert W. Seidel. “crunching numbers” computers and physical research in the AEC laboratories. History and Technology, 15(1–2):31–68, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bowker:1998:ATH

[332] Geoffrey C. Bowker. Archival technology in the historical sciences 1800– 1997. History and Technology, 15(1–2):69–87, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Atkinson:1998:CMH

[333] Paul Atkinson. Computer memories: The history of computer form. History and Technology, 15(1–2):89–120, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Agar:1998:DPT

[334] Jon Agar. Digital patina: Texts, spirit and the first computer. History and Technology, 15(1–2):121–135, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:1998:BRb

[335] Jennifer Alexander, Michel Atten, Frank A. J. L. James, J. Krige, Xavier Roqué, Boel Berner, Luca Guzzetti, K. Austin Kerr, Barbara A. Kimmelmann, and Christiane Sinding. Book reviews. History and Technology, 15(1–2):137–173, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1998:NCc

[336] Anonymous. Notes on contributors. History and Technology, 15(1–2): 174–175, 1998. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:1999:LBP

[337] Jennifer Karns Alexander. The line between potential and working machines: César Nicolas Leblanc and patent engravings, 1811–1835. History and Technology, 15(3):175–212, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jamesa:1999:SPM

[338] Frank A. J. L. Jamesa and Margaret Rayb. Science in the pits: Michael Faraday, Charles Lyell and the Home Office enquiry into the explosion at Haswell Colliery, County Durham, in 1844. History and Technology, 15 (3):213–231, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Chatzis:1999:SSF

[339] Konstantinos Chatzis. Searching for standards: French engineers and time and motion studies of industrial operations in the 1950s. History and Technology, 15(3):233–261, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Marks:1999:HGP

[340] Lara Marks. Human guinea pigs? The history of the early oral contraceptive clinical trials. History and Technology, 15(4):263–288, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hong:1999:HLR

[341] Sungook Hong. Historiographical layers in the relationship between science and technology. History and Technology, 15(4):289–311, 1999. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gardey:1999:STP

[342] Delphine Gardey. The standardization of a technical practice: Typing (1883–1930). History and Technology, 15(4):313–343, 1999. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Berner:1999:WDB

[343] Boel Berner. The worker’s dream of becoming an engineer. History and Technology, 15(4):345–372, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Santesmases:1999:BR

[344] María Jesús Santesmases, Agnes Tandler, Ilana Löwy, Christophe Bonneuil, and Florian Hars. Book reviews. History and Technology, 15(4): 373–389, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Johnson:1999:BAB

[345] Stephen B. Johnson. Building an American bridge over the “management gap”: The adoption of systems management in ESRO and ESA. History and Technology, 16(1):1–32, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Tulkki:1999:TTE

[346] Pasi Tulkki. Two types of engineers in a slowly industrialising Finland. History and Technology, 16(1):33–66, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bulkeley:1999:HSA

[347] Rip Bulkeley. Harbingers of Sputnik: The amateur radio preparations in the Soviet Union. History and Technology, 16(1):67–102, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:1999:BR

[348] Jennifer K. Alexander, David De Vorkin, and Phillip Stevens Thurtle. Book reviews. History and Technology, 16(1):103–110, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1999:NCa

[349] Anonymous. Notes on contributors. History and Technology, 16(1):111, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1999:EBa

[350] Anonymous. Editorial board. History and Technology, 16(1):ebi, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Edgerton:1999:IUT

[351] David Edgerton. From innovation to use: Ten eclectic theses on the historiography of technology. History and Technology, 16(2):111–136, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Russo:1999:SSV

[352] Arturo Russo. Science in space vs space science: The European utilisation of Spacelab. History and Technology, 16(2):137–177, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lecuyer:1999:SIC

[353] Christophe Lécuyer. Silicon for industry: Component design, mass production, and the move to commercial markets at Fairchild Semiconductor, 1960–1967. History and Technology, 16(2):179–216, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1999:NCb

[354] Anonymous. Notes on contributors. History and Technology, 16(2):217, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:1999:EBb

[355] Anonymous. Editorial board. History and Technology, 16(2):ebi, 1999. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Belfanti:2000:IIT

[356] Carlo Marco Belfanti and Fabio Giusberti. Introduction: Institutions and technical change in early Modern Europe. History and Technology, 16 (3):217–222, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Long:2000:IST

[357] Pamela O. Long. Invention, secrecy, and theft: Meaning and context in the study of late medieval technical transmission. History and Technology, 16 (3):223–241, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Heller:2000:PAT

[358] Henry Heller. Primitive accumulation and technical innovation in the French wars of religion. History and Technology, 16(3):243–262, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Davids:2000:PPD

[359] Karel Davids. Patents and patentees in the Dutch Republic between c. 1580 and 1720. History and Technology, 16(3):263–283, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hilaire-Perez:2000:TII

[360] Liliane Hilaire-Pérez. Technical invention and institutional credit in France and Britain in the 18th century. History and Technology, 16(3): 285–306, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:NCa

[361] Anonymous. Notes on contributors. History and Technology, 16(3):307, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:EBa

Macleod:2000:MWR

[363] Christine Macleod, Jeremy Stein, Jennifer Tann, and James Andrew. Making waves: The Royal Navy’s management of invention and innovation in steam shipping, 1815–1832. History and Technology, 16(4):307–333, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Serchinger:2000:WPU

[364] Reinhard W. Serchinger. Wirtschaftswunder in Pretzfeld, Upper Franconia: Interactions between science, technology, and corporate strategies in Siemens semiconductor rectifier research & development, 1945–1956. History and Technology, 16(4):335–381, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schroter:2000:SRA

[365] Harm G. Schröter. Strategic R&D as an answer to the oil crisis. West and East German investment in coal refinement and the chemical industries, 1970–1990. History and Technology, 16(4):383–402, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Steele:2000:BR

[366] Brett D. Steele, Moran Tompkins, and J. Krige. Book reviews. History and Technology, 16(4):403–412, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:NCb

[367] Anonymous. Notes on contributors. History and Technology, 16(4):413– 414, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:EBb

[368] Anonymous. Editorial board. History and Technology, 16(4):ebi, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:E

[369] Anonymous. Editorial. History and Technology, 17(1):i–ii, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Eldridge:2000:EEA

[370] Chris Eldridge. Electronic eyes for the Allies: Anglo–American cooperation on radar development during World War II. History and Technology, 17(1):1–20, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Benson:2000:SAW

[371] Erik Benson. Suspicious allies: Wartime aviation developments and the Anglo–American international airline rivalry, 1939–45. History and Technology, 17(1):21–42, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Engel:2000:WCW

[372] Jeffrey A. Engel. “We are not concerned who the buyer is”: Engine sales and Anglo–American security at the dawn of the jet age. History and Technology, 17(1):43–67, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:2000:BRa

[373] Jennifer Alexander, Jon Agar, and Jon Agar. Book reviews. History and Technology, 17(1):69–77, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:NCc

[374] Anonymous. Notes on contributors. History and Technology, 17(1):78, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:EBc

[375] Anonymous. Editorial board. History and Technology, 17(1):ebi, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Raj:2000:CPV

[376] Kapil Raj. 18th-century Pacific voyages of discovery, “big science”, and the shaping of an European scientific and technological culture. History and Technology, 17(2):79–98, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zavidonov:2000:SEP

[377] Ivan Vladimirovich Zavidonov. Sputniks, explorers and propaganda: The discovery of the Earth’s radiation belts. History and Technology, 17(2): 99–124, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jones-Imhotep:2000:DTE

[378] Edward Jones-Imhotep. Disciplining technology: Electronic reliability, Cold-War military culture and the topside ionogram. History and Technology, 17(2):125–175, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:2000:BRb

[379] Jennifer K. Alexander and James Satter. Book reviews. History and Technology, 17(2):177–181, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:NCd

[380] Anonymous. Notes on contributors. History and Technology, 17(2):182, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2000:EBd

[381] Anonymous. Editorial board. History and Technology, 17(2):ebi, 2000. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kraemer:2001:FIP

[382] Sylvia Katharine Kraemer. Federal intellectual property policy and the history of technology: The case of NASA patents. History and Technology, 17(3):183–216, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Rozwadowski:2001:TOS

[383] Helen M. Rozwadowski. Technology and ocean-scape: Defining the deep sea in mid-nineteenth century. History and Technology, 17(3):217–247, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Johnson:2001:URS

[384] Ann Johnson. Unpacking reliability: The success of Robert Bosch, GmbH in constructing antilock braking systems as reliable products. History and Technology, 17(3):249–270, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Paton-Walsh:2001:BR

[385] Margaret Paton-Walsh and Jonathan Zeitlin. Book reviews. History and Technology, 17(3):271–278, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2001:NCa

[386] Anonymous. Notes on contributors. History and Technology, 17(3):279, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2001:EBa

Jones:2001:TCS

[388] David S. Jones. Technologies of compliance: Surveillance of selfadministration of tuberculosis treatment, 1956–1966. History and Technology, 17(4):279–318, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gardey:2001:MWP

[389] Delphine Gardey. Mechanizing writing and photographing the word: Utopias, office work, and histories of gender and technology. History and Technology, 17(4):319–352, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2001:GWC

[390] Amy Slaton. George Washington Carver slept here: Racial identity and laboratory practice at Iowa State College. History and Technology, 17 (4):353–374, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:2001:BR

[391] Jennifer Karns Alexander, Richard Bellon, Rachel Prentice, J. B. Shank, and Audra J. Wolfe. Book reviews. History and Technology, 17(4):375– 392, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2001:NCb

[392] Anonymous. Notes on contributors. History and Technology, 17(4):393, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2001:EBb

[393] Anonymous. Editorial board. History and Technology, 17(4):ebi, 2001. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lindner:2002:TTG

[394] Stephan H. Lindner. Technology and textiles globalization. History and Technology, 18(1):1–22, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ploeger:2002:ASF

[395] Joanna S. Ploeger. The art of science at Fermi National Accelerator Laboratory: The rhetoric of aesthetics and humanism in the national laboratory system in the late 1960s. History and Technology, 18(1):23–49, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Tone:2002:MRR

[396] Andrea Tone. Making room for rubbers: Gender, technology, and birth control before the pill. History and Technology, 18(1):51–76, 2002. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Calvo:2002:STS

[397] Angel Calvo. The Spanish telephone sector (1876–1924): A case of technological backwardness. History and Technology, 18(2):77–102, 2002. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fletcher:2002:FEI

[398] Amy L. Fletcher. France enters the information age: A political history of Minitel. History and Technology, 18(2):103–119, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hohler:2002:DRO

[399] Sabine Höhler. Depth records and ocean volumes: Ocean profiling by sounding technology, 1850–1930. History and Technology, 18(2):119–154, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Berner:2002:HFM

[400] Boel Berner. “Housewives’ films” and the modern housewife. Experts, users and household modernization: Sweden in the 1950s and 1960s. History and Technology, 18(3):155–179, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Neal:2002:BSF

[401] Valerie Neal. Bumped from the shuttle fleet: Why didn’t enterprise fly in space? History and Technology, 18(3):181–202, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mertens:2002:TSI

[402] Joost Mertens. Technology as the science of the industrial arts: Louis-Sébastien Lenormand (1757–1837) and the popularization of technology. History and Technology, 18(3):203–231, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Carson:2002:NED

[403] Cathryn Carson. Nuclear energy development in postwar West Germany: Struggles over cooperation in the Federal Republic’s first reactor station. History and Technology, 18(3):233–270, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Tone:2002:I

[404] Andrea Tone. Introduction. History and Technology, 18(4):271–276, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Tone:2002:HMM

[405] John Lawrence Tone. How the mosquito (man) liberated Cuba. History and Technology, 18(4):277–308, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See letter [425].

Klawiter:2002:RPB

[406] Maren Klawiter. Risk, prevention and the breast cancer continuum: The NCI, the FDA, health activism and the pharmaceutical industry. History and Technology, 18(4):309–353, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Rosser:2002:OWH

[407] Sue V. Rosser. An overview of women’s health in the U.S. since the mid-1960’s. History and Technology, 18(4):355–369, 2002. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hecht:2003:GMF

[408] Gabrielle Hecht. Globalization meets Frankenstein? Reflections on terrorism, nuclearity, and global technopolitical discourse. History and Technology, 19(1):1–8, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Knowles:2003:LRW

[409] Scott Gabriel Knowles. Lessons in the rubble: The world trade center and the history of disaster investigations in the United States. History and Technology, 19(1):9–28, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Levin:2003:FRT

[410] Miriam Levin and Rosalind Williams. Forum on rethinking technology in the aftermath of September 11. History and Technology, 19(1):29–83, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2003:NCa

[411] Anonymous. Notes on contributors. History and Technology, 19(1):85–86, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Tassava:2003:WSC

[412] Christopher James Tassava. Weak seams: Controversy over welding theory and practice in American shipyards, 1938–1946. History and Technology, 19(2):87–108, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Boersma:2003:SWE

[413] F. Kees Boersma. Structural ways to embed a research laboratory into the company: A comparison between Philips and General Electric 1900–1940. History and Technology, 19(2):109–126, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Macdonald:2003:WDS

[414] Peter Macdonald and Michael Clow. What a difference a skidder makes: The role of technology in the origins of the industrialization of tree harvesting systems. History and Technology, 19(2):127–149, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Twigge:2003:BET

[415] Stephen Twigge. A baffling experience: Technology transfer, Anglo– American nuclear relations, and the development of the gas centrifuge 1964–70*. History and Technology, 19(2):151–163, 2003. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2003:BRa

[416] Anonymous. Book reviews. History and Technology, 19(2):165–172, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Low:2003:I

[417] Morris Low and Robert Kargon. Introduction. History and Technology, 19 (3):175–176, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Forgan:2003:AW

[418] Sophie Forgan. Atoms in Wonderland. History and Technology, 19(3): 177–196, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Low:2003:DFT

[419] Morris Low. Displaying the future: techno-nationalism and the rise of the consumer in Postwar Japan. History and Technology, 19(3):197–209, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Molella:2003:EAC

[420] Arthur Molella. Exhibiting atomic culture: the view from Oak Ridge. History and Technology, 19(3):211–226, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Beyler:2003:DTM

[421] Richard Beyler. The demon of technology, mass society, and atomic physics in West Germany, 1945–1957. History and Technology, 19(3): 227–239, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Badash:2003:SBS

[422] Lawrence Badash. From security blanket to security risk: scientists in the decade after Hiroshima. History and Technology, 19(3):241–256, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hughes:2003:SRB

[423] Jeff Hughes. The Strath Report: Britain confronts the H-bomb, 1954– 1955. History and Technology, 19(3):257–275, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Josephson:2003:TUT

[424] Paul Josephson. Technological utopianism in the twenty-first century: Russia’s nuclear future. History and Technology, 19(3):277–292, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2003:CMM

[425] Anonymous. Correspondence: the mosquito man in Cuba. History and Technology, 19(3):293–298, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See [405].

Anonymous:2003:BRb

[426] Anonymous. Book reviews. History and Technology, 19(3):299–309, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2003:NCb

[427] Anonymous. Notes on contributors. History and Technology, 19(3):311– 312, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wisnioski:2003:ISE

[428] Matt Wisnioski. Inside “the system”: engineers, scientists, and the boundaries of social protest in the long 1960s. History and Technology, 19(4): 313–333, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slayton:2003:SSC

[429] Rebecca Slayton. Speaking as scientists: computer professionals in the star wars debate. History and Technology, 19(4):335–364, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Shulman:2003:SCN

[430] Peter Shulman. “Science can never demobilize”: the United States Navy and petroleum geology, 1898–1924. History and Technology, 19(4):365– 385, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2003:BRc

[431] Anonymous. Book reviews. History and Technology, 19(4):387–391, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2003:NCc

[432] Anonymous. Notes on contributors. History and Technology, 19(4):393, 2003. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Widmalm:2004:SBB

[433] Sven Widmalm. The Svedberg and the boundary between science and industry: Laboratory practice, policy, and media images. History and Technology, 20(1):1–27, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Johnston:2004:TTG

[434] Sean F. Johnston. Telling tales: George stroke and the historiography of holography. History and Technology, 20(1):29–51, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Choi:2004:RGS

[435] Hyungsub Choi. Rationalizing the guerilla state: North Korean factory management reform, 1953–61. History and Technology, 20(1):53–74, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mom:2004:IAT

[436] Gijs Mom. Inter-artifactual technology transfer: Road building technology in The Netherlands and the competition between bricks, macadam, asphalt and concrete. History and Technology, 20(1):75–96, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Siddiqi:2004:DIR

[437] Asif Siddiqi. Deep impact: Robert Goddard and the Soviet ‘space fad’ of the 1920s. History and Technology, 20(2):97–113, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Conway:2004:EGC

[438] Erik M. Conway. Echoes in the grand canyon: Public catastrophes and technologies of control in American aviation. History and Technology, 20 (2):115–134, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mertens:2004:AIT

[439] Joost Mertens. The Annales de l’Industrie (1820–1827): A technological laboratory for the industrial modernization of France. History and Technology, 20(2):135–163, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Evan:2004:VTP

[440] William M. Evan. Voting technology, political institutions, legal institutions and civil society: A study of the hypothesis of cultural lag in reverse. History and Technology, 20(2):165–183, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2004:R

[441] Anonymous. Reviews. History and Technology, 20(2):185–193, 2004. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vanderVleuten:2004:INT

[442] Erik van der Vleuten. Introduction: networking technology, networking society, networking nature. History and Technology, 20(3):195–203, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Verbong:2004:UCM

[443] Geert Verbong and Erik van der Vleuten. Under construction: material integration of The Netherlands 1800–2000. History and Technology, 20 (3):205–226, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hermans:2004:BBS

[444] Janneke Hermans and Onno de Wit. Bourses and brokers: stock exchanges as ICT junctions. History and Technology, 20(3):227–247, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

delaBruheze:2004:MWI

[445] Adri A. Albert de la Bruhèze and Anneke H. van Otterloo. The Milky Way: infrastructures and the shaping of milk chains. History and Technology, 20(3):249–269, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Davids:2004:FPP

[446] Mila Davids. The fabric of production: the Philips industrial network. History and Technology, 20(3):271–290, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vanderVleuten:2004:WWR

[447] Erik van der Vleuten and Cornelis Disco. Water wizards: reshaping wet nature and society. History and Technology, 20(3):291–309, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vandenBelt:2004:NNS

[448] Henk van den Belt. Networking nature, or Serengeti behind the dikes. History and Technology, 20(3):311–333, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Assimakopoulos:2004:IWS

[449] Michalis Assimakopoulos. Introduction: Why have an STS meeting in Greece? History and Technology, 20(4):335–337, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Woolgar:2004:WHP

[450] Steve Woolgar. What happened to provocation in science and technology studies? History and Technology, 20(4):339–349, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pestre:2004:TYS

[451] Dominique Pestre. Thirty years of science studies: knowledge, society and the political. History and Technology, 20(4):351–369, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bijker:2004:SPP

[452] Wiebe E. Bijker. Sustainable policy? A public debate about nature development in The Netherlands. History and Technology, 20(4):371–391, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Downey:2004:KPI

[453] Gary Lee Downey and Juan C. Lucena. Knowledge and professional identity in engineering: code-switching and the metrics of progress. History and Technology, 20(4):393–420, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Picon:2004:EEH

[454] Antoine Picon. Engineers and engineering history: problems and perspectives. History and Technology, 20(4):421–436, 2004. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Misa:2005:IIE

[455] Thomas J. Misa and Johan Schot. Introduction inventing Europe: Technology and the hidden integration of Europe. History and Technology, 21(1):1–19, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vanderVleuten:2005:NE

[456] Erik van der Vleuten and Arne Kaijser. Networking Europe. History and Technology, 21(1):21–48, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Trischler:2005:EEB

[457] Helmuth Trischler and Hans Weinberger. Engineering Europe: big technologies and military systems in the making of 20th century Europe. History and Technology, 21(1):49–83, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Arnold:2005:ETC

[458] David Arnold. Europe, technology, and colonialism in the 20th century. History and Technology, 21(1):85–106, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Oldenziel:2005:EMJ

[459] Ruth Oldenziel, Adri Albert de la Bruhèze, and Onno de Wit. Europe’s mediation junction: technology and consumer society in the 20th century.

History and Technology, 21(1):107–139, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2005:Ra

[460] Anonymous. Reviews. History and Technology, 21(1):141–148, 2005. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ferguson:2005:OTP

[461] Robert G. Ferguson. One thousand planes a day: Ford, Grumman, General Motors and the Arsenal of Democracy. History and Technology, 21 (2):149–175, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

McCray:2005:WSB

[462] W. Patrick McCray. Will small be beautiful? Making policies for our nanotech future. History and Technology, 21(2):177–203, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Brandstetter:2005:MWP

[463] Thomas Brandstetter. “The most wonderful piece of machinery the world can boast of”: The water-works at Marly, 1680–1830. History and Technology, 21(2):205–220, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Buchanan:2005:AMC

[464] Nicholas Buchanan. The atomic meal: The Cold War and irradiated foods, 1945–1963. History and Technology, 21(2):221–249, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2005:RDI

[465] Anonymous. Recent dissertations of interest to historians of technology. History and Technology, 21(2):251–253, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hacker:2005:MWW

[466] Barton C. Hacker. The machines of war: Western military technology 1850–2000. History and Technology, 21(3):255–300, 2005. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2005:OWO

[467] Martin Collins. One world … one telephone: Iridium, one look at the making of a global age. History and Technology, 21(3):301–324, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2005:GEG

[468] Anonymous. A gramophone in every grave. History and Technology, 21 (3):325–329, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kwa:2005:IPE

[469] Chunglin Kwa. Interdisciplinarity and postmodernity in the environmental sciences. History and Technology, 21(4):331–344, 2005. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vanRooij:2005:ECC

[470] Arjan van Rooij. Engineering contractors in the chemical industry. The development of ammonia processes, 1910–1940. History and Technology, 21(4):345–366, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Johnston:2005:AST

[471] Sean F. Johnston. Attributing scientific and technological progress: The case of holography. History and Technology, 21(4):367–392, 2005. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2005:Rb

[472] Anonymous. Reviews. History and Technology, 21(4):393–401, 2005. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Coopersmith:2006:DYM

[473] Jonathan Coopersmith. Does your mother know what you Really do? The changing nature and image of computer-based pornography. History and Technology, 22(1):1–25, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slade:2006:ETR

[474] Joseph W. Slade. Eroticism and technological regression: The stag film. History and Technology, 22(1):27–52, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Harwood:2006:EEB

[475] Jonathan Harwood. Engineering education between science and practice: Rethinking the historiography. History and Technology, 22(1):53–79, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Weiss:2006:STW

[476] Charles Weiss. Science and technology at the World Bank, 1968–83. History and Technology, 22(1):81–104, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Evan:2006:DUT

[477] William M. Evan and Bret B. Hays. Dual-use technology in the context of the non proliferation regime. History and Technology, 22(1):105–113, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Travis:2006:P

[478] Anthony S. Travis. Preface. History and Technology, 22(2):115–118, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Morris:2006:DSM

[479] Peter J. T. Morris. Does the Science Museum, London, have Perkin’s original mauve dye? A critical reassessment of a chemical icon. History and Technology, 22(2):119–130, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Travis:2006:DDC

[480] Anthony S. Travis. Decadence, decline and celebration: Raphael Meldola and the Mauve Jubilee of 1906. History and Technology, 22(2):131–152, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

OReilly:2006:VGC

[481] Declan O’Reilly. Vesting GAF Corporation: The Roosevelt Administration’s decision to Americanise I. G. Farben’s American affiliates in World War II. History and Technology, 22(2):153–186, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Baptista:2006:GFA

[482] Robert J. Baptista and Anthony S. Travis. I. G. Farben in America: The technologies of General Aniline & Film. History and Technology, 22(2): 187–224, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Launius:2006:IML

[483] Roger D. Launius. Interpreting the Moon landings: Project Apollo and the historians. History and Technology, 22(3):225–255, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wormbs:2006:NSP

[484] Nina Wormbs. A Nordic satellite project understood as a trans-national effort. History and Technology, 22(3):257–275, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Uziel:2006:BIR

[485] Daniel Uziel. Between industrial revolution and slavery. Mass production in the German aviation industry in World War II. History and Technology, 22(3):277–300, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lecuyer:2006:MM

[486] Christophe Lécuyer and David C. Brock. The materiality of microelectronics. History and Technology, 22(3):301–325, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2006:RDI

[487] Anonymous. Recent dissertations of interest to historians of technology. History and Technology, 22(3):327–330, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2006:R

[488] Anonymous. Reviews. History and Technology, 22(3):331–335, 2006. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Scranton:2006:TLI

[489] Philip Scranton. Technology-led innovation: The non-linearity of US jet propulsion development. History and Technology, 22(4):337–367, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Grob:2006:WED

[490] Bart Grob. Willem Einthoven and the development of the string galvanometer. How an instrument escaped the laboratory. History and Technology, 22(4):369–390, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:2006:ESU

[491] Jennifer K. Alexander. An efficiency of scarcity: Using food to increase the productivity of Soviet prisoners of war in the mines of the Third Reich. History and Technology, 22(4):391–406, 2006. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Forman:2007:PSM

[492] Paul Forman. The primacy of science in modernity, of technology in postmodernity, and of ideology in the history of technology. History and Technology, 23(1–2):1–152, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2007:R

[493] Martin Collins, Ronald Kline, Chunglin Kwa, and Philip Mirowski. Responses. History and Technology, 23(1–2):153–184, 2007. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Yates:2007:R

[494] Thomas Yates, Greg Downey, and Steven W. Usselman. Reviews. History and Technology, 23(1–2):185–192, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Chatzis:2007:INI

[495] Konstantinos Chatzis. Introduction: The national identities of engineers. History and Technology, 23(3):193–196, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Picon:2007:FES

[496] Antoine Picon. French engineers and social thought, 18–20th centuries: An archeology of technocratic ideals. History and Technology, 23(3):197– 208, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Belhoste:2007:TCT

[497] Bruno Belhoste and Konstantinos Chatzis. From technical corps to technocratic power: French state engineers and their professional and cultural universe in the first half of the 19th century. History and Technology, 23 (3):209–225, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bocquet:2007:ENI

[498] Denis Bocquet. Engineers and the nation in Italy (1750–1922): Local traditions and different conceptions of unity and modernity. History and Technology, 23(3):227–240, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Antoniou:2007:NII

[499] Yiannis Antoniou, Michalis Assimakopoulos, and Konstantinos Chatzis. The national identity of inter-war Greek engineers: Elitism, rationalization, technocracy, and reactionary modernism. History and Technology, 23 (3):241–261, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Saraiva:2007:ITN

[500] Tiago Saraiva. Inventing the technological nation: The example of Portugal (1851–1898). History and Technology, 23(3):263–273, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lucena:2007:CME

[501] Juan C. Lucena. De Criollos a Mexicanos: Engineers’ identity and the construction of Mexico. History and Technology, 23(3):275–288, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Downey:2007:LCM

[502] Gary Lee Downey. Low cost, mass use: American engineers and the metrics of progress. History and Technology, 23(3):289–308, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ashrafi:2007:R

[503] Babak Ashrafi, Kent M. James, and Jennifer L. Croissant. Reviews. History and Technology, 23(3):309–315, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Boudia:2007:IRR

[504] Soraya Boudia and Nathalie Jas. Introduction: Risk and ‘risk society’ in historical perspective. History and Technology, 23(4):317–331, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fressoz:2007:BBC

[505] Jean-Baptiste Fressoz. Beck back in the 19th century: Towards a genealogy of risk society. History and Technology, 23(4):333–350, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Dessaux:2007:CEF

[506] Pierre Antoine Dessaux. Chemical expertise and food market regulation in Belle-Epoque France. History and Technology, 23(4):351–368, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jas:2007:PHP

[507] Nathalie Jas. Public health and pesticide regulation in France before and after Silent Spring. History and Technology, 23(4):369–388, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Boudia:2007:GRC

[508] Soraya Boudia. Global regulation: Controlling and accepting radioactivity risks. History and Technology, 23(4):389–406, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pestre:2007:HHC

[509] Dominique Pestre. The historical heritage of the 19th and 20th centuries: Techno science, markets and regulations in a long term perspective. History and Technology, 23(4):407–420, 2007. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Boyle:2008:HSH

[510] Marjorie O’Rourke Boyle. Harvey in the sluice: from hydraulic engineering to human physiology. History and Technology, 24(1):1–22, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sarasua:2008:TIS

[511] Carmen Sarasúa. Technical innovations at the service of cheaper labor in pre-industrial Europe. The enlightened agenda to transform the gender division of labor in silk manufacturing. History and Technology, 24 (1):23–39, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Boomgaard:2008:TTE

[512] Peter Boomgaard. Technologies of a trading empire: Dutch introduction of water- and windmills in early-modern Asia, 1650s–1800. History and Technology, 24(1):41–59, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fallan:2008:FFF

[513] Kjetil Fallan. Form, function, fiction — translations of technology and design in product development. History and Technology, 24(1):61–87, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vanLente:2008:TOH

[514] Dick van Lente. Three overviews of the history of technology. History and Technology, 24(1):89–96, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gordin:2008:BR

[515] Michael D. Gordin. Book review. History and Technology, 24(1):97–98, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gaudilliere:2008:HPB

[516] Jean-Paul Gaudillière. How pharmaceuticals became patentable: the production and appropriation of drugs in the twentieth century. History and Technology, 24(2):99–106, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gaudilliere:2008:PIO

[517] Jean-Paul Gaudillière. Professional or industrial order? Patents, biological drugs, and pharmaceutical capitalism in early twentieth century Germany. History and Technology, 24(2):107–133, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cassier:2008:PPH

[518] Maurice Cassier. Patents and public health in France. Pharmaceutical patent law in-the-making at the Patent Office between the two world wars. History and Technology, 24(2):135–151, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cassier:2008:PPI

[519] Maurice Cassier and Christiane Sinding. “Patenting in the public interest”: administration of insulin patents by the University of Toronto. History and Technology, 24(2):153–171, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bud:2008:UME

[520] Robert Bud. Upheaval in the moral economy of science? Patenting, teamwork and the World War II experience of penicillin. History and Technology, 24(2):173–190, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slinn:2008:PUP

[521] Judy Slinn. Patents and the UK pharmaceutical industry between 1945 and the 1970s. History and Technology, 24(2):191–205, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hughes:2008:AVG

[522] Thomas P. Hughes. MIT architecture and values: Gehry’s Stata and Holl’s Simmons. History and Technology, 24(3):207–220, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lauwaert:2008:POB

[523] Maaike Lauwaert. Playing outside the box — on LEGO toys and the changing world of construction play. History and Technology, 24(3):221– 237, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lee:2008:TMI

[524] Francis Lee. Technopedagogies of mass-individualization: correspondence education in the mid-twentieth century. History and Technology, 24(3): 239–253, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Godin:2008:IEH

[525] Benoît Godin. The information economy: the history of a concept through its measurement, 1949–2005. History and Technology, 24(3):255–287, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sumner:2008:SSS

[526] James Sumner. Status, scale and secret ingredients: the retrospective invention of London porter. History and Technology, 24(3):289–306, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Heide:2008:PCP

[527] Lars Heide. Punched cards for professional European offices: revisiting the dynamics of information technology diffusion from the United States to Europe, 1889–1918. History and Technology, 24(4):307–320, 2008. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wilson:2008:MSF

[528] Robbie Guerriero Wilson. “The machine should fit the work”: Organisation and Method and British approaches to new technology in business.

History and Technology, 24(4):321–333, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gerovitch:2008:IWS

[529] Slava Gerovitch. InterNyet: why the Soviet Union did not build a nationwide computer network. History and Technology, 24(4):335–350, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Turchetti:2008:TIS

[530] Simone Turchetti, Simon Naylor, Katrina Dean, and Martin Siegert. On thick ice: scientific internationalism and Antarctic affairs, 1957–1980. History and Technology, 24(4):351–376, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vanderVleuten:2008:BR

[531] Erik van der Vleuten. Book review. History and Technology, 24(4):377– 378, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Post:2008:FEB

[532] Robert C. Post and John Krige. Forman: An exchange between Bob Post and John Krige. History and Technology, 24(4):379–381, 2008. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2009:NNE

[533] Martin Collins. A note from the new Editor. History and Technology, 25(1):1–2, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

VanVleck:2009:ACW

[534] Jenifer Van Vleck. An airline at the crossroads of the world: Ariana Afghan Airlines, modernization, and the global Cold War. History and Technology, 25(1):3–24, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hamblin:2009:LTL

[535] Jacob Darwin Hamblin. Let there be light … and bread: the United Nations, the developing world, and atomic energy’s Green Revolution. History and Technology, 25(1):25–48, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Eisler:2009:GPP

[536] Matthew N. Eisler. Getting power to the people: technological dramaturgy and the quest for the electrochemical engine. History and Technology, 25(1):49–68, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Yi:2009:SCM

[537] Doogab Yi. The scientific commons in the marketplace: the industrialization of biomedical materials at the New England Enzyme Center, 1963– 1980. History and Technology, 25(1):69–87, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kershaw:2009:DSH

[538] Michael Kershaw. “Diogenes in search of an honest man”: the genesis of the industrial inch, the first global standard of length. History and Technology, 25(2):89–114, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alcorn:2009:FMM

[539] Aaron L. Alcorn. Flying into modernity: model airplanes, consumer culture, and the making of modern boyhood in the early twentieth century. History and Technology, 25(2):115–146, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Royo:2009:SPI

[540] Jose Antonio Mateos Royo. State policy, institutional framework and technical monopoly in early modern Spain: invention patents in the Crown of Aragon during the seventeenth century. History and Technology, 25(2): 147–162, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2009:ENa

[541] Martin Collins. Editor’s note. History and Technology, 25(3):163–164, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lecuyer:2009:HTM

[542] Christophe Lécuyer and David C. Brock. High tech manufacturing. History and Technology, 25(3):165–171, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Holbrook:2009:CCO

[543] Daniel Holbrook. Controlling contamination: the origins of clean room technology. History and Technology, 25(3):173–191, 2009. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lecuyer:2009:NPS

[544] Christophe Lécuyer and David C. Brock. From nuclear physics to semiconductor manufacturing: the making of ion implantation. History and Technology, 25(3):193–217, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Yost:2009:MMC

[545] Jeffrey R. Yost. Manufacturing mainframes: component fabrication and component procurement at IBM and Sperry Univac, 1960–1975. History and Technology, 25(3):219–235, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Daemmrich:2009:SMC

[546] Arthur Daemmrich. Synthesis by microbes or chemists? Pharmaceutical research and manufacturing in the antibiotic era. History and Technology, 25(3):237–256, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Scranton:2009:SFM

[547] Philip Scranton. The shows and the flows: materials, markets, and innovation in the US machine tool industry, 1945–1965. History and Technology, 25(3):257–304, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2009:ENb

[548] Martin Collins. Editor’s note. History and Technology, 25(4):305, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Quivik:2009:ENS

[549] Fredric L. Quivik. Engineering nature: the Souris River and the production of migratory waterfowl. History and Technology, 25(4):307–323, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gardner:2009:TTP

[550] Robert Gardner. Trees as technology: planting shelterbelts on the Great Plains. History and Technology, 25(4):325–341, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jorgensen:2009:OWD

[551] Dolly Jørgensen. An oasis in a watery desert? Discourses on an industrial ecosystem in the Gulf of Mexico Rigs-to-Reefs program. History and Technology, 25(4):343–364, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Brookes:2009:TGB

[552] Barbara Brookes and Catherine Smith. Technology and gender: barbers and hairdressers in New Zealand, 1900–1970. History and Technology, 25 (4):365–386, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kessler:2009:IIT

[553] Elizabeth A. Kessler. Introducing ‘images, technology, and history’: a note from the image editor. History and Technology, 25(4):387–389, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Marcaida:2009:PTG

[554] José Ramón Marcaida. Portraying technology in gallery paintings. History and Technology, 25(4):391–397, 2009. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2010:ENa

[555] Martin Collins. Editor’s note. History and Technology, 26(1):1–2, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guse:2010:NTT

[556] John C. Guse. Nazi technical thought revisited. History and Technology, 26(1):3–33, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Herf:2010:CJH

[557] Jeffrey Herf. Comment by Jeffrey Herf. History and Technology, 26(1): 33–37, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guse:2010:JCG

[558] John Guse. John C. Guse replies. History and Technology, 26(1):38, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hohler:2010:ELS

[559] Sabine Höhler. The environment as a life support system: the case of Biosphere 2. History and Technology, 26(1):39–58, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2010:MCF

[560] Martin Collins. Matthew Connelly’s Fatal Misconception: The Struggle to Control World Population: Introduction. History and Technology, 26 (1):59–60, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Engerman:2010:RP

[561] David C. Engerman. Reproducing power? History and Technology, 26 (1):61–67, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Murphy:2010:TGP

[562] Michelle Murphy. Technology, governmentality, and population control. History and Technology, 26(1):69–76, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Rao:2010:IGH

[563] Anupama Rao. India and global history. History and Technology, 26 (1):77–84, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Connelly:2010:ARA

[564] Matthew Connelly. Author response: all biopolitics is global. History and Technology, 26(1):85–88, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Nielsen:2010:TLP

[565] Henry Nielsen and Henrik Knudsen. The troublesome life of peaceful atoms in Denmark. History and Technology, 26(2):91–118, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jeon:2010:FWM

[566] Chihyung Jeon. Flying weather men and robot observers: instruments, inscriptions, and identities in US upper-air observation, 1920–1940. History and Technology, 26(2):119–145, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2010:COV

[567] Martin Collins. Chris Otter’s The Victorian Eye: A Political History of Light and Vision in Britain, 1800–1910: Introduction. History and Technology, 26(2):147–149, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See response [572].

Gunn:2010:IVC

[568] Simon Gunn. Illuminating the Victorian city. History and Technology, 26 (2):151–156, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Morus:2010:IV

[569] Iwan Rhys Morus. Illuminating the Victorians. History and Technology, 26 (2):157–162, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ward:2010:WNE

[570] Janet Ward. Weimar and Nazi eyes. History and Technology, 26(2):163– 171, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Haard:2010:VEB

[571] Mikael Hård. The Victorian Eye and its blind spot: toward a cultural assessment of technology. History and Technology, 26(2):173–177, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Otter:2010:ARB

[572] Chris Otter. Author response: between liberty and discipline. History and Technology, 26(2):179–185, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See [567].

Collins:2010:ENb

[573] Martin Collins. Editor’s note. History and Technology, 26(3):187, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Moon:2010:PVI

[574] Suzanne Moon. Place, voice, interdisciplinarity: understanding technology in the colony and postcolony. History and Technology, 26(3):189–201, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Freed:2010:NCP

[575] Libbie Freed. Networks of (colonial) power: roads in French Central Africa after World War I. History and Technology, 26(3):203–223, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kusiak:2010:TTA

[576] Pauline Kusiak. “Tubab” technologies and ‘African’ ways of knowing: nationalist techno-politics in Senegal. History and Technology, 26(3):225– 249, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Smith:2010:LPC

[577] Laurel C. Smith. Locating post-colonial technoscience: through the lens of indigenous video. History and Technology, 26(3):251–280, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2010:ENc

[578] Martin Collins. Editor’s note. History and Technology, 26(4):281, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

McCray:2010:GHI

[579] W. Patrick McCray. “Globalization with hardware”: ITER’s fusion of technology, policy, and politics. History and Technology, 26(4):283–312, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Spinardi:2010:RFS

[580] Graham Spinardi. The rise and fall of Safeguard: anti-ballistic missile technology and the Nixon Administration. History and Technology, 26 (4):313–334, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Maarald:2010:MFF

[581] Erland Mårald. Methanol as future fuel: efforts to develop alternative fuels in Sweden after the Oil Crisis. History and Technology, 26(4):335–357, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2010:CSR

[582] Martin Collins. Cotten Seiler’s Republic of Drivers: A Cultural History of Automobility in America: Introduction. History and Technology, 26 (4):359–360, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Packer:2010:AAC

[583] Jeremy Packer. Automobility and apparatuses: commentary on Cotten Seiler’s Republic of Drivers. History and Technology, 26(4):361–368, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gudis:2010:DC

[584] Catherine Gudis. Driving consumption. History and Technology, 26(4): 369–378, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Flonneau:2010:RTD

[585] Mathieu Flonneau. Read Tocqueville, or drive? A European perspective on US ’automobilization’. History and Technology, 26(4):379–388, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Seiler:2010:ARE

[586] Cotten Seiler. Author response: the ends of automobility. History and Technology, 26(4):389–397, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2010:EB

[587] Anonymous. Editorial board. History and Technology, 26(4):ebi, 2010. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zachmann:2011:FTT

[588] Karin Zachmann and Per Østby. Food, technology, and trust: an introduction. History and Technology, 27(1):1–10, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Spiekermann:2011:RFS

[589] Uwe Spiekermann. Redefining food: the standardization of products and production in Europe and the United States, 1880–1914. History and Technology, 27(1):11–36, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Petrick:2011:PLH

[590] Gabriella M. Petrick. “Purity as life”: H. J. Heinz, religious sentiment, and the beginning of the industrial diet. History and Technology, 27(1):37–64, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zachmann:2011:APR

[591] Karin Zachmann. Atoms for peace and radiation for safety — how to build trust in irradiated foods in Cold War Europe and beyond. History and Technology, 27(1):65–90, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kvaal:2011:SDN

[592] Stig Kvaal and Per Østby. Sweet danger — negotiating trust in the Norwegian chocolate industry 1930–1990. History and Technology, 27(1):91–111, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

DeLue:2011:ADP

[593] Rachael Z. DeLue. Arthur Dove, painting, and phonography. History and Technology, 27(1):113–121, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Johnston:2011:SSI

[594] Sean F. Johnston. Security and the shaping of identity for nuclear specialists. History and Technology, 27(2):123–153, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Tiemeyer:2011:TGI

[595] Phil Tiemeyer. Technology and gay identity: the case of the pre-Second World War male flight attendant. History and Technology, 27(2):155–181, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2011:CMI

[596] Martin Collins. Chandra Mukerji’s Impossible Engineering: Technology and Territoriality on the Canal du Midi: Introduction. History and Technology, 27(2):183–186, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Roberts:2011:VHT

[597] Lissa Roberts. A very human tale. History and Technology, 27(2):187–195, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wise:2011:CIC

[598] M. Norton Wise. Collective intelligence and its corollaries. History and Technology, 27(2):197–203, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Appuhn:2011:WME

[599] Karl Appuhn. What makes an early modern French technology? History and Technology, 27(2):205–212, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mukerji:2011:AR

[600] Chandra Mukerji. Author response. History and Technology, 27(2):213– 222, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Weems:2011:IAS

[601] Jason Weems. Interpreting a 1930s aerial survey photograph: the artfulness of technological images. History and Technology, 27(2):223–231, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jacobson:2011:BMF

[602] Brian R. Jacobson. The Black Maria: film studio, film technology (cinema and the history of technology). History and Technology, 27(2):233–241, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2011:ENa

[603] Martin Collins. Editor’s note. History and Technology, 27(3):243, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schipper:2011:IEP

[604] Frank Schipper and Johan Schot. Infrastructural Europeanism, or the project of building Europe on infrastructures: an introduction. History and Technology, 27(3):245–264, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schot:2011:DTR

[605] Johan Schot, Hans Buiter, and Irene Anastasiadou. The dynamics of transnational railway governance in Europe during the long nineteenth century. History and Technology, 27(3):265–289, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lagendijk:2011:EFT

[606] Vincent Lagendijk. “An experience forgotten today”: examining two rounds of European electricity liberalization. History and Technology, 27 (3):291–310, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Laborie:2011:FLF

[607] Léonard Laborie. Fragile links, frozen identities: the governance of telecommunication networks and Europe (1944–53). History and Technology, 27(3):311–330, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Henrich-Franke:2011:CEC

[608] Christian Henrich-Franke and Isabel Tölle. Competition for European competence: the Central Commission for Navigation on the Rhine and the European Economic Community in the 1960s. History and Technology, 27 (3):331–352, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Griset:2011:HWW

[609] Pascal Griset and Valérie Schafer. Hosting the World Wide Web Consortium for Europe: from CERN to INRIA. History and Technology, 27 (3):353–370, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Soppelsa:2011:VVP

[610] Peter Soppelsa. Visualizing viaducts in 1880s Paris. History and Technology, 27(3):371–377, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hirsh:2011:WMP

[611] Max Hirsh. What’s missing from this picture? Using visual materials in infrastructure studies. History and Technology, 27(3):379–387, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2011:ENb

[612] Martin Collins. Editor’s note. History and Technology, 27(4):389, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fridlund:2011:BBB

[613] Mats Fridlund. Buckets, bollards and bombs: towards subject histories of technologies and terrors. History and Technology, 27(4):391–416, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Roberts:2011:GSM

[614] Lissa Roberts. Geographies of steam: mapping the entrepreneurial activities of steam engineers in France during the second half of the eighteenth century. History and Technology, 27(4):417–439, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kalba:2011:HMW

[615] Laura Anne Kalba. How media were made: chromolithography in Belle Époque France. History and Technology, 27(4):441–453, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Nielsen:2011:DDD

[616] Kristian H. Nielsen. Detachment, death, and destruction in Gerhard Richter’s Strontium (2004). History and Technology, 27(4):455–459, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2011:EB

[617] Anonymous. Editorial board. History and Technology, 27(4):ebi, 2011. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2012:ENa

[618] Martin Collins. Editor’s note. History and Technology, 28(1):1, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Plaiss:2012:WGD

[619] Adam Plaiss. Who gets to draw the map? The contentious creation of the American road/map system, 1917–1926. History and Technology, 28 (1):3–24, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Krebs:2012:SCS

[620] Stefan Krebs. Standardizing car sound — integrating Europe? International traffic noise abatement and the emergence of a European car identity, 1950–1975. History and Technology, 28(1):25–47, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Neufeld:2012:NAE

[621] Michael J. Neufeld. The Nazi aerospace exodus: towards a global, transnational history. History and Technology, 28(1):49–67, 2012. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2012:Ia

[622] Martin Collins. Introduction. History and Technology, 28(1):69–71, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Graham:2012:IPP

[623] Margaret B. W. Graham. Intellectual property and the professionals. History and Technology, 28(1):73–82, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:2012:PGC

[624] Jennifer Karns Alexander. The power to give credit and blame. History and Technology, 28(1):83–92, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Scranton:2012:TAW

[625] Philip Scranton. Thinking about work, knowledge, law and property. History and Technology, 28(1):93–100, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fisk:2012:ARL

[626] Catherine L. Fisk. Author response: Law and the problem of embodied knowledge. History and Technology, 28(1):101–106, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Coombs:2012:UET

[627] David Sweeney Coombs. An untrained eye: the tachistoscope and photographic vision in early experimental psychology. History and Technology, 28(1):107–117, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Otremba:2012:IIE

[628] Eric Otremba. Inventing ingenios: experimental philosophy and the secret sugar-makers of the seventeenth-century Atlantic. History and Technology, 28(2):119–147, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Vinsel:2012:CCE

[629] Lee Jared Vinsel. The crusade for credible energy information and analysis in the United States, 1973–1982. History and Technology, 28(2):149–176, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Apelian:2012:MML

[630] Colette Apelian. Modern mosque lamps: electricity in the historic monuments and tourist attractions of French colonial Fez, Morocco (1925– 1950). History and Technology, 28(2):177–207, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Conekin:2012:FMT

[631] Becky E. Conekin. Fashioning Mod Twiggy and the moped in ‘swinging’ London. History and Technology, 28(2):209–215, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2012:EN

[632] Anonymous. Editor’s note. History and Technology, 28(3):217, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Geppert:2012:RSA

[633] Alexander C. T. Geppert. Rethinking the space age: astroculture and technoscience. History and Technology, 28(3):219–223, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Brandau:2012:CCS

[634] Daniel Brandau. Cultivating the cosmos: spaceflight thought in Imperial Germany. History and Technology, 28(3):225–254, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Poole:2012:CSA

[635] Robert Poole. The challenge of the spaceship: Arthur C. Clarke and the history of the future, 1930–1970. History and Technology, 28(3):255–280, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Macauley:2012:CFE

[636] William R. Macauley. Crafting the future: envisioning space exploration in post-war Britain. History and Technology, 28(3):281–309, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Farry:2012:UWT

[637] James Farry and David A. Kirby. The Universe will be televised: space, science, satellites and British television production, 1946–1969. History and Technology, 28(3):311–333, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Geppert:2012:EEU

[638] Alexander C. T. Geppert. Extraterrestrial encounters: UFOs, science and the quest for transcendence, 1947–1972. History and Technology, 28(3): 335–362, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Edgington:2012:ASF

[639] Ryan Edgington. An “all-seeing flying eye”: V-2 rockets and the promises of Earth photography. History and Technology, 28(3):363–371, 2012. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2012:ENb

[640] Martin Collins. Editor’s note. History and Technology, 28(4):373, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Brock:2012:ASV

[641] David C. Brock. From automation to Silicon Valley: the automation movement of the 1950s, Arnold Beckman, and William Shockley. History and Technology, 28(4):375–401, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2012:Ib

[642] Martin Collins. Introduction. History and Technology, 28(4):403–405, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kline:2012:BCW

[643] Ronald R. Kline. Beyond the closed world. History and Technology, 28 (4):407–413, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Dennis:2012:STK

[644] Michael A. Dennis. Scientific and technical knowledge and the making of political order. History and Technology, 28(4):415–421, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Saraiva:2012:HCM

[645] Tiago Saraiva. The history of cybernetics in McOndo. History and Technology, 28(4):423–430, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Medina:2012:AR

[646] Eden Medina. Author response. History and Technology, 28(4):431–441, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kostelnick:2012:VTP

[647] Charles Kostelnick. Visualizing technology and practical knowledge in the Encyclopédie’s plates: rhetoric, drawing conventions, and enlightenment values. History and Technology, 28(4):443–454, 2012. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2013:E

[648] Anonymous. Erratum. History and Technology, 29(1):iii–??, 2013. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2013:ENa

[649] Martin Collins. Editor’s note. History and Technology, 29(1):1–2, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wellock:2013:CCE

[650] Thomas R. Wellock. The children of Chernobyl: Engineers and the campaign for safety in Soviet-designed reactors in Central and Eastern Europe. History and Technology, 29(1):3–32, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hultman:2013:ETE

[651] Martin Hultman and Christer Nordlund. Energizing technology: expectations of fuel cells and the hydrogen economy, 1990–2005. History and Technology, 29(1):33–53, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Behrent:2013:FT

[652] Michael C. Behrent. Foucault and technology. History and Technology, 29 (1):54–104, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gaudilliere:2013:PFC

[653] Jean-Paul Gaudillière and Ulrike Thoms. Pharmaceutical firms and the construction of drug markets: from branding to scientific marketing. History and Technology, 29(2):105–115, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Huntelmann:2013:DMM

[654] Axel C. Hüntelmann. A different mode of marketing? The importance of scientific articles in the marketing process of Salvarsan. History and Technology, 29(2):116–134, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bonah:2013:SQE

[655] Christian Bonah. “The Strophanthin question”: early scientific marketing of cardiac drugs in two national markets (France and Germany, 1900– 1930). History and Technology, 29(2):135–152, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kessel:2013:DCS

[656] Nils Kessel. “Doriden von Ciba”: sleeping pills, pharmaceutical marketing, and Thalidomide, 1955–1963. History and Technology, 29(2):153–168, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Thoms:2013:SSP

[657] Ulrike Thoms. Standardizing selling. Pharmaceutical marketing, the pharmaceutical company and the marketing expert (1900–1980). History and Technology, 29(2):169–187, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gaudilliere:2013:PSM

[658] Jean-Paul Gaudillière. From Propaganda to scientific marketing: Schering, cortisone, and the construction of drug markets. History and Technology, 29(2):188–209, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Greene:2013:MBF

[659] Jeremy A. Greene. The materiality of the brand: Form, function, and the pharmaceutical trademark. History and Technology, 29(2):210–226, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2013:EN

[660] Anonymous. Editor’s note. History and Technology, 29(3):227–228, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pike:2013:HF

[661] David Pike. Headlong into futurity. History and Technology, 29(3):229– 246, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Grossman:2013:CGT

[662] Jonathan H. Grossman. The character of a global transport infrastructure: Jules Verne’s Around the World in Eighty Days. History and Technology, 29(3):247–261, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Williams:2013:LBH

[663] Rosalind Williams. The lantern-bearers of the history of technology. History and Technology, 29(3):262–277, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Trotter:2013:TW

[664] David Trotter. Telephone wires. History and Technology, 29(3):278–283, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Groening:2013:AS

[665] Stephen Groening. Aerial screens. History and Technology, 29(3):284–303, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ravenscroft:2013:TCL

[666] Michael Ravenscroft. Through a cracked lens: CCTV and the urban crisis in HBO’s The Wire. History and Technology, 29(3):304–307, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2013:ENb

[667] Martin Collins. Editor’s note. History and Technology, 29(4):309–310, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Derksen:2013:HST

[668] Maarten Derksen and Tjardie Wierenga. The history of “social technology”, 1898–1930. History and Technology, 29(4):311–330, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Leslie:2013:ASC

[669] Stuart W. Leslie. Aerospaces: Southern California architecture in a Cold War world. History and Technology, 29(4):331–368, 2013. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schuring:2013:ANV

[670] Michael Schüring. Advertising the nuclear venture: the rhetorical and visual public relation strategies of the German nuclear industry in the 1970s and 1980s. History and Technology, 29(4):369–398, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Golec:2013:PPR

[671] Michael J. Golec. Poster power: rural electrification, visualization, and legibility in the United States. History and Technology, 29(4):399–410, 2013. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2014:ENa

[672] Martin Collins. Editor’s note. History and Technology, 30(1–2):1–2, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bud:2014:ASN

[673] Robert Bud. “Applied science” in nineteenth-century Britain: public discourse and the creation of meaning, 1817–1876. History and Technology, 30(1–2):3–36, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gouyon:2014:MSH

[674] Jean-Baptiste Gouyon. Making science at home: visual displays of space science and nuclear physics at the Science Museum and on television in postwar Britain. History and Technology, 30(1–2):37–60, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Giffard:2014:PDT

[675] Hermione Giffard. The politics of donating technological artifacts: technonationalism and the donations of the world’s first jet engines. History and Technology, 30(1–2):61–82, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cozzoli:2014:PER

[676] Daniele Cozzoli. Penicillin and the European response to post-war American hegemony: the case of Leo-penicillin. History and Technology, 30 (1–2):83–103, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Andersson:2014:GFS

[677] Jenny Andersson and Anne-Greet Keizer. Governing the future: science, policy and public participation in the construction of the long term in The Netherlands and Sweden. History and Technology, 30(1–2):104–122, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kryczka:2014:TTN

[678] Anna Kryczka. Television and taste on the New Frontier: “A Tour of the White House With Mrs. John F. Kennedy”. History and Technology, 30 (1–2):123–132, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Roberts:2014:POD

[679] Lissa Roberts. Practicing oeconomy during the second half of the long eighteenth century: an introduction. History and Technology, 30(3):133– 148, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Serrano:2014:MOP

[680] Elena Serrano. Making oeconomic people: The Spanish Magazine of Agriculture and Arts for Parish Rectors (1797–1808). History and Technology, 30(3):149–176, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vanDriel:2014:AAS

[681] Joppe van Driel. Ashes to ashes: the stewardship of waste and oeconomic cycles of agricultural and industrial improvement, 1750–1800. History and Technology, 30(3):177–206, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Stockland:2014:PON

[682] Etienne Stockland. Policing the oeconomy of nature: The oiseau martin as an instrument of oeconomic management in the eighteenth-century French maritime world. History and Technology, 30(3):207–231, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wakefield:2014:BNH

[683] Andre Wakefield. Butterfield’s nightmare: the history of science as Disney history. History and Technology, 30(3):232–251, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See note [700].

Collins:2014:AVA

[684] Martin Collins. Adelheid Voskuhl’s Androids in the Enlightenment: Mechanics, Artisans, and Cultures of the Self. History and Technology, 30 (3):252–254, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mukerji:2014:EQC

[685] Chandra Mukerji. Entangled in questions of cultural analysis. History and Technology, 30(3):255–260, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Crawford:2014:DWG

[686] Katherine B. Crawford. Designed women: gender and the problem of female automata. History and Technology, 30(3):261–268, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gooday:2014:MOH

[687] Graeme Gooday. Museums, objects, and historical meaning. History and Technology, 30(3):269–274, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Voskuhl:2014:SKH

[688] Adelheid Voskuhl. Self-knowledge in the history of technology: industrialism, cultural analysis and desire. History and Technology, 30(3):275–279, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2014:ENb

[689] Martin Collins. Editor’s note. History and Technology, 30(4):281–282, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Martland:2014:SSW

[690] Samuel J. Martland. Standardizing the state while integrating the frontier: the Chilean telegraph system in the Araucanía, 1870–1900. History and Technology, 30(4):283–308, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sumner:2014:DCV

[691] James Sumner. Defiance to compliance: Visions of the computer in postwar Britain. History and Technology, 30(4):309–333, 2014. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Choi:2014:EON

[692] Hyungsub Choi. Emerging opportunities: nanoelectronics and engineering research in a South Korean university. History and Technology, 30(4): 334–353, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Scranton:2014:PFH

[693] Philip Scranton. Projects as a focus for historical analysis: surveying the landscape. History and Technology, 30(4):354–373, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Shell:2014:SHM

[694] Hanna Rose Shell. Shoddy heap: a material history between waste and manufacture. History and Technology, 30(4):374–394, 2014. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2014:EB

[695] Anonymous. Editorial board. History and Technology, 30(4):ebi, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2015:ENa

[696] Martin Collins. Editor’s note. History and Technology, 31(1):1, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

vanderHoogte:2015:SIC

[697] Arjo Roersch van der Hoogte and Toine Pieters. Science, industry and the colonial state: a shift from a German- to a Dutch-controlled cinchona and quinine cartel (1880–1920). History and Technology, 31(1):2–36, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lavine:2015:SAX

[698] Matthew Lavine. “Something about X-rays for everyone”: emerging technologies and open communities. History and Technology, 31(1):37–54, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kasperski:2015:NDR

[699] Tatiana Kasperski. Nuclear dreams and realities in contemporary Russia and Ukraine. History and Technology, 31(1):55–80, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Popplow:2015:NDN

[700] Marcus Popplow. Nightmares in Disneyland — a note on Andre Wakefield’s critique of non-actors’ terms and alleged Cold War rhetoric. History and Technology, 31(1):81–82, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See [683].

Collins:2015:ENb

[701] Martin Collins. Editor’s note. History and Technology, 31(2):83, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hanel:2015:ITI

[702] Tilmann Hanel and Mikael Hård. Inventing traditions: interests, parables and nostalgia in the history of nuclear energy. History and Technology, 31 (2):84–107, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kochetkova:2015:HFI

[703] Elena Kochetkova. A history of failed innovation: continuous cooking and the Soviet pulp industry, 1940s-1960s. History and Technology, 31 (2):108–132, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hundebol:2015:PCA

[704] Nils Randlev Hundebøl and Kristian H. Nielsen. Preparing for change: acid rain, climate change, and the Electric Power Research Institute (EPRI), 1972–1990s. History and Technology, 31(2):133–159, 2015. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Buss:2015:VWS

[705] Jared S. Buss. Virtual witnessing and space-age media: a case study of The Conquest of Space (1949). History and Technology, 31(2):160–170, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Krige:2015:NKI

[706] John Krige and Jessica Wang. Nation, knowledge, and imagined futures: Science, technology, and nation-building, post-1945. History and Technology, 31(3):171–179, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wang:2015:CDS

[707] Zuoyue Wang. The Chinese developmental state during the Cold War: the making of the 1956 Twelve-Year Science and Technology Plan. History and Technology, 31(3):180–205, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Smith:2015:AYD

[708] Jenny Leigh Smith. The awkward years: defining and managing famines, 1944–1947. History and Technology, 31(3):206–219, 2015. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Leslie:2015:ASA

[709] Stuart W. Leslie. Atomic structures: the architecture of nuclear nationalism in India and Pakistan. History and Technology, 31(3):220–242, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mateos:2015:WRC

[710] Gisela Mateos and Edna Suárez-Díaz. “We are not a rich country to waste our resources on expensive toys”: Mexico’s version of Atoms for Peace. History and Technology, 31(3):243–258, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hirshberg:2015:NSU

[711] Lauren Hirshberg. “Navigating sovereignty under a Cold War military industrial colonial complex: US military empire and Marshallese decolonization”. History and Technology, 31(3):259–274, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Laveaga:2015:BNF

[712] Gabriela Soto Laveaga. Building the nation of the future, one waiting room at a time: hospital murals in the making of modern Mexico. History and Technology, 31(3):275–294, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mavhunga:2015:GHI

[713] Clapperton Chakanetsa Mavhunga. Guerrilla healthcare innovation: creative resilience in Zimbabwe’s chimurenga, 1971–1980. History and Technology, 31(3):295–323, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mukharji:2015:TIM

[714] Projit Bihari Mukharji. Technospatial imaginaries: Masud Rana and the vernacularization of popular Cold War geopolitics in East Pakistan, 1966– 1971. History and Technology, 31(3):324–340, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2015:E

[715] Anonymous. Erratum. History and Technology, 31(3):i, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Siddiqi:2015:TSA

[716] Asif A. Siddiqi. Technology in the South Asian imaginary. History and Technology, 31(4):341–349, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mitra:2015:TTF

[717] Durba Mitra. Translation as techné: female sexuality and the science of social progress in colonial India. History and Technology, 31(4):350–375, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mukharji:2015:PPF

[718] Projit Bihari Mukharji. Profiling the profiloscope: facialization of race technologies and the rise of biometric nationalism in inter-war British India. History and Technology, 31(4):376–396, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sackley:2015:RSE

[719] Nicole Sackley. The road from serfdom: economic storytelling and narratives of India in the rise of neoliberalism. History and Technology, 31 (4):397–419, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Siddiqi:2015:SGN

[720] Asif A. Siddiqi. Science, geography, and nation: the global creation of Thumba. History and Technology, 31(4):420–451, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Phalkey:2015:ATT

[721] Jahnavi Phalkey and Sumandro Chattapadhyay. The Aakash tablet and technological imaginaries of mass education in contemporary India. History and Technology, 31(4):452–481, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pauwels:2015:RCC

[722] Erin Pauwels. Resetting the camera’s clock: Sarony, Muybridge & the aesthetics of wet-plate photography. History and Technology, 31(4):482–491, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2015:EB

[723] Anonymous. Editorial board. History and Technology, 31(4):ebi, 2015. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Collins:2016:EN

[724] Martin Collins. Editor’s note. History and Technology, 32(1):1, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guerrero:2016:HSR

[725] Saúl Guerrero. The history of silver refining in New Spain, 16c to 18c: back to the basics. History and Technology, 32(1):2–32, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cantoni:2016:EUG

[726] Roberto Cantoni. The enemy underground: geostrategic intelligence and the war in Algeria. History and Technology, 32(1):33–69, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lieberman:2016:MFA

[727] Jennifer L. Lieberman. The myth of the first African–American electrical engineer: Arthur U. Craig and the importance of teaching in technological history. History and Technology, 32(1):70–90, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bate:2016:FEP

[728] Jason Bate. Facilitating exchanges: photography as a link between dentistry and surgery in the First World War. History and Technology, 32 (1):91–103, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Muller:2016:IUC

[729] Simone M. Müller and Heidi J. S. Tworek. Imagined use as a category of analysis: new approaches to the history of technology. History and Technology, 32(2):105–??, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ellis:2016:MMH

[730] Heather Ellis. Marconi, masculinity and the heroic age of science: wireless telegraphy at the British Association meeting at Dover in 1899. History and Technology, 32(2):120–??, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Noakes:2016:TSW

[731] Richard Noakes. Thoughts and spirits by wireless: imagining and building psychic telegraphs in America and Britain, circa 1900–1930. History and Technology, 32(2):137–??, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Davies:2016:SFC

[732] Hannah Catherine Davies. Spreading fear, communicating trust: writing letters and telegrams during the Panic of 1873. History and Technology, 32(2):159–??, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Tworek:2016:HBW

[733] Heidi J. S. Tworek. How not to build a world wireless network: German– British rivalry and visions of global communications in the early twentieth century. History and Technology, 32(2):178–??, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mosley:2016:SSC

[734] Stephen Mosley. Selling the smokeless city: advertising images and smoke abatement in urban-industrial Britain, circa 1840–1960. History and Technology, 32(2):201–??, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bruton:2016:LCS

[735] Elizabeth Bruton and Graeme Gooday. Listening in combat — surveillance technologies beyond the visual in the First World War. History and Technology, 32(3):213–226, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Judkins:2016:SFS

[736] Phil Judkins. Sound and fury: sound and vision in early U.K. air defence. History and Technology, 32(3):227–244, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bruton:2016:LDA

[737] Elizabeth Bruton and Paul Coleman. Listening in the dark: audio surveillance, communication technologies, and the submarine threat during the First World War. History and Technology, 32(3):245–268, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schirrmacher:2016:SRW

[738] Arne Schirrmacher. Sounds and repercussions of war: mobilization, invention and conversion of First World War science in Britain, France and Germany. History and Technology, 32(3):269–292, 2016. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Marklund:2016:LSC

[739] Andreas Marklund. Listening for the state: censoring communications in Scandinavia during World War I. History and Technology, 32(3):293–314, 2016. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bjorck:2017:DSF

[740] Henrik Björck. A distinguished scientific field? Pursuing resources and building institutions for engineering research in Sweden, 1890–1945. History and Technology, 32(4):315–348, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pettersson:2017:TSC

[741] Ingemar Pettersson. The technoscientists: a Swedish construction of basic research in World War II. History and Technology, 32(4):349–381, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Crowley:2017:TCP

[742] Mark J. Crowley. Technological change and Post Office communications in Britain, 1918–1945. History and Technology, 32(4):382–408, 2017. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Morawski:2017:TSP

[743] Jill G. Morawski and Matthew Donahue. Technological simulacrum: psychologists’ invention of a bogus pipeline to the soul. History and Technology, 32(4):409–418, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2017:EB

[744] Anonymous. Editorial board. History and Technology, 32(4):ebi, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2017:E

[745] Amy Slaton and Tiago Saraiva. Editorial. History and Technology, 33(1): 1–3, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mercelis:2017:CSN

[746] Joris Mercelis, Gabriel Galvez-Behar, and Anna Guagnini. Commercializing science: nineteenth- and twentieth-century academic scientists as consultants, patentees, and entrepreneurs. History and Technology, 33(1): 4–22, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mercelis:2017:CAK

[747] Joris Mercelis. Commercializing academic knowledge and reputation in the late nineteenth and early twentieth centuries: photography and beyond. History and Technology, 33(1):23–52, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Konig:2017:EPE

[748] Wolfgang König. Engineering professors as entrepreneurs: the case of Franz Reuleaux (1829–1905) and Alois Riedler (1850–1936). History and Technology, 33(1):53–69, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Guagnini:2017:ITC

[749] Anna Guagnini. Ivory towers? The commercial activity of British professors of engineering and physics, 1880–1914. History and Technology, 33 (1):70–108, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Katzir:2017:TEP

[750] Shaul Katzir. Technological entrepreneurship from patenting to commercializing: a survey of late nineteenth and early twentieth century physics lecturers. History and Technology, 33(1):109–125, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Dick:2017:CMB

[751] Brian Dick and Mark Jones. The commercialization of molecular biology: Walter Gilbert and the Biogen startup. History and Technology, 33(1): 126–151, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Zhang:2017:MFU

[752] Jun Zhang. Materializing a form of urban governance: when street building intersected with city building in Republican Canton (Guangzhou), China. History and Technology, 33(2):153–174, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pereira:2017:TIT

[753] Hugo Silveira Pereira. The technodiplomacy of Iberian transnational railways in the second half of the nineteenth century. History and Technology, 33(2):175–195, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Johnston:2017:TPI

[754] Sean F. Johnston. Technological parables and iconic illustrations: American technocracy and the rhetoric of the technological fix. History and Technology, 33(2):196–219, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Matala:2017:SNB

[755] Saara Matala and Aaro Sahari. Small nation, big ships: winter navigation and technological nationalism in a peripheral country, 1878–1978. History and Technology, 33(2):220–248, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Thompson:2017:CSP

[756] Peter Thompson. The chemical subject: phenomenology and German encounters with the gas mask in the World War I. History and Technology, 33(3):249–271, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Munns:2017:LAS

[757] David P. D. Munns and Kärin Nickelsen. To live among the stars: artificial environments in the early space age. History and Technology, 33 (3):272–299, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Heinze:2017:RSN

[758] Thomas Heinze and Olof Hallonsten. The reinvention of the SLAC National Accelerator Laboratory, 1992–2012. History and Technology, 33 (3):300–332, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Grevsmuhl:2017:VHO

[759] Sebastian Vincent Grevsmühl. A visual history of the ozone hole: a journey to the heart of science, technology and the global environment. History and Technology, 33(3):333–344, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Choi:2017:IMG

[760] Hyungsub Choi. Imported machines in the garden: the kyongun’gi (power tiller) and agricultural mechanization in South Korea. History and Technology, 33(4):345–366, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Carone:2017:FVH

[761] Justin Carone. Fixing value: history, ethnography, and material ontologies of deservingness in a Philadelphia repair shop. History and Technology, 33 (4):367–395, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cramer:2017:LEE

[762] Katharina Christine Cramer. Lightening Europe: Establishing the European Synchrotron Radiation Facility (ESRF). History and Technology, 33 (4):396–427, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fletcher:2017:STP

[763] Curtis Fletcher. The school of tomorrow: promoting electronic multimedia education in the 1960s. History and Technology, 33(4):428–440, 2017. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hayton:2018:C

[764] Darin Hayton. On the cover. History and Technology, 34(1):1–3, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2018:EN

[765] Amy Slaton and Tiago Saraiva. Editors’ note. History and Technology, 34 (1):4, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

SotoLaveaga:2018:I

[766] Gabriela Soto Laveaga and Pablo F. Gómez. Introduction. History and Technology, 34(1):5–10, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Gomez:2018:CSC

[767] Pablo F. Gómez. Caribbean stones and the creation of early-modern worlds. History and Technology, 34(1):11–20, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

SotoLaveaga:2018:LDC

[768] Gabriela Soto Laveaga. Largo dislocare: connecting microhistories to remap and recenter histories of science. History and Technology, 34(1): 21–30, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mukharji:2018:OM

[769] Projit Bihari Mukharji. Occulted materialities. History and Technology, 34(1):31–40, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

deLuna:2018:ILK

[770] Kathryn de Luna. Inciteful language: knowing and naming technology in south central Africa. History and Technology, 34(1):41–50, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Prasad:2018:TMC

[771] Amit Prasad. Taj Mahal, circulations of science, and (post) colonial present. History and Technology, 34(1):51–60, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Canizares-Esguerra:2018:BIM

[772] Jorge Cañizares-Esguerra. Bartolomé Inga’s mining technologies: Indians, science, cyphered secrecy, and modernity in the new world. History and Technology, 34(1):61–70, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Podgorny:2018:HCU

[773] Irina Podgorny. A horse-cloth for Uganda, or how an account by a transhumant veterinary connects histories, animal diseases and continents. History and Technology, 34(1):71–78, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ragab:2018:TSC

[774] Ahmed Ragab. Two students and a corpse: the semantics of disgust in the making of colonial knowledge. History and Technology, 34(1):79–88, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Saraiva:2018:ASC

[775] Tiago Saraiva. Anthropophagy and sadness: cloning citrus in São Paulo in the Plantationocene era. History and Technology, 34(1):89–99, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sorlin:2018:ETT

[776] Sverker Sörlin and Nina Wormbs. Environing technologies: a theory of making environment. History and Technology, 34(2):101–125, 2018. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Epstein:2018:IPN

[777] Katherine C. Epstein. Intellectual property and national security: the case of the Hardcastle superheater, 1905–1927. History and Technology, 34(2): 126–156, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Frumer:2018:CEJ

[778] Yulia Frumer. Cognition and emotions in Japanese humanoid robotics. History and Technology, 34(2):157–183, 2018. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2019:ENa

[779] Amy Slaton and Tiago Saraiva. Editors’ note. History and Technology, 34(3–4):185–186, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Trischler:2019:PTN

[780] Helmuth Trischler and Robert Bud. Public technology: nuclear energy in Europe. History and Technology, 34(3–4):187–212, 2019. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ruuskanen:2019:EBM

[781] Esa Ruuskanen. The emergence of Baltic Moorkultur: visions of scientifictechnological mastery of peatlands in the age of great social change, 1850– 1914. History and Technology, 34(3–4):213–234, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Valentines-Alvarez:2019:SLF

[782] Jaume Valentines-Álvarez. Seeing like a factory: Technocratic nationalism in Catalonia, 1930–1939. History and Technology, 34(3–4):235–258, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Perkins:2019:IAT

[783] Emma Perkins. Instruments of authority: Tycho Brahe’s technological illustrations. History and Technology, 34(3–4):259–272, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Saraiva:2019:C

[784] Tiago Saraiva. On the cover. History and Technology, 35(1):1–2, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ash:2019:ONE

[785] Eric H. Ash. By any other name: early modern expertise and the problem of anachronism. History and Technology, 35(1):3–30, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Skyggebjerg:2019:KMC

[786] Louise Karlskov Skyggebjerg. Knowledge making and corporate environmentalism from the perspective of the egg tray. History and Technology, 35(1):31–57, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Petersen:2019:NLT

[787] Sonja Petersen. The notebook and the laboratory — types of knowledge in German piano-making, 1880–1930. History and Technology, 35(1):58–80, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Paris:2019:BEC

[788] Ivan Paris. Between efficiency and comfort: the organization of domestic work and space from home economics to scientific management, 1841– 1913. History and Technology, 35(1):81–104, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Balbi:2019:TTH

[789] Gabriele Balbi and Christiane Berth. Towards a telephonic history of technology. History and Technology, 35(2):105–114, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

John:2019:CVH

[790] Richard R. John and Léonard Laborie. ‘Circuits of victory’: how the First World War shaped the political economy of the telephone in the United States and France. History and Technology, 35(2):115–137, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

McGuire:2019:CHL

[791] Coreen Anne McGuire. The categorisation of hearing loss through telephony in inter-war Britain. History and Technology, 35(2):138–155, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Weber:2019:ERT

[792] Heike Weber. Everyman’s radio-telephone: how CB radio users mobilized West Germany’s communication culture. History and Technology, 35(2):

156–176, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wasiak:2019:TNB

[793] Patryk Wasiak. Telephone networks, BBSes, and the emergence of the transnational ‘warez scene’. History and Technology, 35(2):177–194, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jansen:2019:EKI

[794] Axel Jansen, John Krige, and Jessica Wang. Empires of knowledge: introduction. History and Technology, 35(3):195–202, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wang:2019:PIB

[795] Jessica Wang. Plants, insects, and the biological management of American empire: tropical agriculture in early twentieth-century Hawai‘i. History and Technology, 35(3):203–236, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hendrickson:2019:AAE

[796] Mark Hendrickson. Advance agent of expanding empires: George F. Becker and mineral exploration in South Africa and the Philippines. History and Technology, 35(3):237–265, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Webel:2019:TTM

[797] Mari K. Webel. Trypanosomiasis, tropical medicine, and the practices of inter-colonial research at Lake Victoria, 1902–07. History and Technology, 35(3):266–292, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ventura:2019:PPP

[798] Theresa Ventura. Prison, plantation, and peninsula: colonial knowledge and experimental technique in the post-war Bataan Rice Enrichment Project, 1910–1950. History and Technology, 35(3):293–315, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lagendijk:2019:SKR

[799] Vincent Lagendijk. Streams of knowledge: river development knowledge and the TVA on the river Mekong. History and Technology, 35(3):316–337, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schulze:2019:SDU

[800] Frederik Schulze. In search of El Dorado: U.S. experts and the promise of development in the Guayana region of Venezuela. History and Technology, 35(3):338–363, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2019:ENb

[801] Amy Slaton and Tiago Saraiva. Editors’ note. History and Technology, 35(4):365, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Rader:2019:ISI

[802] Karen A. Rader. Introduction to the special issue Biology and Technology Reframed: historiographical reflections and opportunities. History and Technology, 35(4):366–373, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Berry:2019:MDB

[803] Dominic J. Berry. Making DNA and its becoming an experimental commodity. History and Technology, 35(4):374–404, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lee:2019:WTC

[804] Victoria Lee. Wild toxicity, cultivated safety: aflatoxin and koji classification as knowledge infrastructure. History and Technology, 35(4):405–424, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Myelnikov:2019:TGE

[805] Dmitriy Myelnikov. Tinkering with genes and embryos: the multiple invention of transgenic mice c. 1980. History and Technology, 35(4):425–452, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ling:2019:DKS

[806] Vivian Ling and Lijing Jiang. A different kind of synthesis: artificial synthesis of insulin in socialist China. History and Technology, 35(4): 453–480, 2019. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2020:AC

[807] Amy Slaton and Tiago Saraiva. About the cover. History and Technology, 36(1):1–2, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jones-Imhotep:2020:GFH

[808] Edward Jones-Imhotep. The ghost factories: histories of automata and artificial life. History and Technology, 36(1):3–29, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Robertson:2020:RAB

[809] Beth A. Robertson. ‘Rehabilitation aids for the blind’: disability and technological knowledge in Canada, 1947–1985. History and Technology, 36(1):30–53, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lachney:2020:HMD

[810] Michael Lachney and Ellen K. Foster. Historicizing making and doing: Seymour Papert, Sherry Turkle, and epistemological foundations of the maker movement. History and Technology, 36(1):54–82, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Peyerl:2020:AGB

[811] Drielli Peyerl and Silvia Fernanda de Mendonça Figueirôa. Applied geophysics in Brazil and the development of a national oil industry (1930 — 1960). History and Technology, 36(1):83–104, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Tarr:2020:ISA

[812] Joel A. Tarr. Illuminating the streets, alleys, parks and suburbs of the American City: non-networked technologies, 1870–1920. History and Technology, 36(1):105–128, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Esselborn:2020:NSN

[813] Stefan Esselborn and Karin Zachmann. Nuclear safety by numbers. Probabilistic risk analysis as an evidence practice for technical safety in the German debate on nuclear energy. History and Technology, 36(1):129–164, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:2020:IET

[814] Jennifer Karns Alexander. Introduction: the entanglement of technology and religion. History and Technology, 36(2):165–186, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Stoneman:2020:PDR

[815] Timothy H. B. Stoneman. Presencing the divine: religion and technology in the Latin West. History and Technology, 36(2):187–204, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bix:2020:RSH

[816] Amy Sue Bix. ‘Remember the Sabbath’: a history of technological decisions and innovation in Orthodox Jewish communities. History and Technology, 36(2):205–239, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Moon:2020:SOU

[817] Suzanne Moon. A sociotechnical order for the umma: connecting Islam and technology in Suharto’s Indonesia. History and Technology, 36(2): 240–262, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Alexander:2020:ECI

[818] Jennifer Karns Alexander. An engineering career as industrial mission: Jack Keiser in post-war Britain. History and Technology, 36(2):263–291, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mateos:2020:DIS

[819] Gisela Mateos and Edna Suárez-Díaz. Development interventions: science, technology and technical assistance. History and Technology, 36(3–4): 293–309, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wang:2020:AER

[820] Jessica Wang. Agricultural expertise, race, and economic development: small producer ideology and settler colonialism in the Territory of Hawai‘i, 1900–1917. History and Technology, 36(3–4):310–336, 2020. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Laveaga:2020:SOG

[821] Gabriela Soto Laveaga. The socialist origins of the Green Revolution: Pandurang Khankhoje and domestic ‘technical assistance’. History and Technology, 36(3–4):337–359, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mizuno:2020:MRA

[822] Hiromi Mizuno. Mutant rice and agricultural modernization in Asia. History and Technology, 36(3–4):360–381, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

DiMoia:2020:RTI

[823] John P. DiMoia. Reconfiguring transport infrastructure in post-war Asia: mapping South Korean container ports, 1952–1978. History and Technology, 36(3–4):382–399, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Vargha:2020:TAS

[824] Dora Vargha. Technical assistance and socialist international health: Hungary, the WHO and the Korean War. History and Technology, 36(3–4): 400–417, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mateos:2020:CNM

[825] Gisela Mateos and Edna Suárez-Díaz. Creating the need in Mexico: the IAEA’s technical assistance programs for less developed countries (1958– 68). History and Technology, 36(3–4):418–436, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hamblin:2020:AMN

[826] Jacob Darwin Hamblin. Aligning missions: nuclear technical assistance, the IAEA, and national ambitions in Pakistan. History and Technology, 36(3–4):437–451, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Siddiqi:2020:WIS

[827] Asif A. Siddiqi. Whose India? SITE and the origins of satellite television in India. History and Technology, 36(3–4):452–474, 2020. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2021:EN

[828] Amy Slaton and Jesse Smith. Editors’ note. History and Technology, 37(1):1–3, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ito:2021:CPN

[829] Kenji Ito and Maria Rentetzi. The co-production of nuclear science and diplomacy: towards a transnational understanding of nuclear things. History and Technology, 37(1):4–20, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ivacs:2021:PFT

[830] Gabriella Ivacs. From paper files to terabytes: the evolution of IAEA documentation in the nuclear age. History and Technology, 37(1):21–24, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kyrtsis:2021:LBD

[831] Alexandros-Andreas Kyrtsis and Maria Rentetzi. From lobbyists to backstage diplomats: how insurers in the field of third party liability shaped nuclear diplomacy. History and Technology, 37(1):25–43, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hoey:2021:CCN

[832] Fintan Hoey. The ‘conceit of controllability’: nuclear diplomacy, Japan’s plutonium reprocessing ambitions and US proliferation fears, 1974–1978. History and Technology, 37(1):44–66, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ito:2021:TTU

[833] Kenji Ito. Three tons of uranium from the International Atomic Energy Agency: diplomacy over nuclear fuel for the Japan Research Reactor-3 at the Board of Governors’ meetings, 1958–1959. History and Technology, 37(1):67–89, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mateos:2021:AAI

[834] Gisela Mateos and Edna Suárez-Díaz. Atomic ambassadors: the IAEA’s first Preliminary Assistance Mission (1958). History and Technology, 37 (1):90–105, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Aaberg:2021:WMI

[835] Anna Åberg. The ways and means of ITER: reciprocity and compromise in fusion science diplomacy. History and Technology, 37(1):106–124, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Krige:2021:MTH

[836] John Krige and Sabina Leonelli. Mobilizing the transnational history of knowledge flows. COVID-19 and the politics of research at the borders. History and Technology, 37(1):125–146, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Fretwell:2021:TTT

[837] Elizabeth Ann Fretwell. The tools of tailoring as technologies-in-use in twentieth century Benin, West Africa. History and Technology, 37(2): 147–171, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pohl:2021:UEE

[838] Dennis Pohl. Uranium exposed at Expo 58: the colonial agenda behind the peaceful atom. History and Technology, 37(2):172–202, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Clifton-Morekis:2021:FLF

[839] Alice Shackelford Clifton-Morekis. Front-line fowl: Messenger pigeons as communications technology in the U.S. Army. History and Technology, 37 (2):203–246, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Turnbull:2021:EHH

[840] Thomas Turnbull. Energy, history, and the humanities: against a new determinism. History and Technology, 37(2):247–292, 2021. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Rijke-Epstein:2021:HTC

[841] Tasha Rijke-Epstein. On humble technologies: containers, care, and water infrastructure in northwest Madagascar, 1750s–1960s. History and Technology, 37(3):293–328, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Park:2021:RCI

[842] Seohyun Park. Reassembling colonial infrastructure in Cold War Korea: the Han River Basin Joint Survey Project (1966–71). History and Technology, 37(3):329–354, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Barton:2021:WWS

[843] Jenifer Barton. ‘We were shot down!’: Earth observing satellites, data surveillance, and NASA’s 1982 Global Habitability initiative. History and Technology, 37(3):355–378, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Garrido:2021:MNT

[844] Francisco Garrido and Ricardo Paredes. Modernizing a nation through its radio and television industry: RCA Victor in Chile, 1928–1973. History and Technology, 37(3):379–395, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hasenohrl:2021:HRI

[845] Ute Hasenöhrl and Patrick Kupper. Historicizing renewables: issues and challenges. History and Technology, 37(4):397–410, 2021. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Manuel:2021:LFF

[846] Jeffrey T. Manuel. Lessons from a forgotten fuel: assessing the long history of alcohol fuel advocacy and use in the United States. History and Technology, 37(4):411–428, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Uekotter:2021:RCB

[847] Frank Uekötter. The revolt of the chemists: biofuels, agricultural overproduction, and the chemurgy movement in New Deal America. History and Technology, 37(4):429–445, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Hesse:2021:WPR

[848] Nicole Hesse. Wind power and rural modernization: wind-powered water supply systems in northern Germany and southern France, 1880–1950. History and Technology, 37(4):446–467, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kaijser:2021:DWS

[849] Arne Kaijser. Driving on wood: the Swedish transition to wood gas during World War Two. History and Technology, 37(4):468–486, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Frey:2021:POW

[850] Felix Frey. Putting oceans to work: tidal energy in the USA and the USSR, 1930–1970. History and Technology, 37(4):487–504, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Pallua:2021:MSH

[851] Irene Pallua. The materiality of space heating: heat pumps and heating transitions in Twentieth-century Switzerland. History and Technology, 37 (4):505–526, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Melsted:2021:EFF

[852] Odinn Melsted. Eliminating fossil fuels: Iceland’s transition from coal and oil to geothermal district heating, 1930–1980. History and Technology, 37 (4):527–547, 2021. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2022:AC

[853] Amy Slaton and Tiago Saraiva. About the cover. History and Technology, 38(1):1–2, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sxenchez:2022:TFS

[854] Esther M. Sánchez Sánchez. The training in France of Spanish nuclear personnel, c. 1950s–1990s. History and Technology, 38(1):3–30, 2022. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Clarke:2022:TDD

[855] Sabine Clarke and Thomas Lean. Turning DDT into ‘Didimac’: Making insecticide products and consumers in British farming after 1945. History and Technology, 38(1):31–61, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Mirza:2022:SAI

[856] Priya Mirza. ‘Sovereignty of the air’: The Indian princely states, the British Empire and carving out of air-space (1911–1933). History and Technology, 38(1):62–83, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Jeon:2022:DCS

[857] Chihyung Jeon, Scott Gabriel Knowles, and Sang-Eun Park. Disaster (continued): Sewol Ferry investigations, state violence, and political history in South Korea. History and Technology, 38(1):84–106, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Seow:2022:MHT

[858] Victor Seow and Dagmar Schäfer. Making history: technologies of production and the estate of knowledge in East Asia. History and Technology, 38(2–3):107–125, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kobiljski:2022:MRM

[859] Aleksandra Kobiljski and Sarah Teasley. Making raw materials: innovation and imported technology in Meiji Japan. History and Technology, 38 (2–3):126–143, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Kang:2022:REH

[860] Hyeok Hweon Kang. Reverse engineering as history and method: The Portuguese espingarda in Choson Korea. History and Technology, 38(2– 3):144–166, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Ghosh:2022:MMC

[861] Arunabh Ghosh. Multiple makings at China’s first hydroelectric power station at Shilongba, 1908–1912. History and Technology, 38(2–3):167– 185, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lee:2022:MWO

[862] Jung Lee. Making waste one’s own: transformations in production by resting paper, or hyuji, in Choson Korea. History and Technology, 38(2– 3):186–204, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Choi:2022:BLS

[863] Hyungsub Choi. Before localization: the story of the electric rice cooker in South Korea. History and Technology, 38(2–3):205–221, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Chen:2022:IMT

[864] Kaijun Chen. Imperial models: technology and design in state-controlled porcelain manufacture in early modern China. History and Technology, 38(2–3):222–238, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Frumer:2022:MHR

[865] Yulia Frumer. Manufacturing hands: robot fingers and human labour in post-war Japan. History and Technology, 38(2–3):239–256, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Seow:2022:PTI

[866] Victor Seow. Psychology as technology: industrial psychology for an industrializing China. History and Technology, 38(2–3):257–273, 2022. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bray:2022:GTP

[867] Francesca Bray and Barbara Hahn. ‘The Goddess Technology is a polyglot’: a critical review of Eric Schatzberg, Technology: critical history of a concept. History and Technology, 38(4):275–316, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See response [884].

Bruno:2022:UUA

[868] Karl Bruno. Use and users of artificial insemination in Swedish dairy cattle breeding, 1935–1955. History and Technology, 38(4):317–343, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Picabea:2022:AWL

[869] Facundo Picabea. Aircraft without wings: local design and serial production of utilitarian vehicles in Argentina (1952–1955). History and Technology, 38(4):344–365, 2022. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Bulstrode:2023:BMM

[870] Jenny Bulstrode. Black metallurgists and the making of the industrial revolution. History and Technology, 39(1):1–41, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See correction [888].

Cropper:2023:SLM

[871] John Cropper. “The sparrow loves millet, but labors not“: Energy use and infrastructure in the Senegal Valley, 1450–1760. History and Technology, 39(1):42–64, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See correction [887].

Paterson:2023:FPP

[872] Mark Paterson. Fatigue as a physiological problem: experiments in the observation and quantification of movement and industrial labor, 1873-1947. History and Technology, 39(1):65–90, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Perez-Zapico:2023:EFR

[873] Daniel Pérez-Zapico. Electrical futures for a regenerated Spain: electricity, engineering and national reconstruction after the 1898 ‘Disaster’. History and Technology, 39(1):91–125, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2023:E

[874] Amy E. Slaton and Tiago Saraiva. Editorial. History and Technology, 39 (2):127–136, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Slaton:2023:I

[875] Amy E. Slaton and Tiago Saraiva. Introduction. History and Technology, 39(2):137–140, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Cucu:2023:OLC

[876] Alina-Sandra Cucu and Bridget Kenny. The ordinary lives of crisis: transformations in the realm of work in South Africa and Romania. History and Technology, 39(2):141–155, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lachney:2023:IME

[877] Michael Lachney and Madison C. Allen Kuyenga. Irreplicability in methodology: embracing the historical contingencies of educational technology research during the 2020–2021 United States school year. History and Technology, 39(2):156–175, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Szczygielska:2023:TVT

[878] Marianna Szczygielska and Agata Kowalewska. Tracing viral trajectories. epistemic and bodily reservoirs in interspecies health. History and Technology, 39(2):176–192, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Older:2023:CTP

[879] Malka Older and Scott Gabriel Knowles. CORONALAG: time, place, and power in Pandemic Year One. History and Technology, 39(2):193–207, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wernimont:2023:PMD

[880] Jacqueline D. Wernimont. Processing mortality data otherwise: making history in a turbulent sea. History and Technology, 39(2):208–224, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sturman:2023:THN

[881] Rachel Sturman. Technologies of a humble natural resource: The sand mining industry and marginal value in Bombay/Mumbai, 1920–2020. History and Technology, 39(3–4):225–253, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Esposito:2023:CCC

[882] James Esposito. Canaries, camouflets, and carbon monoxide: making ‘Proto Man’ in Britain’s tunnelling war 1915–1918. History and Technology, 39(3–4):254–279, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Lynam:2023:SNM

[883] John Lynam, Megan McNeil Zandstra, and Derek Byerlee. The search for new models for organizing regional agricultural research in the postcolonial era: Rice in West Africa. History and Technology, 39(3–4):280– 303, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Schatzberg:2023:MMR

[884] Eric Schatzberg. Missing the mark: a response to Bray and Hahn. History and Technology, 39(3–4):304–315, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See [867].

Beltramini:2023:WRM

[885] Enrico Beltramini. When religion meets history of technology: Secularism and the problem of the sacred. History and Technology, 39(3–4):316–327, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Wickberg:2023:CDA

[886] Adam Wickberg and Johan Gärdebo. Computation, data and AI in Anthropocene history. History and Technology, 39(3–4):328–346, 2023. CO-DEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Anonymous:2023:CSL

[887] Anonymous. Correction: [“The sparrow loves millet, but labors not“: Energy use and infrastructure in the Senegal Valley, 1450–1760]. History and Technology, 39(3–4):347, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See [871].

Anonymous:2023:CBM

[888] Anonymous. Correction: [Black metallurgists and the making of the industrial revolution]. History and Technology, 39(3–4):348, 2023. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic). See [870].